A Complete Bibliography of Publications in the Journal of Cell Biology: 1960–1964

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

04 June 2019 Version 1.00

Title word cross-reference

(λ_{26}) [Ryt60]. 2 [LA64, SS62b]. 3 [HK61a]. 5 [HZB63, HS61a]. 8 [Jac64]. ^{14} [CG61, vdD63]. ^{3} [HQ61, KF60b, MP64a, Sim64, Ton61]. ^{35} [CP60]. ^{∘} [SNE60, WA60]. _{z} [dBdMP62]. α [KF60a]. β [Ozz64]. • [Owe63, OM63].

-79 [WA60]. -acylamido [KF60a]. -Bromodeoxyuridine [HZB63]. -estradiol [Ozz64]. -Ethoxycaffeine [Jac64]. -Labeled [CG61]. -Leucyl-4-Methoxy-2-Naphthylamide [NMRS60]. -Me-Dab [LA64]. -Mercaptoethanol [SS62b]. -Methionine [CP60]. -Methyldeoxycytidine [HS61a]. -Phosphate [HK61a]. -Thymidine [HQ61, MP64a]. -Tyrosine [Sim64].

12 [DK60]. 17- [Ozz64].

3 [SNE60].

5 [KB60]. 5-Fluorodeoxyuridine [KB60].

732 [Erl64]. 79 [WA60].

929 [Fra62].

Abbau [Bie63]. Abdominal [Uch62]. Aberrations [Jac64]. Ability [Das62]. Abnormal [TY60]. Absorbed [Pal60]. Absorption [FS64, GOE60, Mil60, MP64b, NP60, Str63, FP60]. Absorptive [OVLG64]. Accumulation [BKS61, GRL64, Pea64, dW62]. Acer [WN64]. Acetabularia [KC63]. Acetobacter [CR64, ODH62]. Acetylcholinesterase [Bar62]. Acid [BS64a, Bas62, Beh63, Bie63, Dal63, DCPN64, Eds60a, Eds60b, EGS61, EK61, EG63, EN61, FT62, GB63, HH61a, HS61b, HS63b, HZ61, HE62, Ive64, KC63, KMTH60, McD62, MK60a, MF63, Mon64, MPH62, Osi64, Pai61, RC62, Sea61, SD64, Str64a, WL63, WSB64, WRS61, Ozz64, BM60, DK60]. Acid-Containing [GB63, HZ61]. Acid-Deficient [WL63]. Acid-Schiff [FT62]. Acidophilic [RS61b]. Acids [AK63, BB61, EB62, EH61a, SP64, WA61, Yam62, vdD63]. Acinar [Car61a, HF62a, HF62b, JM61, WHF62, Wes64, Cla60]. Acinous [Par62a]. Acoustic [Ros62b]. Acrididae [GR64]. Acridine [RM63, RMG64]. Acrolein [MJ63]. Acrosomal [CC61b]. Acrosome [Kay62b]. Across [LK63]. Acth [LM63]. Acting [FR61]. Actinomycin [JE64, PP64]. Action [GDLD63]. Activated [Dir61]. Active [GRL64, Nap63a, O’B64a]. Activities [BGSN60, EJ62, HM63b, MP64b, OVLG64, LY60]. Activity [Ack60, EN61, HD60, HK61a, HG63, Kar64, MB63a, MB63b, NM64, NdTBB62, Osi64, OVLG63, RBF^{+}61, RB62a, SC63a, Sea61, SF61, Ste61, TB62, WLD63, dBdMP62, CC60a, HMR60, KWCL60]. acylamido [KF60a]. Additional [Lan63]. Adenine [Hil64]. Adenosine [ALS63, EL62, EH63a, EH63b, Har64, PG63]. Adenosinetriphosphatase [TB62]. Adhesion [Cur64]. Adipose [BB60, Nap63b, Wil64]. Adjacent [Aro63b]. Administration [EMNB62, RBF^{+}61]. Adminstration [SS62a]. Adrenal [CZ63]. Adrenocortex [SD61]. Adrenocorticotrophin [Sip63]. Adrenocorticotrophin-Producing [Sip63]. Adult [BS62, Ber62, BBR61, CG63, CP60, DL63, GP63, KS64, LK61, BBR60]. Aedes [RP64]. Aegypti [RP64]. Aeshna [Smi61a]. Affect [GG62]. Affecting [RR64]. After [CZ63, EMNB62, KL64, LC63b, MF62, PL64, RBF^{+}61, Rah62, SJ64, Tro64, LA60]. Agar [MPH61]. Age [Lan64, MS63]. Agent [O’B64a]. Aggregated [MS60]. Aggregates [Ric61]. Aggregation [Sha64]. Aging [EB64, SBdlTL61, Ton61]. Air [WK64]. Air-Blood [WK64]. Albino [DG62]. Albumin [HHC64, vdD63, CAH60]. Albumin-Like [vdD63]. Alga [Bou62, Ech63, HG60c, RS61b, Sha60]. Algae [Gib62b, RS61a]. Algal [HC63]. Alkaline [Hal61, RBF^{+}61, RB62a, MDvD60]. Allium [CG61]. Allomyces [RS64b, AP60, BT60]. Alloxan [OM62b]. Alone [CZ63]. Alterations [AGZ61, LFB64a, RDM64, WS61b, AGC60]. Alternative [Nic61]. Alveolar [DBWK63]. Amacronucleate [KP64]. Amblystoma [Hay61]. Ameba [All60, BP60c, BP62, FM62]. Amebae [BPG63a, BPG63b, ROD60, RD62, Gri60]. Ameboid [All60, AC62]. Ameloblast [KSW63]. Ameloblasts [Rei61, Rei63]. Americana [And64, Hes60]. Amidase [HG63]. Amines [CZ63]. Amino [DCPN64, MPH62, SP64, vdD63]. Aminopeptidase [NMRS60]. Amoeba [CAH60, DR64, Ive64, RP62a, RP62b, RD61, WP64b]. Amphibian [JE64, Kar63, O’B64a, Too64, WK63, Wis60a]. Amphioxus [EW62, Pea61]. Amyloid [CC64b]. Anabaena [HG60c, Sha60]. Anacystis [Ech63]. Anal [Cop64]. Analysis [ABM63, BPA64a, BPA64b, Eds60b, GG64, Hir60, KC63, MPH61, Sko61, TS63b, WQSW60, LD60]. Anatomy [Mat60, Wis63, Wis60b]. Androgen [PF63b]. Androgen-Stimulated [PF63b]. Angular [GTL61]. Animal [Kon63, dT64]. Animals [CP60, Yas60, YTT60]. Annelida [CC60a, CC60b, CC61c, CC61a, CC61b]. Annulate [CW64, Kes63, Wis60a]. Annuli [Fel62b, Fel64]. Anomalous [HZB63]. Anoxia [OM62b]. Anterior [San62]. Anthers [HS61a]. Anthoceros [Wil63, Man60]. Antibodies [TH61]. Antibody [BPA62, EGG62, LF64, Pep61, PF61, PFH61, RTPM63, SS61a]. Antibody-Ferritin [RTPM63]. Antigenic [LF64]. Antigens [BPA62, CKP61, EGG62, MPH61, PCK64, TH61]. Anuran [CD61b]. Anurans [Mat60]. Aorta [MSBN63, Kee60b, Kee60a]. Aortic [WS61b]. Apatite [JP60]. Aplysia [Bat61, Ros63b]. Apparatus [BB64b, Har62, Har64, KSW63, Kan62a, MMMH61, MB63b, MW63a, MM64, WM63, Kan62b, WD61]. Appareil [Lal62]. Apparent [HL62a]. Appearance [SSK62, Pal60]. Applicable [Mun61a]. Application [ALM64, Lee62, Mit60]. Applications [Reb61]. Applied [HG63]. après [Ryt60]. Araldite [Par61, Stä63]. Arbacia [Kan60]. Arbusculus [RS64b]. Areas [MCT61, RP62c, RB62a]. Armadillidium [TC60]. Aroused [RM64]. Arrangement [Her64b, Tay64]. Arrays [TS64]. Arrest [Amo63]. Arterial [RB62a, WP64a]. Arthropod [Aro63a]. Articular [BCM64]. Artificial [Sch64]. Artificially [Dir61]. Ascidian [BS64b]. Ascites [Bas62, Bor60, EKY^{+}60]. Ase [LL61b, NdTBB62, OVLG63]. Aspects [Har62, Mer62, Rob61a]. aspersa [BH60b, MB63b, BH60a]. Assembly [Sha64]. Associated [AH63, GC63, Par63, Sea61, Ste60, Wat62, WA60]. Association [Dal62]. Astrocytes [Fri64]. Astrocytic [Mug64]. Asynchronous [ATD62, Pai61, Tay60]. Asynchrony [MD63, Pet64b]. ATP [BAS64, Epe63, HL62a, LL61b, NdTBB62, OVLG63]. Atpase [dBdMP62]. Atrial [JP64]. Atrophy [PF63a]. Attachment [GKH63]. Aurelia [FKKH60, KP64]. Aureus [Sug61, Sug64]. Autonomic [GP63]. Autoradiogaphy [Car62]. Autoradiographic [Mon62, MPH61]. Autoradiographically [Ton61]. Autoradiography [Bas62, CD63, CvT62, HQ61, KMF63, MSP64, MPH62, Nak64, SB64, SMV64, SHSG61, Sip63, WQSW60]. aux [Ryt60]. Avian [HB61, NdTBB62, TP64]. Axenic [NGD64]. Axonal [dW62]. Azo [LA64, RBF^{+}61]. Azotobacter [WNS61].

Bacilli [FJ60]. Bacillus [Han61, HD60, Mon61, OM62a, RLF62, WOM63b, Hol60, Van61]. Bacteria [BS63, CBPK64, PCB64, VR61, vT61]. Bacterial [Sch64, Tho64b, WOM63a, WS64, YFJ62, HG60a]. Bacterium [BD61, CR64, vIL64b, vIL64a]. Bai [NdTBB62]. Baileyana [Bou62]. Band [KH63]. bands [dBdMP62]. Barrier [WK64]. Bars [MP62]. Basal [RS64b]. Base [EB62, EG63, EH61a, HE62, Ive64]. Based [ZD62]. Basement [Fre62, Tho64a]. Bases [Las61]. Basidiobolus [Rob63b]. Basophilia [BT60]. Bat [IW63, Rev62]. Bats [RM64]. Beam [SHSS62]. Bearing [All60, HS63a]. Beat [Gib61]. Bee [Gol62]. Beef [SC64]. Beetle [SKdBH63]. Before [Rah62]. Behavior [Bou63, Ros62a, TJA63, HPWJ60]. Belonging [PCB64]. Berberine [Yam62]. Beta [PF63b]. Between [BH61, BS64b, BBP62, Dal62, DCPN64, ES63, Gib61, MP62, Nic61, Ols62, PKC62, Ros63c, SF64, See64, Str64a, Tor64, TG64, Yam62, KF60a, BT60]. Binding [Fel64]. Biochemical [HM63b, PM62, MRC60]. Biological [FM64a, DZ60]. Biopsy [Yas62a]. Birds [Cam64]. Birefringent [Aro63b]. Bismuth [AK63]. Bladder [Cho63, PR61, Sko61]. Blastocyst [HS60, HS62]. Blastomeres [Har61b]. Bleu [Lav60]. Blocks [Dan61]. Blood [Ack60, Ger64, Kar60c, MB63a, PT62, Tel61, TC64, WK64, Ham60a]. Blowfly [Smi63b]. Blue [Ech63, HC63, HG60c, Lav60, RBF^{+}61, RS61a, RT63, Sha60]. Blue-Green [Ech63, HC63, HG60c, RS61a]. Bodies [Beh63, Her64b, KS60, MB62, RS64b, RP61, HG60d, Man60, SPB61]. Body [BT62b, FS61a, Han61, Pea61, Rah62, SC63b, TT64, Voe64]. Bond [SHSS62]. Bone [Ack60, Boo64, CD63, DS61c, GK61, Sto62a, JP60]. Border [EM61, MS64b]. Botrytis [JE62]. Boundary [Ech63]. Bovine [HW61, SBdlTL61, TG64]. Brain [BSB63, BGSN60, BP63, Gra60, GY64, Mat63, SWD62, SSNB63]. Brains [Rob63a, RBS63]. Brasiliensis [WMG61]. Brat [Baw64]. Breakdown [Bie63, TS64]. Breaking [Fah63a]. Breaks [SNE60]. Breast [Cos64]. Briggsae [NGD64]. Bromodeoxyuridine [HZB63]. Brought [EB60]. Brown [CS64a]. Bufo [Cho63]. Bull [BZ64]. Bullfrog [Bon63, Sed61a, Sed61b, Sed62, Yam63b]. Burnsi [SM64]. Bursa [Ack62].

C [CG61, CS64b, SNE60, WA60, dBdMP62, vdD63]. C3H [LD60, SPB61, WD61]. C3H/Crgl [SPB61, WD61]. Caenorhabditis [NGD64]. Calcification [GSW64, MSBN63]. Calcified [Boo64]. Calcium [Cos64, EL62, GRL64, SF64]. Caldarium [MBM62]. Calf [KT63]. Californica [Bat61, Ros63b]. Called [Pel64]. Calliphora [Smi63b]. Canal [FW62b]. Canals [Loc61]. Capacities [Das62]. Capacity [PP60].

Capillaries [MB63a]. Carbohydrates [PL64]. Carbon [Epe63, Hum63, SHSS62]. carboxyl [KF60a]. Carboxypoly [SH60]. Carboxypoly-glucose [SH60]. Carcinogen [PH64, RNS63]. Carcinogen-Induced [PH64, RNS63]. Carcinogenesis [AGZ61, AGC60]. Carcinoma [Buc61]. Cardiac [GE60, Kar64, MS63, SS61b]. Cartilage [CD61a, CD63, SR60, SK62]. Cases [Gré61]. Casting [ODH63]. Cat [BBR61, MB61, BBR60]. Catechol [CZ63]. Catesbiana [Sed61b, Sed62, Sed61a]. Cations [Pea64]. Caulobacteraceae [PCB64]. Caused [HF62a]. Cavity [Rie62]. Cell [AH63, AGZ61, AP62, BP61, Bir61, Bou63, Bro62, BS63, CC64a, CP64, Ech63, EH63b, FP62, FM64b, Her63, HL62a, JE64, Kar60a, LP63, LK63, LK64, MMR62, Mal63, Mau61, Mon62, MPH62, Nev60, O’B64a, O’B64b, Ols62, OS64, Owe63, OM63, PDLP61, PBD^{+}61, Pel64, Per64, PHH63, Rey63a, RDM64, Rob61a, RM63, RMG64, RG64b, RG64a, Rob63a, RBS63, Rob63b, Ros63a, RP61, Sak60a, Sak60b, San62, SA64, STA64, SF61, Sed61a, Sed61b, Sed62, See64, SL62, Smi61b, SWD62, SP64, SL64, TMT^{+}61, Tay63, Tho64a, TC64, WM63, WSL64, You62, AP60, KG60, AGC60, Blo60, Fra60, Gre60, GOE60, Law60a, LY60, MK60b, Sha60, dHF60]. Cell-Associated [AH63]. Cell-Free [MPH62]. Cells [Ack60, Ack61, AB62, Bas62, BTAD60, BM64a, BH61, BM64b, BKS61, BGM64, BS64b, BS62, Ber62, BV63, BB61, BGW63, Bry63, Buc61, BT62a, CG61, Cam64, Car61a, CW64, CBB^{+}63, CF61b, Cop64, Cur64, DS61a, DF62, Dal63, DBWK63, DB64a, DL63, DS61c, EGG62, EK61, EB62, EH61a, EMNB62, End62, EES63, EL64, Eps61, Eps62, EH63a, FW62a, Fra62, FS61a, Ger64, GP63, Hag61, HHC64, HQ61, Har64, Hay61, HZB63, HN64, HF62a, HF62b, HW61, HH61a, HG63, HR62, HS63b, Ito60, Ito61, JS63, JP64, KG64, Kes62, KK62a, KMTH60, Kon63, LA64, LCOVG64, LW63, MS64b, MCC62, Mil60, MF62, Mil64, MM63a, MD63, MM64, MP62, MW63b, Nap63a, Nap63b, NB63, NGD64, NMW60, NMWVF61, OFS^{+}63, Osi64, OVLG63, OVLG64]. Cells [Pan62, Par62a, Par63, PC63, PH64, Ric61, Rie62, RR63, Ros62a, Ros63b, RCA62, RAC62, SF64, SJ64, SY60, SSK62, SS62b, SO62, Sip63, Sor62, Sto62a, Tho62, TT64, TG63, Wal63, WLD63, Wat62, WHF62, WA60, Wes64, WQSW60, Wis63, WG61, Yas62b, ZD62, dT64, MMB61, SSKD60, Tay62, WD61, Bor60, BP60a, BP60b, Cla60, DK60, EKY^{+}60, EB60, HPWJ60, LD60, MS60, RGB60, Tay60, VO60, Wis60b]. Cellulaires [BN60]. Cellular [Buc62, GSW64, KC63, LA60, MPH61, SL60, Ton61, Too64]. Cellule [Dro60b]. Cellulose [Col63, ODH62, OD64]. Centers [BM60, MHB60]. Central [BBP62, Hes60, NHK60, Pet60a, Pet60b, Pet61, Pet64a]. Centrifugation [Rah62]. Centrifuged [Zal61, GPN60]. Centriole [Gal61]. Centrioles [Ack61, Dir61, Sor62, Szo64]. Centrosomes [Ack61]. Cerebellum [Her64a, RBL62, Wol64]. Cerevisiae [VNL61, VL61]. Cernuum [CG61]. Certain [Beh63, Gra60, Nic61, GG60b]. Cervical [Kar60a]. Chaetognath [EW64]. chain [KF60a]. Chamber [She62]. Change [KL64]. Changes [BB60, BH60a, BL62, Bon63, CC61c, CC61a, DF62, EH61a, EB60, EMNB62, HH63, HF62a, HE62, KB64, KK62a, LA64, Nag63, SD64, Too64, Web64, WHF62, WA60, WL63, Tay62]. Chaos [FM62, Nac64]. Character [KC63]. Characteristics [MK60a, NN63a, NN63b, EKY^{+}60]. Characterization [MPH62, Gri60]. Characterized [MMR62]. Chelation [PSH60]. Chemical [AGZ61, Cha61, LGC61, MSBN63, SGW64, YFJ62, AGC60]. Chemically [HW61]. Chick [Ack62, BT63, DCPN64, OS64, MK60b, dL60]. Chicken [Nag62, DK60]. Chickens [CC64a, Cos64]. Chicks [HKA64]. Chinese [Tay60]. Chironomus [EB62]. Chlamydomonas [RP62c]. Chloride [PC63]. Chlorophyll [EK62, Sha60]. Chloroplast [GA63, Gib62a, Gib62b, RP62c, Wil63]. Chloroplasts [EK62, Kle60, MNP62, WT62, GOE60]. Cholinesterase [BMT63, Kar64]. Chondriome [Ste64]. Chondrocyte [GL64]. Chondrogenesis [GP60]. Chondroitin [Dzi62, TD63a]. Choroid [PT62]. Chortophaga [BB64a]. Chromatid [MP64a, SNE60]. Chromatin [ATD62, KS60, NTM60, Wat62]. chromatinic [HG60d]. Chromatography [HM62]. Chromatophore [Bou62, WS64]. Chromosomal [HM63a, MK60a, Pai61, Pet64b, TD63b]. Chromosome [BL62, Das62, ES63, HM63a, Jac64]. Chromosomes [CBBH63, CBB^{+}63, CS64b, CM64, EB62, EG63, Ger64, Hsu64, MD63, Rob61a, Tay64, Tay60]. Chylomicra [CS62]. Cilia [Afz61, Gib61, LL61a, LL61b, LMM61, RS64c, Sat63, Sor62]. Ciliary [Ols62, PT62, Tor63, Tor64, dL60]. Ciliate [Sea61]. Ciliated [GP63]. Cipangopaludina [Yas62b]. Cisterns [Ros62c]. citrate [Rey63b]. Clam [RS64a]. Clarkii [Mos61a, Mos61b]. Cleavage [BT62a, GC63]. Clotting [RPM63]. Club [Rob63a]. Coagulans [OM62a]. Coated [BH62]. Cobalt [PSH60]. Coccidioides [BHE61]. Cockroach [Hes60, WG61]. Coelicolor [GH61, GH60, HG60b, HG60d]. Cold [CS64a]. Cold-Exposed [CS64a]. Coleoptera [Smi60, Smi63a, SKdBH63]. Coleoptile [TT64]. Coleus [HN64]. coli [Car61b, CF61a, Ryt60]. Collagen [Cha61, GG64, Kee61, LGC61, MR63a, RB61, RB62b, RB62c, RB64, Tho64a, Yam60]. Collagenase [CC64b]. Collagens [JB60]. Collecting [Mil61a]. Collection [O’B64b]. Collets [Mun60]. Colloid [WSB64]. Colloidal [FM62, PT62]. Colonic [LCOVG64, OVLG63, OVLG64]. Color [Lee62]. Colorado [SKdBH63]. Colored [Hir60]. Colorimetric [Str62]. Column [HM62]. Combined [HQ61, HH61a, Str64a]. Combining [Stä63]. Comparative [DML61]. Comparison [DF62, Mil64, PKC62, SGW64, WMG61, UV60]. Compatible [FL62]. Competition [Str62]. Complement [Ton61]. Complementary [SS62b]. Complex [Amo62b, DS61b, MZ60, PL64, Ros61, SF64]. Complexes [AB62, FP63]. Component [Aro63b, BV63, FS61b, HR63, Lan63, Pet61, RWG60]. Components [AP62, SA64, STA64, TC64, WP64a, KZ60]. Composition [Eds60a, EB62, EG63, EH61a, HE62, Mer61, WOM63a, WOM63b]. Compound [Gol62]. Concentration [Car61a, EL62, MK60b]. Concentric [RV62]. Concerning [JSB63]. Conchiolin [Gré61]. Conductance [LK63, PP60]. Configuration [Fah64, Tor64]. Conjugated [BPA62]. Conjugates [EGG62, LF64, RTPM63]. Conjugation [SS61a]. Connections [BS64b, BBP62, Ove63]. Consideration [CD61a]. Consistency [All60]. Constant [CG63, Cam64]. Constituent [FL62]. Constituting [FaP64]. Containing [Beh63, BGM64, GB63, HZ61, MB62, RP62c, TT64, Sha60]. Content [EK62, EH61b, FHI63, Gal62, Ive64, Nic61]. Continue [BN60]. Continuities [NB63]. Contracted [CKB61, TH61]. Contraction [Gil62, GR64, LFB64b]. Contrast [Buc62, Gon62, Par61, RP60, Ros63c]. Contraste [BN60]. Contribution [Fah63a]. Control [Gon62, Riz63, SKdBH63]. Controlled [Hei62, Kan62b]. Conversion [KPM61]. Convoluted [Mil60]. Cooling [FM64a, WA60]. Copepod [Fah63b]. Cord [BBR61, Hes60, BBR60]. Cordylophora [Ove63]. Cornea [HL62a, KP62, KPDM62, Kay62a]. Corpus [FS61a, Mug64]. corpuscles [CR60]. Correction [Per64]. Correlated [Das62, LC63a, Pac63a, Pac63b]. Correlation [MBH63, SF61]. Cortex [RBF^{+}61, Sim64]. Cortical [Lan64, Mil61a]. cortisol [SPB61]. Course [MHB60]. Coverslips [NR60]. Crab [Lav60]. Crabe [Lav60]. Crassa [Luc63]. Crayfish [BK62, BK63, Mos61a, Mos61b, PP61, Pet63, Uch62]. Crgl [SPB61, WD61]. Cricket [Kay62b, Kay64]. Cricothyroid [Rev62]. Cristam [FS61a]. Cristatus [TD63b]. Croft [HS61a]. Cross [MS64a]. Cross-Linked [MS64a]. Crustacea [TC60]. Crypt [Tri63]. Cryptorchidism [Nag63]. Crypts [Bry63]. Crystal [NEO63, TT64, Wat63]. Crystal-Containing [TT64]. Crystal-Matrix [Wat63]. Crystalline [LH62, Sto62b, TT61, Gri60]. Crystallites [JP60]. Crystalloid [HS60]. Crystals [FM64b, NHK60, TG64]. Ctenophore [Afz61]. Cultivated [KK62a, NMWVF61]. Culture [Har64, HKA64, KG64, KMTH60, MCC62, NGD64, Ros62a, RBL62, TG63, SSKD60, NR60]. Cultured [BHE61, HZB63, Jac64, MMB61, RGB60, Tay60]. Cultures [CP60, DL63, Rob61a, RP60, Tay63, Wol64, DK60, LY60]. Cuticle [Loc61, MR63a]. Cuticular [EM61]. Cyanidium [MBM62]. Cyanophora [HC63]. Cycle [CC64a, Mon62, RG64b, SK61b, Ste61, Tay63, WRS61, EKY^{+}60]. Cycles [DF62]. Cyclopoid [Fah63b]. Cylindrica [HG60c]. Cyst [WNS61]. Cysteine [EMNB62]. Cytidine [RC62]. Cytochemical [AB62, FA64, Fra62, Hal62a, Hal62b, Har64, HD60, HH61a, HH61b, JSB63, LW63, PG63, RT63, RM64, SP62, Str64a, WS61a, WS63, vIL64b, vIL64a, HHP60, Mit60, PSH60, SP60a, SP60b]. Cytochemistry [LB61, MJ63, SBB63, ZB61a]. Cytochrome [LY60]. Cytodifferentiation [RR62, Riz63]. Cytological [CS64a, EN62]. Cytology [MR63b]. Cytolysis [O’B64a, GG60c]. Cytolysomes [Nap63a, NE62]. Cytonucleoproteins [BPG63a, BPG63b]. Cytopathology [NMWVF61]. Cytophotometry [FS64]. Cytoplasm [All60, Buc62, EB62, EG63, FS61b, GG62, HZB63, HN64, Nac64, RD61, BT60, GPN60, MK60b]. Cytoplasmic [AP62, BS64a, Beh63, BV63, Dru63, Far64, FJ60, HM62, Ito60, KZ60, Lou62, MK60a, MCC62, MB62, MZ60, Per61, PF63b, PBK62, RP62a, RP62b, SSK62, Sla63, SPB61, WN61, WP64a, WP64b, dT64, BB64a, HMR60, RWG60]. Cytostome [SN60].

D [JE64, PP64]. D.E.R.(R) [Erl64]. Dab [LA64]. Damage [ES63, Par62b]. Danica [Gib62a]. Dans [Dro60b, Lav60]. Darstellung [Sto62a]. Daudin [Pet60b]. Deamination [HS61a]. Decalcification [LK61, Wat63]. Decemlineata [SKdBH63]. Decrease [SBdlTL61]. Deer [MR63b]. Deferens [MBH63]. Deficiency [MNP62, SD64]. Deficient [WL63]. Defined [HW61, LW62]. Degenerating [Hes60]. Degeneration [BM64b, NE62, SKdBH63, TS64, dW62, Ehr60]. Degenerative [HF62a]. Degradation [TS63a]. Degrees [CKB61]. Dehydration [CHG61, Fin60b, Fin60c]. Dehydrogenase [Ack60, FA64, HD60, HK61a, RT63, WS61a, LY60]. Dehydrogenases [Wal63, WS63, PSH60]. Demineralisation [Boo64]. Demonstration [Ack60, Beh63, BBP62, EGG62, HK61a, MB63a, OVLG63, Pea62, WS61a, WS63, MDvD60]. Demyelination [BBR60]. Denervated [BMT63]. Denervation [PF63a]. Dense [FS61a, Law60b]. Densitometric [TS63b]. Density [BS60, FHI63, Rah62]. Density-Gradient [Rah62]. Dentine [JP60]. Deoxycytidine [HS61a]. Deoxyribonuclease [Ste61]. Deoxyribonucleic [BB61, BM60, Dal63, EK61, HS61b, KMTH60, McD62, MF63, Pai61, DK60]. Deoxyribosides [HS61b]. Dependence [PP64]. Dependent [WS63]. Depletion [SF64]. Deposition [HN64, PCDMP61, Tho64a]. Deprivation [WHF62]. Deprived [SP64]. derivative [Man60]. Derived [TS64]. Description [Fra62]. Desmosome [TS63b]. Desmosomes [GE60, Ham60b]. Detecting [BPA62]. Detection [Pan62]. Determination [EK61, HLG63, NTM60, WLD63]. Determined [HC61, TH61]. Developing [Aro61, BK63, BBP62, Cos64, DCPN64, HS61a, KF60b, Nag62, MK60b]. Development [BT63, Bou62, BT62a, DS61a, Dal63, Gib62a, GTWB63, GSA62, Kle60, KBB64, KMTH60, MW63b, RS64b, TG63, WN64, WNS61, BH60a]. Developmental [CKP61, PCK64]. Device [NR60]. Dextran [BCM64, DS61b]. Diabetes [OM62b]. Diamox [Tor63]. Diaphorase [HD60]. Diatom [Dru63]. Diazotype [Lee62]. Didelphis [RL64]. Differences [RM64, Tor64]. Different [CKB61, DML61, LMM61, Str64b, dT64]. Differentiating [BM64a, KG64]. Differentiation [And64, Beh63, GG62, Nap63b, PP64, TD61, Wes64, Wil63, Wol64]. Differentiations [HM63a]. Differing [SSK62]. Diffraction [Fin60b, Fin60c]. Diffusion [She62]. Digestion [Str64b, TW61]. Dimensional [JP60]. Dimensions [OD64]. Dimer [MR63a]. Dimethylnitrosamine [EB60]. Diphosphate [EJ62, Luc61]. Diphosphopyridine [DHV61]. Diplococcus [TJO64]. Diploid [KS60, MD63]. Diptera [Smi63b]. Direct [MMMH61]. Direction [Gib61]. Directional [Flo64]. Dirt [PD60]. Dirt-Free [PD60]. Disc [GE60, KC62, dBdMP62]. Discharge [CP64]. Discocephalinids [Nic61]. Disease [SSK62]. Disintegration [RPM63]. Disposition [Tha63]. Dissociability [KMTH60]. Dissociated [DL63]. distinguishing [KF60a]. Distribution [Cos64, DB64b, DCPN64, EGS61, Fel62a, JSB63, LW62, RW64, SS62a, TH61]. Distributions [Per64]. Distrophy [LD60]. Disturbance [SJ64]. Divalent [Pea64]. Dividing [EKB62, Ito60, Pel64, BM60]. Division [GA63, Pel64, Rob63b, Sak60a, Sak60b, SP64, Tay63]. Dl [HF62a]. Dl-Ethionine [HF62a]. DNA [AB62, ATD62, BGM64, Bry63, CG63, Cam64, CM64, EKY^{+}60, ES63, FHI63, GJ60, Gal62, Gar64, Ger64, HC61, dFB62, Mon62, MD63, NN63a, NN63b, Nic61, Pel64, Pet64b, PP64, PKC62, PBK62, RP62a, RP62b, Rie62, RP62c, SBdlTL61, Sch64, See64, SK61b, Sto61, SP64, TW61, Wes64, WP64b]. DNA-Containing [BGM64, RP62c]. DNP [HR63]. Dog [SF61]. Dogs [WS61b]. Dollar [SK61a, KB60]. Domestic [CM64, Nag62]. Domestica [Baw64, Bie63, Bie63]. Domesticus [CM64]. Dose [HR62]. Doses [MP64a]. Dpn [WS63]. Dpn- [WS63]. Dragonfly [Smi61a]. Dried [SY60]. droplet [WD61]. Droplets [MP64b, NP60, WSB64, FP60, Pal60]. Drosophila [Aro62, Aro63b, DKG64, HM63a, Ito60, MK60a, Riz61, RR62, RR63, Sha63a, Sha63b]. Dry [BOS62]. Drying [FM64a]. Duct [EM61, She62]. Ductal [MB61]. Dugesia [Mac63]. Duodenal [OS64]. Duodenum [Beh63]. Duplication [BM60, MHB60, Tay60]. Duration [CG63, Cam64]. Durcupan [Stä63]. During [Amo62b, AGZ61, Aro63b, BS64a, Beh63, BT63, Bon63, BB64b, CC61c, CC61a, DF62, EB64, Gil62, HH63, HS63a, HS63b, HE62, KB64, Kon63, LC63a, Mon64, Rei61, Rei63, RDM64, RG64b, RPM63, Sak60a, Sak60b, SD64, Tho64a, TS64, Ton61, Too64, WHF62, WRS61, You62, dW62, BH60a, BB64a, Her60, AGC60]. Dye [LA64]. Dyes [RBF^{+}61]. Dynamic [Mer62]. Dynamics [LC63b, RM63, RMG64]. Dystrophic [Cos64, HKA64]. Dystrophy [DS62b].

E. [Ryt60]. Earle [DS61a]. Early [Dal62, GC63, KBB64, Rei63, dW62, BH60a]. Earthworm [MR63a, Ham60a]. Eberth [CD61b]. Eccrine [EM61, MB61, Mun61b]. Echo [NMW60, NMWVF61]. EDTA [KG60, LFB64b]. Edulis [Gré61]. Effect [CC64b, EK62, Fel64, GRL64, HR62, Kle60, LD63, MNP62, OM62b, SC64, WN61, KG60, MMB61, Kan60]. Effective [Hol60]. Effects [CZ63, CS64b, Epe63, Fin60b, Fin60c, HW61, JE64, KB60, LK61, Ozz64, RLF62, SF64, SS62b, Str62, SL64, TC64, Tor63]. Effets [Lal62]. Egg [BPA62, BPA64a, BPA64b, BM64b, Bel61, CC60a, Epe63, NP60, Sak60a, Sak60b, Str62, GPN60]. Eggs [Amo63, CKP61, Dir61, Hin62, MF63, PCK64, BM60, Kan60]. Ehrlich [Bor60, EKY^{+}60]. Eighth [RP61]. Einzelzellen [SSKD60]. Eisenia [Ham60a]. Electrical [LK63, PPS60]. Electron [Ack62, AMFW60, AK63, BNH62, BC61, Baw64, BPA64a, BPA64b, Bay64, BK63, BH62, BM64a, Bel61, BL62, BP60c, BP62, BHE61, BT62a, BBR60, BBP62, BS60, CH64, Car61a, CSC61, Cla60, CFG62, DS61a, Dal62, DR64, DML61, Dan61, Dro60a, Ehr60, EM61, EES63, EH63a, EH63b, EN61, FL62, Far64, FR62, Fin60b, Fin60c, Fra60, FS62, FM64a, GP60, GK61, Gon62, GTWB63, GY64, GPN60, Hag61, HC61, HQ61, HB61, Har61b, Hei62, HH61a, HH61b, HZ61, IC63, JS63, JP60, Kar64, Kar60b, Kee60b, Kee60a, Kes63, Klu62, KT63, KMF63, LC63a, LA64, Lar62, LdR60, Las61, Lee62, LC62, Mal63, MB63a, Mar61, Mat63, Mat60, MCT61, MS60, MF62, Mun60, MB61, MW63b, Nak64, NdTBB62, NMW60, NMWVF61, Odo60, Osi64, PMRGG62, PT62]. Electron [PD60, PDLP61, PBD^{+}61, Par62b, Pea64, PF63a, Pep61, PF61, PFH61, Per64, PCDMP61, Rah62, RTPM63, Reb61, RNF60, RS61a, RG64a, RLF62, RBL62, ROD60, RD62, SBB63, SB64, San62, Sch64, Sed61a, Sed61b, Sed62, SHSS62, SJ64, SY60, Sha63a, Sha63b, SMV64, SHSG61, SS61a, Stä63, SW64, Sto61, Sto62b, SC63b, TKMT61, TS63b, TW61, VR61, VO60, WMG61, Wat63, WA61, WR64, Wet61, Yas62a, Zal61, ZM60, ZB61b, vT61, Erl64, Kar61, KG60, Rey63b, Blo60, DZ60, Ess60, Fin60a, Hal60, Her60, Hol60, HA60b, Law60b, LBB60, Mit60, MDvD60, Sto62a, Yas60, YTT60, dHF60]. electron-opaque [Rey63b]. Electron-optical [Sto62a]. électronique [Ryt60, Dro60b, Lav60]. Electrophoretic [Eds60b, Hir60]. Electroplaques [WMG61]. Electroscopic [Lav60]. Elektronenoptische [Sto62a]. Elemental [Yas62b]. Elements [NTM60, Uch62]. Elongated [ODH63]. Elongation [Aro62]. Embedded [MHR61, Law60b]. Embedding [BP61, CHG61, CFG62, FL62, Luf61, SKF63, Stä63, NR60]. Embedment [FS62, Erl64]. Embedments [LC62]. Embryo [DCPN64, TG63, DK60, KB60]. Embryogenesis [BT63]. embryonal [Ozz64]. Embryonic [Ack62, HB61, JE64, See64, TG64, BH60a]. Embryos [Har62, Hin62, SK61a]. Emission [Hal62a]. Emulsion [Bas62]. Enamel [KSW63, NEO63, Rei61, Rei63, TG64, Wat60]. Encephalomyocarditis [DF62]. Encystment [BS64a]. End [BMT63]. Endings [DPRG61, DdI61, Rob63a]. Endochondral [You62]. Endometrium [EES63]. Endoplasmic [BK62, BM64a, EMNB62, ESP62, Her64b, Ito61, Lav60, MB62, PY60, RP60, SJ64, PM60]. Endoplasmique [Lav60]. Endothelia [WP64a]. Endothelial [MP62]. Engaged [Sko61]. Enhanced [HE62]. Entamoeba [BS64a]. Enteropneusta [CC63a, CC63b]. Entry [Tel61, CC60a, CC60b]. Enucleated [MF63]. Envelope [Gib62b, MM61, Mer61, Mer62, WML60, Pal60]. Envelope-Chloroplast [Gib62b]. Envelopes [BS63]. Environment [But62]. Enzymatic [DBWK63, Fri64, HH63, MB63a, TW61]. Enzymatically [GOE60]. Enzyme [BGSN60, ESP62, HG63, KH63, PBK62, RR63, SS62a]. Enzyme-Structure [ESP62]. Enzymes [DB64a, ER62, EH63a, EH63b, EH61b, HK61b, HT64, NSD61]. Eosinophilia [Lit64]. Ependyma [BP63]. Epidermal [Dro60a, Fre62]. Epidermis [CD61b, FS61a, Hay61, PM64]. Epineurial [Ros63b]. Epithelia [CG63, FP63, FS61b, LC63b, SL60, LA60]. Epithelial [Ack62, Bro62, HQ61, LK64, MF62, Mil64, OVLG63, Wal63, WSL64, WD61, BP60a, BP60b]. Epithelium [Beh63, Bon63, DG62, Far64, HL62a, Lar62, MB62, SB62, Tor63, Tor64, Tri63, TP64, WSB64, Kar60a]. Epithelium-Stroma [HL62a]. Epon [CFG62]. Epoxy [Fin60a, FS62, Luf61, MHR61, Stä63, Law60b]. Epoxy-Embedded [MHR61]. Erectum [HS63b]. Ergastoplasm [Par62a]. Ergosterol [TT61]. Erythroblast [BT62b]. Erythrocephala [Smi63b]. Erythrocyte [DS62a, O’B64a, TD63b]. Erythrocytes [Dav61, GSA62, LF64, Too64]. Erythroid [RDM64]. Erythrophagocytosis [Ess60]. Escherichia [Car61b, CF61a]. Especially [BM64a, EM61]. Essential [SD64, SP64, WL63]. Essentially [CG63]. Established [GG64, TG63]. Esterase [HG63, WSB64]. Estimation [Lou62]. Estradiol [WN61, Ozz64]. Etched [MM63a]. Ethionine [HF62a, HF62b, SS62a]. Ethoxycaffeine [Jac64]. Etiolated [EK62, KPM61]. Étude [Lav60, Ryt60]. Euglena [GG62]. Euplotes [Klu62, PKC62]. Eurystomus [Klu62, PKC62]. Evaluation [CHG61]. Events [EKB62]. Evidence [AB62, BB64a, CG63, FP62, MF63, MB61, TJA63, CC60b, Dro60b]. Evolutionary [Sat62]. Examination [Bay64]. Examined [DS61a]. Excess [GDLD63]. Exchanges [Fel62b, FM62, Fel64, MP64a]. excretory [SL60]. Existence [FP62, SN60]. Exocrine [CP64, KG64, WLD63]. Exoerythrocytic [HPWJ60]. Exogenous [RW64]. Experimental [BBR61, Lit64, Nag63, TH64]. Experimentally [EH61a, BBR60]. Experiments [SM64]. Exposed [CS64a]. Exposure [MP64a]. Extensibility [Ros63a]. Extracellular [LW62, Wat60]. Extract [CC60a]. Extractable [JB60]. Extraction [Eds60b, Kan60]. Extramitochondrial [DHV61]. Extraperiosteal [TC64]. Extremely [BS63]. Extrinsic [Reg61]. Extrusion [HS62, Str64b]. Extrusions [Cla60]. Exudate [DBWK63, DB64a]. Eye [EW60, EW64, Gol62, Reg61, WG61].

Faba [CG61, LC63a, WRS61, SNE60]. Fabricius [Ack62]. Factors [GG62, RR64]. fallax [HPWJ60]. Family [PCB64]. Far [KBB64]. Far-Red [KBB64]. Farlow [Bou62]. fascifrons [SL60]. Fast [FR61]. Fast-Acting [FR61]. Fat [CS64a, Str63, Pal60]. Fatbody [Riz61]. Fate [BGM64, EB64, Sim64]. Fatty [SD64, War62a, War62b, WL63]. Feather [BT63, FR62]. Features [Eps61, Van61]. Feeding [CC64a, EK62]. Fern [GA63]. FernÁNdez [Bul60]. Ferritin [BPA62, EGG62, Fel62a, LF64, RTPM63, Ric61, RW64, SS61a, FP60]. Ferritin-Antibody [EGG62, LF64]. Ferritin-Conjugated [BPA62]. Ferrocyanide [LW62]. Ferrous [Amo62b]. Fertilization [CC61c, CC61a, CC63a, CC63b, BH60a]. Fertilized [SK61a]. Fetal [FVK61, GSA62]. Fetus [RG62, SS62a]. Feulgen [FS64, Gar64, SBdlTL61]. Feulgen-DNA [Gar64]. Feulgen-Positive [SBdlTL61]. Fiber [CKB61, MA61, TC60, Ver61]. Fibers [FaP64, Hes60, Lav60, Mau61, NN63a, NN63b, PG63, PH62, PF63a, RWG60]. Fibre [Bat61]. Fibres [BMT63, Lav60]. Fibrillar [Nac64, Smi60, Smi61c].

Fibrils [Aro63b, CC64b, HN64, Kee61, Pea63, Yam60]. Fibrinogen [HHC64]. Fibroblast [DS61b, GG64]. Fibroblasts [Sor62, Voe64, Ozz64]. Fibrous [TMT^{+}61]. Field [Nak64]. Figures [CD61b]. Filament [PH63]. Filamentous [OB64c, AP60]. Filaments [MS64b, Mug64, RS64c]. Films [Hum63, Lee62]. Filter [DL63]. Filter-Well [DL63]. Filtrates [dHF60]. Fine [Afz61, AGC60, AGZ61, ALS63, Bar62, BTAD60, BS64b, BS62, BD61, Boa64, Bon63, Bou62, BP63, BT62b, Cam61, CD61b, Cho63, CF61b, CR64, CBK63, CBPK64, CM64, CC61b, DF62, DG62, Dro60a, Dru63, DS61c, EW60, EW62, EW64, Ech63, EK62, End62, EL64, Eps61, FR62, Fre62, FM64b, Gib61, GH60, GH61, GKH63, Gol62, Hay61, HR63, Her63, Her64a, Hes60, HG60c, HA60a, IW63, KG64, Kay64, KS64, Kes62, KPM61, KBB64, KB64, KT61, LL61a, LP63, Mac63, Mat60, MJ63, Mer61, MBH63, MM63a, MM63b, Nad64, Nag62, Nag63, NM64, OS64, PM64, PP61, PC63, PCB64, RNS63, Reg61, RK63, RS61b, RP61, Ros62b, Ros63b, RL64, SC63a, SF61, SHT61, SO62, Tho62, TB62, TJO64, Tor63, Tro64, TC60, VR61]. Fine [Ver61, Yam63b, ZB61a, CR60, HG60b, BT60, BP60a, BP60b, Ehr60, EB60, Ham60b, Ham60a, HGCN60, Law60a, Mer60, MR60, TY60, VO60, dL60]. fingers [CR60]. Fire [Baw64]. Fire-Brat [Baw64]. Firefly [Smi63a]. Fish [FR61, FW62a, FW62b, FaP64, RT62]. Fixation [CHG61, DS62a, HH61b, MZ60, PMRGG62, Rob61a, RLF62, Sch64, Too64, WS61a, WS63, Wet61]. Fixative [MJ63]. Fixed [JM61, Mun61a, OVLG63, Ros63c, Tor64, ZM60]. Flagella [Kay64, Lan63, RS64b]. Flagellar [Pea63, GG60b]. Flagellates [GG60b]. Flagellum [GKH63, Nag63]. Flight [Aro62, Sha63a, Sha63b, Smi60, Smi61a, Smi61c, Smi63b, SKdBH63]. Fluorescein [TH61]. Fluorescein-Labeled [TH61]. fluorescent [CAH60]. Fluorodeoxyuridine [KB60]. Fly [KF60b]. Focal [dW62]. Foetal [Beh63]. foetida [Ham60a]. Follicle [Fra60]. Follicular [FM64b, Wis63, Wis60b]. Followed [Dal63]. Following [BCM64, CC64a, DF62, Fel62a, HF62b, MP64a, NHK60, SS62a]. Foot [MB61]. Formaldehyde [Hir60]. Formalin [HH61b, OVLG63, WS61a, WS63]. Formalin-Fixed [OVLG63]. Formation [Cha61, CC60b, FJ60, KSW63, Kay62b, Kee61, LGC61, Luc63, Mon61, OM62a, Pet60a, Rei61, RB61, RB62b, RB62c, RB64, RS64c, Sor62, WW61, WM63, YFJ62, WD61]. Formations [Par62a]. Formed [HM63a]. Found [TC60]. Four [HC61]. Four-Stranded [HC61]. Fowler [Han61]. Fraction [EL62, LBDL63, Nev60, Sha60]. Fractionation [HW61]. Fractions [HW61, MPH61, Rah62, WN61, WS64]. Fracture [Hag61]. Fractures [GK61]. Fragments [Das62]. Free [CMRS62, Fel62c, KK62a, MPH62, PD60, WT62, GOE60, Gri60, dHF60]. Freeze [Bul60, FM64a, Reb61]. Freeze-Drying [FM64a]. Freeze-Substitution [Bul60, Reb61]. Freezing [MMWFW61]. Freezing-Ultramicrotome [MMWFW61]. French [Lav60]. Frequency [MP64a]. Frog [KS64, MA61, MACAvdD61, PH62, SC63a, Ham60b]. Frogs [SM64]. Frozen [Hag61, MM63a, SY60]. Frozen-Dried [SY60].

Frozen-Etched [MM63a]. Fruit [KF60b]. Fully [Boo64]. Function [BB64b, EMNB62, FMOBG64, FW62a, HH63, Rot60b, SC64, Szo64, Rot60a]. Functional [EN62, FP62, Har62, SF61, Sko61]. Functioning [MW63a]. Functions [DBWK63]. Fundulus [PC63, Reg61]. Fungus [Rob63b, AP60, BT60]. Furrow [BT62a]. Further [EW60, Pet64a, San62, TMT^{+}61, dHF60]. Fusion [Jen64].

gallinaceum [HPWJ60]. Gallus [CM64]. Gamete [CC63a, CC63b]. Gametophyte [GA63]. Gamma [SNE60]. Ganglia [Ros62b, Ros63c, RW64, TC60]. Ganglion [BTAD60, RP61, Uch62, Yam63b, dL60]. Gas [Hei62]. Gastric [Ito61, IW63, SF61, Sed61a, Sed61b, Sed62, SF64, Law60a]. Geer [BB64a]. Gene [SM64]. Genes [Riz64]. Genetic [Riz63, LD60]. German [Bie63, SSKD60, Sto62a]. Germination [WNS61]. Gewebekultur [SSKD60]. Giant [Bir61, DR64]. Gill [Gib61]. Given [CZ63]. Gland [DdI61, EB62, EES63, JS63, LK64, WSL64, Wis63, Wis60b]. Glands [CZ63, EM61, MK60a, MB61, Mun61b, MR63b, PM62, Sed61a, Sed61b, Sed62, SF64]. Glass [Cur64, Fah63a]. Glia [BBP62, HH63, HL62b, Las61]. Glial [RV62]. Globular [Hal60]. Glomerular [FP60]. Glomerulus [FP62, Yam60]. Glucose [Luc61, PL64, SH60]. Glucuronidase [PF63b]. Glucuronidase-Rich [PF63b]. Glutaraldehyde [Tor64]. Glutaraldehyde-Fixed [Tor64]. Glutinosa [WR64]. Glyceraldehyde [HK61a]. Glyceraldehyde- [HK61a]. Glycerin [Hum63]. Glycerinated [GR64, NM64]. Glycogen [Luc61, RNF60, SSK62]. GlycogÉNe [Dro60b]. Goblet [LCOVG64, OVLG64]. Gold [FM62]. Golden [Str63]. Goldfish [LM63, Rob63a, RBS63, RP61]. Golgi [Car61a, KSW63, Man60, MB63b, MZ60, MW63a, MM64, PL64, Ros61, RM64, WD61, WM63, ZD62]. Gonadotropic [HMR60]. Gradient [FHI63, Rah62]. Gram [CR64, vIL64b, vIL64a]. Gram-Negative [CR64, vIL64a]. Gram-Positive [vIL64b]. Grana [WT62]. Granular [FS61b, Her64b]. Granule [Lil61, LBDL63, SL62, Wat62]. Granules [BK62, Beh63, But62, Dro60a, HM62, HM63b, JP64, Mil61a, Pea64, PM62, Tho62, WSB64, HMR60]. Graphite [BH62]. Graphite-Coated [BH62]. Grasshopper [TD61, Tay64, BB64a]. Green [CBPK64, Ech63, HC63, HG60c, RS61a, Sha60]. Gregarium [Szo64]. Grids [BH62, PD60]. Ground [GPN60]. Groups [BH61, Sak60a, Sak60b, KF60a]. Growing [PHH63, Rie62]. Grown [HW61, NR60]. Growth [BS64a, FJ60, NEO63, Rob63b, TG63, Ton61, MMB61, Ozz64, SSKD60, Gre60]. Guinea [AH64, DL63, MBH63, SP62, SP60a, SP60b].

H [Ton61, HQ61, KF60b, MP64a, Sim64]. Haematolœchus [SHT61]. Haemoglobin [CSC61, TD63b]. Hair [Nak64, Sim64]. Hairs [FW62a]. Hale [GB63]. Halogenated [KB60]. Halophilic [BS63]. Hamster [RNS63, Str63, HA60b, Tay60]. Harv [Bou62]. HCl [Sug64]. Healing [GK61, RB61, RB62b, RB62c, RB64]. Heart [DHV61, OM62b, RK63, SC64, TB62, dBdMP62]. Heavy [BG64, Sto61]. HeLa [SSKD60, Eps61, Eps62, EH63a, EH63b, Har64, HZB63, KK62a, LW63, Pai61, Ros62a, SSKD60, SS62b, SL64]. Helical [Mug64]. Helium [Bul60]. Helix [BH60a, BH60b, MB63b]. Hemoglobin [Mil60]. Hepatectomy [KL64, Tro64]. Hepatic [AP62, EN61]. HÉPatique [Dro60b]. Hepatoma [PH64, KZ60]. Hepatomas [EN62]. Herpes [Eps62, EH63a, LW63]. Heterochromatic [KS60]. Heterochromatin [dFR63]. Heteroclitus [Reg61]. Heterogeneity [EGS61]. Heteroptera [Nic61]. Heterozygous [HKA64]. Hexagonus [CC61c, CC61a, CC61b, CC60a, CC60b]. Hibernating [RM64]. High [BH62, CvT62, Car62, FHI63, RB62a, Kar61, Rey63b]. High-Resolution [CvT62, Car62]. Highly [Pea62]. Histoautoradiographic [DK60]. Histochemical [Ack60, DBWK63, HK61a, HG63, WMG61, KF60a, NMRS60]. Histochemistry [Ack61, Bir61, GG60a]. Histological [FT62]. Histometric [TS63b]. Histone [DKG64, BB64a, BH60b]. histones [BH60a]. Histophysiology [Mun61b]. Historadiographic [Hal61]. Historadiography [Hal62b]. Holder [Mun60]. Holes [CC60b]. homoptera [SL60]. Honey [Gol62]. Honey-Bee [Gol62]. Hormonal [HM63b, SKdBH63, Her60, KWCL60]. hormone [Ozz64]. Hormones [FT62]. Horseradish [Str62, Str64a, Str64b]. Host [HPWJ60]. Hosts [Dal62]. House [Kay62b, Kay64]. Housefly [Bie63]. Human [Ack60, Ack61, ATD62, BNH62, Bro62, CS64b, Ger64, HHC64, HW61, HG63, IC63, Jac64, KS60, LF64, Lil61, dFR63, MS63, MCC62, MD63, MB61, Mun61b, NTM60, NB63, OB64c, RK63, Sto62a, Yas62a, ZB61a, CR60, Ozz64, JP60, Kar60a, RGB60]. Hyaluronidase [CC64b]. Hydei [HM63a]. Hydrazide [CG61]. Hydroides [CC61c, CC61a, CC61b, CC60a, CC60b]. Hydrolysis [Eds60b]. Hydrolytic [DB64a, HK61b]. Hydroxide [Fel62c, DZ60]. Hyperplastic [FS61a]. Hypertonic [MF62]. Hypertrophic [Pan62]. Hypertrophy [RK63]. Hyphae [Zal61]. Hypo [MF62]. Hypo- [MF62]. Hypophysis [Sip63]. Hypothalamic [Osi64]. Hypoxia [HH63].

Identical [Amo63]. Identification [CS62, Hal61, Hal62a, Hal62b, PH62, RNF60, Sip63, Gri60]. Ii [Bul60, OM63]. Ileum [LC63b, LA60]. Im [Bie63]. Imaging [Sto62a]. Immature [Wis60a]. Immitis [BHE61]. Immune [GG60c]. Immuno [BPA64a, BPA64b, RTPM63]. Immuno-Electron [BPA64a, BPA64b, RTPM63]. Immunological [FKKH60]. Impregnation [Mar61]. Improved [SW64]. Improvement [NMRS60]. Improvements [Luf61]. Inclusion [HS60, SPB61]. Inclusions [HS62, Kes62, Mon61, TP64, Gri60]. Incorporation [CG61, CP60, DCPN64, KF60b, MPH62, NGD64, Per61, RC62, Sim64, WQSW60]. Increase [RBF^{+}61]. Increased [HH63, Par61]. Indirect [Aro62, Sha63a, Sha63b].

Indium [WA61]. Individual [EK61, WQSW60]. Induced [Aro63b, BB60, Bir61, EH61a, KK62b, LA64, PH64, RNS63, SNE60, BBR60, dHF60]. Inducibility [Riz64]. Inducible [RR63]. Induction [LM63, Ste61, TD61]. Infected [Eps62, EH63a, LW63, DK60]. Infection [DF62, NHK60, NMW60, NMWVF61, RGB60]. Infective [RD61]. Inflammation [MP61, MPS61]. Influence [But62, DS61b, TD61]. Influenza [Bay64]. Infolded [Tor63]. Infoldings [Bat61]. Inherited [DS62b]. Inhibition [Epe63, HS63a, PP64]. Inhibitors [Amo63]. Inhibitory [PP61]. Initiation [CC64a]. Injected [Str62, Str64a, Str64b]. Injection [BCM64, PL64, WN61]. Injections [TC64]. Injured [Buc62]. Injury [Buc62, HL62a, Rey63a]. Inner [MM61, OB64c]. Innervation [MBH63, Smi60, Smi63a]. Inorganic [CD61a, TG64]. Insect [Baw64, GR64, Loc61, Smi60, Smi61c, SL60]. Insoluble [Stä63]. Instrument [Dan61]. Insulin [BB60, LBDL63, OM62b]. Integrity [BH61, HHP60]. Intensity [EK62]. Interaction [Bay64, HLG63, HL62a, RM63, RMG64, Yam62]. Interactions [TKMT61]. Intercalated [GE60, dBdMP62]. Intercellular [Bie63, Ove63]. Interconnections [Kar60a]. Intercostal [ZB61a]. Interference [AC62, SSKD60]. Interferenzmikroskopische [SSKD60]. Internal [PP60]. Interphase [Amo62a, ATD62]. Interplay [Flo64]. Interstitial [CF61b]. Interzellulärer [Bie63]. Intestinal [Bon63, Bry63, HQ61, HT64, LC63b, MS64b, Mil64, Tri63, LA60]. Intestine [ALS63, MB62, SL62, Str63]. Intra [BCM64, FS61a]. Intra-Articular [BCM64]. Intracellualr [RR64]. Intracellular [Cos64, DCPN64, Dzi62, EGS61, ER62, Fel62a, FM62, FG64, Hay61, HS63a, NHK60, Riz61, Rot60b, SHSG61, TD63a, TT61, KWCL60, Rot60a]. Intracisternal [BK62, HS62]. Intracytoplasmic [KT61]. Intramitochondrial [FM64b, NN63a, NN63b, Pea64]. Intranuclear [Ric61, SSK62]. Intravenously [Str62, Str64a, Str64b]. Invadens [BS64a]. Invaginations [FaP64, SO62]. Inverse [HH63]. Invertebrate [Reb61]. Investigation [Str62, KG60, SSKD60]. Investigations [Ver61]. Iodine [SHSS62]. Iodoalbumin [SHSS62]. ionantha [Ehr60]. Ionogenic [O’B64a]. Ions [EL62, Fri64, GRL64]. Iproniazid [CZ63]. Iridescent [TW61]. Iron [BCM64, DS61b, Ric61, Yas62b]. Iron-Dextran [DS61b]. Irradiated [BL62]. Irradiation [BN60, HR62, Rah62, Ryt60]. Isolated [BS62, Ber62, BH63, BAS64, DHV61, EH61a, FKKH60, GTL61, HS63b, KPM61, KT63, MSP63a, MSP63b, Mal63, MS63, Rah62, RC62, RCA62, RAC62, SA64, STA64, SD64]. Isolation [CMRS62, CBBH63, CC64b, DOB63, DPRG61, FHI63, HMR60, Hin62, HM62, HM63b, LBDL63, MMR62, MMMH61, MM64, Nev60, O’B64b, PM62, RS64a, SWD62, Kan62b, Gri60]. Isopoda [TC60]. Isotopes [FG64].

Jejunal [Bro62, CS62]. Jejunum [MF62]. Jellyfish [Szo64]. Junction [Bar62, Csi63, LK64, WSL64]. Junctional [FP63, SF64]. Junctions [MP62, Reg61, ZB61a].

K12S [Ryt60]. Kb [MW63b]. Keratin [BT63, FR62, Nak64]. Keratinizing [FS61b]. Kidney [ALS63, BAS64, Mil60, Mil61a, NP60, NMW60, PF63b, RBF^{+}61, Str62, Str64a, Tho62, HA60b, LBB60]. Kidneys [RM64]. Kinetic [HG63, HL62b]. Kinetics [Per61, Owe63, OM63]. Kinetonucleus [Ste60]. Knife [Fah63a]. Knochenmarks [Sto62a]. Known [Gib63]. Korschikoff [HC63]. Kowalevskii [CC63a, CC63b]. Kv [SHSS62]. Kynurenine [Riz61, RR64, Riz64].

L [GA63, BKS61, DS61a, DF62, Dal63, DS61b, Fra62, NMRS60, RP64, WR64]. L-929 [Fra62]. Label [vT61]. Labeled [CG61, Dal63, HQ61, PL64, TS63a, TH61]. Labeling [vdD63]. Labelling [Pel64]. Laboratory [RG62]. Lacteals [CS62]. Lactic [FA64]. laevis [Pet60b]. Lamellae [Buc61, CW64, Kes63, Wis60a]. Lamellar [Her64b, Ito60, Kle60]. Lamellibranch [Gib61]. Laminated [TP64]. Lamination [RV62]. Lampbrush [EG63, HM63a]. Large [MCT61, Gri60]. Larva [BS64b]. Larvae [Cop64, Web64]. Larval [CD61b, RR63]. Lateral [FW62a, FW62b, Mil64]. Lathyrism [TH64]. Lathyritic [Kee60a]. Layer [BPA62]. Layered [MP62]. Layers [BPA64a, BPA64b]. Lead [Fel62c, Mil61b, PD60, Kar61, Rey63b, DZ60]. Leaflets [Sha64]. Leaves [EK62]. Lemm [HG60c]. Lengths [PH63]. Lens [RWG60]. Leptinotarsa [SKdBH63]. Lesion [BBR61, BBR60]. Lesions [WR64]. Leucine [NMRS60]. Leucyl [NMRS60]. Leukemia [Ack61, dHF60]. Leukemic [Ack61]. Leukocytes [Jac64, dFR63, Mit60]. Level [Epe63, GSW64, HL62a, RK63]. Leydig [Hay61]. Light [EK62, GTL61, KBB64, LC63a, LA64, Mun61a, TKMT61]. Light-Scattering [GTL61]. Like [EH63a, HG63, HS63b, vdD63, UV60]. Lilium [HS61a]. Line [FW62a, FW62b]. Lines [GG64, TG63]. Linked [EL62, MS64a]. Linne [Gré61]. Lipid [BKS61, LH62, MMR62, Sto62b, Tho62, Tro64, MMB61]. Lipid-Rich [MMR62, MMB61]. Lipid-Water [LH62, Sto62b]. Lipofer [Rob61b]. Lipomyces [Rob61b]. Lipoprotein [HLG63, Mil61a]. Lipoproteins [CS62]. Liquid [LH62, Sto62b]. Liquid-Crystalline [LH62, Sto62b]. Liver [ALS63, BZ62, BS62, Ber62, Bry63, CMRS62, DOB63, DCPN64, EMNB62, ESP62, EJ62, FHI63, GSA62, GRL64, HHC64, HH61a, HR62, LA64, LK61, LD63, MMR62, MPH61, MPH62, Nev60, O’B64a, PHH63, RBF^{+}61, Rey63a, Ric61, SSK62, SD64, Str64a, TJA63, Tro64, WL63, Yas62a, vdD63, DK60, EB60, MRC60, SH60]. Lividus [Lal62]. Living [CP60, RP60, Ros61, Gri60, HPWJ60]. Lizard [Gra60]. Lizards [EW60]. Localization [AC62, ALS63, BS64a, Bar62, BGSN60, Car61b, CF61a, ER62, EH63b, EN61, FA64, HHC64, Har64, HD60, HH61a, Kar64, LL61b, MB63b, Nak64, NSD61, Osi64, OVLG64, Riz61, RT63, SC63a, Tel61, TW61, Tho64b, TB62, WSB64, vIL64b, vIL64a, vT61, KWCL60, NMRS60]. Localized [RB62a, Yas62b]. Location [Hil64, WOM63b]. Locus [JE62]. Lomentaria [Bou62]. Longiflorum [HS61a]. Longitudinal [Pea61]. Loss [CP60]. Lota [FW62b]. Low [Kle60]. Lucifugus [RM64]. Lumbrical [Mer60]. Luminescent [Smi63a]. Lung [TP64, Kar60b]. Lutein [End62, EL64]. Lympho [Ack62]. Lympho-Epithelial [Ack62]. Lymphocyte [Bry63]. Lymphocytes [IC63, She62]. Lymphocytic [Rie62]. Lymphomatosis [DK60]. Lysis [SS62b]. Lysosomal [HT64]. Lysosome [Rah62]. Lysosome-Rich [Rah62]. Lysosomes [AP62, BM64a, EN61, KWB64, Str64a, Str64b, WBT64]. Lytic [MP64b, CC60a, CC60b].

M [KH63]. Macaca [EM61]. Macroglobulin [Sto62a]. macrogynus [BT60]. Macromolecular [FMOBG64]. Macromolecule [BG64]. Macromolecules [Car61b, CF61a]. Macronuclear [EKB62, PKC62]. Macronuclei [McD62]. Macronucleus [Klu62]. Macrophages [DBWK63, SC63b]. Macrosteles [SL60]. Macula [Flo64]. Maculosus [HK61b]. Magnetic [Nak64]. Maintain [Das62]. Makroglobulin [Sto62a]. Makroglobulin-Retention [Sto62a]. Male [dFR63]. Maleic [CG61]. Malignant [BH61]. Malleata [Yas62b]. Mammalian [BGM64, BB61, BT62a, BBP62, CBB^{+}63, Dal62, EGG62, GSA62, Hsu64, MZ60, RG64b, SS61b, Tay63, WA60, MMB61]. Mammals [Cam64]. mammary [SPB61, WD61]. Man [MR63b]. Manganese [MNP62]. Manipulating [Mun60]. Mapping [MCT61]. Maraglas [SKF63, Erl64]. Marinus [Cho63]. Marker [SM64]. Marrow [Ack60, Sto62a]. Masking [Gon62]. Mass [BS60]. Mast [HG63]. Material [Das62, Gra60, HG60c, LC63a, Lee62, SBdlTL61, Yas62a, DZ60]. Materials [MSP64]. Matrix [CD63, KSW63, Rei61, TG64, Wat63]. Matter [Tho64b]. Matteuccia [GA63]. Maturation [FVK61, Rei61, Rei63, RDM64]. Mature [BM64a, Tay64, TD63b]. Maturing [OS64]. Maturity [RG62]. Mauthner [Rob63a, RBS63]. Mays [JSB63]. Me [LA64]. Means [EN61]. Measles [RGB60]. Measurement [BS60, Hal60, ODH63]. Measurements [FS64, Too64]. Mechanism [All60, Cur64, RW64]. Mechanisms [Amo63]. Medium [HW61, HR62, KK62a, Tay62]. Mega [SN60, Ste60, Ste64]. Megakaryocytes [Gar64]. Megaterium [HD60, RLF62]. Meiosis [HS63a, HS63b, dFB62, Mon64]. Meiotic [Nic61]. Meissner [CR60]. Melanin [Dro60a]. Melanization [LM63]. Melanocytes [Dro60a, RNS63]. Melanogaster [DKG64, RR62, RR63, Sha63a, Sha63b]. Melanotic [RNS63]. Membrane [BNH62, Cha62, Csi63, FJ60, Fre62, KL64, LK63, MM61, Mil64, MP62, Nev60, Ros62c, SGW64, Sug61, Sug64, Tho64a, Tor64, TP64, Yam63a, CC60a, CC60b, PP60]. Membranes [Bat61, CC63a, CC63b, Ito60, Lan64, Lil61, MB62, Rob63a, Tor63, PPS60]. Mengo [DF62]. Mengovirus [Fra62]. Menschlichen [Sto62a]. Mercaptoethanol [SS62b]. meristem [Man60]. Mesothelium [Ham60b]. Messenger [HS63b]. Messenger-Like [HS63b]. Metabolic [BB60, Ber62, GJ60, LFB64a, MK60a, NM64, OFS^{+}63]. Metabolically [Nap63a]. Metabolism [DL63, VNL61, VL61, WRS61]. Metal [Sto61].

Metals [LFB64b]. Metamorphosis [Bon63, Web64]. Metaphase [CBBH63, Rob61a]. Metaphysis [Cam61]. Metastatic [GSW64]. Metchnikovii [GKH63]. Methacrylate [LC62]. Methionine [CP60]. Method [ABM63, AH63, BP61, BC61, Bul60, Fel62c, FG64, Gib63, Lou62, MHR61, Men61, O’B64b, Par61, SHSG61, SW64, KF60a, DZ60, Hol60, PSH60]. Methods [ALM64, Luf61, Mun61a, TS63b, WA61, Kar61]. Methoxy [NMRS60]. Methyldeoxycytidine [HS61a]. Mice [NHK60, Par62a, Ric61, WL63, SPB61, KWCL60, LD60]. Micro [McD62]. Micro- [McD62]. Microanalysis [Hal62a, Yas62a, Yas62b]. Microcentrifuge [Mal63]. Microchemical [EK61]. Microfibril [Col63]. microfibrillar [AP60]. Microfibrils [OD64]. Micrographs [Gon62, RNF60]. Microincineration [Tho64b]. Microinjection [Fel62a]. Micronuclear [PKC62]. Microorganisms [HG60a, HGCN60]. Microphoresis [Eds60b]. Microscope [AMFW60, Baw64, BPA64a, BPA64b, Bay64, BK63, BM64a, BS60, CSC61, Dal62, Dro60a, EES63, EH63a, Far64, FR62, Fin60b, Fin60c, GP60, GTWB63, Hag61, HB61, Har61b, JS63, Kes63, KT63, LC63a, LA64, Las61, Lav60, MB63a, Mat63, Mat60, MCT61, MF62, Mun60, NMW60, PD60, Par62b, PF63a, Reb61, RS61a, SB64, SJ64, SY60, SHSG61, SC63b, TKMT61, VR61, Wat63, WR64, Ryt60, Cla60, Her60, HA60b, Kee60b, Kee60a, LBB60, Mit60, MDvD60, VO60, dHF60]. Microscopic [ABM63, BL62, BP60c, BP62, BT62a, BBP62, Car61a, Eds60b, Hei62, KMF63, Lee62, MB61, MW63b, Nak64, NdTBB62, NMWVF61, Osi64, PT62, Pea64, PCDMP61, RBL62, ROD60, RD62, Sha63a, Sha63b, SMV64, WMG61, KG60, BBR60, Ess60, JP60, Kar60b, Odo60, ZM60]. Microscopical [DML61, Sto62b]. Microscopie [Dro60b]. Microscopy [Ack62, AC62, BNH62, BC61, BH62, Bel61, BHE61, CH64, CFG62, DS61a, DR64, Dan61, EM61, EH63b, EN61, FL62, FS62, FM64a, GK61, GY64, HC61, HQ61, HH61a, HH61b, HZ61, IC63, JM61, Kar64, Klu62, Lar62, LC62, Mal63, Mar61, Mun61a, PMRGG62, PDLP61, PBD^{+}61, Pep61, PF61, PFH61, Per64, Rah62, RTPM63, RG64a, RLF62, SBB63, SSKD60, San62, Sch64, Sed61a, Sed61b, Sed62, SS61a, Stä63, SW64, Sto61, TS63b, TW61, Wet61, Yas62a, Zal61, vT61, Erl64, Kar61, Rey63b, Blo60, DZ60, Ehr60, Fin60a, Fra60, GOE60, GPN60, Hal60, HG60a, Hol60, LdR60, MS60, Yas60, YTT60]. Microsomes [CMRS62, EJ62, LD63, TJA63, MRC60, SH60]. Microspectrophotometry [Men61]. Microsporocyte [NTM60]. Microtubulation [MM61]. Microtubule [LP63]. Microtubules [HN64, Sla63, dT64]. Microvilli [Bro62, MF62]. Microvillous [MS64b]. Microvillus [Mil64]. Mid [BT62b]. Mid-Body [BT62b]. Migration [WLD63]. Milk [WD61]. Mineral [Tho64b]. Mink [EES63]. Mirabile [LBB60]. Mise [Dro60b]. Mite [Aro61]. Mitochondria [BZ62, BSB63, BM64b, BAS64, DHV61, EB64, EH61b, FS61a, GTL61, Gra60, GRL64, LVN62, Luc63, Mal63, MB62, Mug64, Nad64, NP60, Pac63a, Pac63b, PCDMP61, RT62, SD64, SSNB63, SC64, Ste60, Sto63, WS61a, WS63, dW62, Bor60, PSH60, ZM60]. Mitochondrial [Cha62, Cop64, FMOBG64, KK62b, LFB64a, LFB64b, NSD61, NE62, OVLG63, Pea62, RFP63, WL63, WS61b, PPS60, PP60, PSH60]. Mitomycin [CS64b]. Mitosis [Amo62a, Amo62b, BG64, CC64a, DR64, Epe63, Har61b, Kon63, LC63a, Mon64, Rob61b, Rob63b, ROD60, RD62, Wes64]. Mitotic [Amo63, Buc61, Har62, Har64, Kan62a, MMMH61, RG64b, RG64a, SK61b, Ste61, WRS61, Kan62b, EKY^{+}60, MHB60]. Mitotique [Lal62]. Mm [NHK60]. Mode [Eps62, GKH63, GDLD63, San62]. Modifications [SKF63]. Modified [Mil61b]. Molecules [CSC61, JE62, SS61a, Sto61, Hal60]. Molitor [Smi60]. Mollusc [Bat61, Gib61, MB63b, Wat63]. Monkey [NMW60]. Monochromatic [HG60a]. Mononuclear [DBWK63, DB64a]. Monotetrazolium [PSH60]. Monoxide [Epe63]. Montana [Ech63]. Months [LC63b, LA60]. MorÁN [Bul60]. Morphogenesis [TY60]. Morphological [Cha61, CZ63, Flo64, LA64, Las61, LGC61, MSBN63, YFJ62, MRC60]. Morphology [DHV61, ZD62, HPWJ60, JP60]. Morphometric [WK64]. Mosaic [Sha64, WR64]. Mosquito [Cop64, RP64]. Moths [Tel61]. Motion [HB61]. Mould [MS60]. Mounting [BC61]. Mouse [BS62, Ber62, Bul60, CG63, FS61a, GG64, HQ61, KG64, dFB62, Mat63, Mil60, Mil61a, Mil64, Mon62, Mon64, PDLP61, PBD^{+}61, Par62b, PF63b, RBL62, RG62, San62, SL62, SC63b, Tho62, TG63, Tro64, Wat62, Wol64, ZB61a, WD61, dBdMP62, BP60a, BP60b, GJ60, UV60, dHF60]. Movement [All60, AC62, FG64]. Mucins [GB63]. Mucopeptide [WOM63b]. mucopolysaccharides [Ozz64]. Mucoprotein [Bay64]. Mucosa [HT64, IW63, OS64]. Mudpuppy [BGW63]. Mulatta [EM61]. Multinet [Gre60]. Multinuclearity [KMTH60]. Multinucleation [Cap60]. Multiplication [DF62, Fra62]. Multiplicity [MHB60]. Musca [Bie63]. Muscle [Aro62, Aro63b, Aro63a, BS64b, BMT63, CKB61, Cos64, DHV61, EL62, FA64, Fah63b, FR61, FaP64, Har61a, HKA64, Hil64, JP64, KH63, Kar64, KC62, KMTH60, Lav60, Mac63, Mau61, MA61, MBH63, MACAvdD61, NM64, PG63, PH63, PH62, PF63a, Rev62, Ros63b, Smi60, Smi61a, Smi61b, Smi61c, Smi63b, Sor62, SKdBH63, SS61b, Tha63, TB62, TH61, Ver61, dBdMP62, Cap60, Mer60]. Muscles [Aro61, Pea61, Reg61, Sha63a, Sha63b, Sko61, Web64, Mer60]. Musculaires [Lav60]. Musculature [Kar60c]. Mutant [RR62, Riz64]. Mycobacteria [KT61]. Mycobacterial [TMT^{+}61]. Mycobacteriophage [TKMT61]. Mycobacterium [TKMT61]. Mycota [MM63b]. Myelin [BBP62, CHG61, Fin60b, Fin60c, Pet60a, Pet60b, Pet61, Pet64a, RP61, TS64]. Myelinating [Wol64]. Myeloblastic [Ack61]. Myeloblastosis [NdTBB62]. Myocardial [NB63, SO62]. Myocardium [WS61b]. Myofibril [Gil62]. Myofibrils [Aro62, GR64, TB62, TH61]. Myofilaments [Pet63]. Myometrium [WN61]. Myoneural [Bar62, Csi63]. Myotis [RM64]. Mytilus [Gré61].

N [FL62]. N-Vinylpyrrolidone [FL62]. Naphthylamide [NMRS60]. Narcine [WMG61]. Nature [KH63, Wat60]. Necturus [BGW63, HK61b, Kes63]. Needle [Yas62a]. Negative [AH63, CR64, Par63, vIL64a]. Negatively [Nad64, Sto63]. Nematoda [NGD64]. Nematode [Bir61]. Nematode-Induced [Bir61]. Neohydei [HM63a]. Neoplastic [RNS63]. Nephron [HK61b, LMM61, Str64b]. Nerve [Bat61, DPRG61, DdI61, EH61a, Fin60b, Fin60c, Hes60, Mat60, Pan62, RP61, Tha63, Tho64a, TS64, TC60, MK60b]. Nerves [UV60]. Nervous [BBP62, GP63, NHK60, Pet60a, Pet60b, Pet64a]. Network [Hay61]. Neural [PMRGG62]. Neurofilaments [Pan62]. Neuroglia [Gra60]. Neurohypophysis [RL64]. Neuromuscular [Reg61, ZB61a]. Neuron [HL62b]. Neuron-Glia [HL62b]. Neuronal [Ros62c, Wol64]. Neurons [HH63, HE62, RW64]. Neurosecretory [Osi64]. Neurospora [Luc63, TT61, Zal61]. Newborn [MB62]. Newt [SJ64, TD63b]. Nexus [DB64b]. Nicotiana [WT62, WR64]. Nitella [Gre60]. Nitzschia [Dru63]. Noctiluca [Afz63]. Node [UV60]. Nodes [RBS63, UV60]. Nodules [Ack62]. Non [Amo62b, Amo63, O’B64a, Pel64]. Non-Dividing [Pel64]. Non-Identical [Amo63]. Non-Ionogenic [O’B64a]. Non-Respiratory [Amo62b]. Normal [Ack60, BH61, CF61b, Cos64, HKA64, MR63b, NM64, RNS63, See64, Kar60a, Kee60b]. Note [NTM60, Pet63, GPN60]. Notochord [EW62, SJ64]. Novikoff [KZ60]. nucléaires [Ryt60]. Nuclear [ALM64, EG63, Fel62b, Fel64, Gib62b, HS62, HG60c, HE62, KL64, KT63, MM61, Mer61, Mer62, MCC62, Per61, SBdlTL61, SM64, WML60, BH60a, BB64a, Cla60, Pal60]. Nuclear-Cytoplasmic [MCC62]. Nucleated [Dav61]. Nuclei [AB62, ALM64, ATD62, DS62a, FHI63, Gal62, Hin62, JM61, Jen64, KS60, LK63, MSP63a, MSP63b, O’B64b, RC62, RG62, See64, SWD62, TD63b, VR61, Wat62]. Nucleic [AK63, BS64a, Bas62, EB62, HZ61, WA61, WRS61, Yam62]. Nucleocytoplasmic [Fel62b, Fel64]. Nucleolar [Das62, LC63a, MK60a]. Nucleoli [BH63, EB62, EG63]. Nucleolus [KK64, Mac64, MK60b]. Nucleoproteins [LW63, SL64]. Nucleoside [EJ62, Per61, SC63a]. Nucleotide [DHV61, Hil64]. Nucleus [Afz63, HR63, JE64, MK60b]. Nursing [Las61]. Nutritive [Yas62b].

Oat [TT64]. Observation [BN60, Hei62, MW63a]. Observations [Aro63a, Boa64, EW60, End62, EL64, Eps62, EH63a, GSA62, HG60d, HA60a, Jen64, Las61, Law60a, Mac63, Nag62, NM64, Pea64, Pet64a, RGB60, RWG60, RFP63, Rob63b, RP60, RS61b, RBL62, SY60, SK61a, Sto62b, Sto63, Str64a, SC63b, SR61, VR61, Wat62, WML60, Wis60a, Yam63b, ZB61a, HG60a, JP60, VO60, ZM60]. Observed [Eps61, Per64, Sim64]. Obtaining [Gib63, Par61]. Occurrence [JSB63, Rob63a, SN60, Str64b, TD63a, TD63b]. Ochromonas [Gib62a]. Octopus [GY64]. Odonata [Smi61a]. Oeuf [Lal62]. Off [FS64]. Off-Peak [FS64]. Oligodendroglial [EH61a]. Oligodendrogliomas [RV62]. Onion [BG64]. Oocyte [And64, Par62b, RP64, RG62, BK62, Fra60].

Oocytes [AH64, BK63, Eds60a, EGS61, EG63, Kes63, KF60b, Reb61, Tel61, Wis60a, Odo60]. opaque [Rey63b]. Opisthobranch [Bat61]. Opossum [CF61b, RL64]. Optic [Mat60, UV60]. Optical [BL62, Sto62a]. Oral [Far64]. Orange [RM63, RMG64]. Orange-Cell [RM63, RMG64]. Ordered [Pea62]. Organ [CP60, FW62a, FW62b, Smi63a]. Organelle [TD61, Van61]. Organelles [Bou63, KT61]. Organic [CD63, TG64]. Organisation [Gra60]. Organisms [RD61]. Organites [BN60]. Organization [Csi63, EH61b, Smi61a, Smi63a, TG64]. Organizations [Smi61b]. Organized [Mac64]. Organs [KS64]. Orientation [FW62a, Gib63, JE62, TS63b]. Origin [BK63, Kes63, LVN62, War62a, War62b, CC60b]. Orthoptera [GR64]. Oscillation [SC64]. Osmiophilic [TP64]. Osmium [HLG63, Mun61a, PMRGG62, Rob61a, Ros63c, Too64, Tor64]. Osmium-Fixed [Mun61a, Ros63c]. Osmotic [LFB64a, TJA63]. Osmunda [JE62]. Osteoblasts [Cam61, Ros61]. Osteoclast [HB61]. Osteoclasts [GK61]. Osteogenesis [You62]. osteogenic [Owe63, OM63]. Other [Bou62, MZ60, Ols62, RBF^{+}61]. Oursin [Lal62]. Outgrowing [Ove63]. Ova [Odo60]. Ovar [Bie63]. Ovarian [Odo60]. Ovariectomized [WN61]. Ovary [Bie63, Kes62, Fra60]. Overlap [Aro63b]. Oxidase [LY60]. Oxidation [DHV61]. Oxidative [HK61b, NSD61, Pac63a, Pac63b, VNL61, VL61]. Oxide [Ric61]. Oxygen [Amo62a]. Oxyntic [Sed61a, Sed61b, Sed62, SF64].

P [HPWJ60]. Pack [Baw64]. Pads [MB61]. Paired [TMT^{+}61]. Palea [Dru63]. Paludosa [CG61]. Pancreas [Bul60, SP62, WLD63, SP60a, SP60b]. Pancreatic [Car61a, CP64, HF62a, HF62b, JM61, KG64, WHF62, Wes64, Cla60]. Paneth [BM64a, SL62]. Papillae [Cop64]. Paracentrotus [Lal62]. Paradoxa [HC63]. Paramecium [FKKH60, KP64]. Parasporal [Han61, Mon61]. Parathyroid [MR63b]. Parenchymal [Rey63a]. Parencyhma [O’B64a]. Parietal [EW60, Ito61, SF61, Law60a, VO60]. Parotid [Par62a]. Partial [KL64, LC63b, O’B64a, Tro64, Kan60, LA60]. Participation [Amo62b, FJ60]. Particles [AGZ61, BGM64, MMR62, ODH63, PT62, PF63b, vdD63, MMB61, AGC60]. Particular [Ack62, RNS63, RM64, Tor63]. Particulate [EL62]. Particulates [FKKH60, PSH60]. Parts [Eds60a]. Passage [CS62, PT62]. Pathologic [SGW64]. Pathways [Fel62b]. Pattern [Ger64, MPH62, Pea62, PHH63, PH64, Rob63a, Sat62]. Patterns [Nic61, OFS^{+}63]. Pea [Amo63, BH63, RC62]. Peak [FS64]. Pectin [AMFW60]. Penetration [BB61, SR61]. Pennsylvanica [Smi63a]. Pentatomidae [Nic61]. Pepsinogen [Lil61]. Perfusion [PMRGG62]. Perikaryon [Pan62]. Perineural [SB62]. Period [HE62, dFB62]. Periodic [FT62, Ste61]. Periosteal [TC64]. Peripheral [Tho64a, TS64]. Periplaneta [And64, Hes60]. Peritoneal [Rie62]. Permanganate [JM61, MZ60, Par61, Ros63c, Wet61, Law60b, ZM60]. Permanganate-Fixed [Ros63c]. Permeability [BH61, CD61a, SH60]. Peroxidase [Str62, Str64a, Str64b, Mit60]. Persistence [RG62]. pH [Kan62b, Kar61, MMB61, Rey63b, Tay62]. Phago [Str64b]. Phago-Lysosomes [Str64b]. Phagocytic [DBWK63]. Phagocytized [BGM64]. Phagocytizing [OFS^{+}63]. Phagocytosis [Kar60b]. Phagosomes [Str64a, Str64b]. Phagotrophy [Sea61]. Pharmacological [CZ63]. Pharyngeal [Mac63]. Phase [BN60, Buc62, RP60]. Phases [LH62, Sto62b]. Phialidium [Szo64]. Phlorizin [BAS64]. Phosphatase [Beh63, EN61, Hal61, HH61a, KC63, Osi64, OVLG64, RB62a, Sea61, Str64a, WSB64, MDvD60]. Phosphatase-Containing [Beh63]. Phosphate [GRL64, HK61a, KK62b]. Phosphate-Induced [KK62b]. Phosphorylation [HS61b, Pac63a, Pac63b]. Photodestruction [KB64]. Photometric [TJA63]. Photoreceptive [Uch62]. Photoreceptor [JE62, WG61]. Photoreceptors [LdR60, MR60]. Photosynthetic [BD61]. Photuris [Smi63a]. Physiology [MBH63, PR61]. Phytohaemagglutinin [IC63]. Picture [HB61]. Pig [AH64, DL63, MBH63, SP62, SP60a, SP60b]. Pigment [BGW63, DG62, Kes62, KPM61, MS63]. Pigments [KB64]. Pineal [DdI61, KS64]. Pinocytic [MB63a]. Pinocytosis [BP60c, BP62, EGG62, SN60, CAH60]. Pipiens [KS64, MF63, PM64, War62a, War62b]. Pits [Bou62]. Pituitary [FT62, HM62, HM63b, PM62, San62, HMR60, KWCL60]. Placenta [SS62a]. Placental [Gal62]. Planarian [Mac63]. Plant [BM64b, BV63, Bou63, LP63, MM64, Nad64, TS63a, KG60]. Plants [Jen64]. Plasma [PDLP61, PBD^{+}61, Ros62c, Sto62a, Sug61, Sug64]. Plasma-Cell [PDLP61, PBD^{+}61]. Plasmazellen [Sto62a]. Plasmodium [HPWJ60]. Plasms [Sch64]. Plastic [Dan61]. Plastid [KBB64, PP64]. Plastids [GG62, JSB63]. Plate [WM63]. Platelet [RPM63]. Plates [Afz61, BMT63, MPH61]. Plexus [PT62]. Plurivesicular [DdI61]. Pneumoniae [TJO64]. Polarization [ABM63, Flo64]. Polioencephalitis [NHK60]. Pollen [Ehr60]. Polyampholytes [MS64a]. Polyester [LC62]. Polyester-Methacrylate [LC62]. Polymerase [PBK62]. Polymorphonuclear [DBWK63]. polymyxa [Hol60]. Polyoma [HA60a, HA60b]. Polyploid [KS60]. Polyribosome [Per64]. Polyribosomes [RDM64]. Population [LC63b, Owe63, OM63]. Pore [BB64b, Loc61]. Pores [FA63]. Positive [SBdlTL61, vIL64b]. Possible [Fel64, MR63a, TD63b, Uch62]. Postfixation [Par61]. Postsynaptic [Csi63]. Postulate [WM63]. Potassium [JM61, Par61, Law60b, ZM60]. Potato [SKdBH63]. Potential [LK63]. Precipitate [Fel62c]. Precipitate-Free [Fel62c]. Precipitin [MPH62]. Preferential [HZ61]. Preliminary [HG60a]. Preparation [BH62, Blo60, Bor60, Hum63, Mal63, RTPM63, WS64]. Preparations [KT63, WQSW60, NR60]. Prepared [Bul60, SS61a]. Preparing [Hol60]. Presence [Dir61, HZB63]. Preservation [CHG61, HHP60]. Pressure [Hei62]. Prevention [KK62b]. Principal [MF62]. Prior [Dan61, FG64, dFB62]. Prisms [Gré61]. Problem [Boo64, CHG61]. Procambarus [Mos61a, Mos61b]. Procedure [Mil61b]. Process [Kar60b]. Processes [DdI61, PT62, RV62, Tha63]. Produced [Hir60]. Producing [Sip63]. Product [Hir60]. Production [HE62, Lee62, Sko61, Ozz64]. Products [Hal62a, Hal62b, TS63a]. Projection [Lee62]. Proliferation [TC64, You62, LA60]. Prolonged [She62]. Pronuclei [SK61a]. Properties [Ber62, DOB63, HM63b, MMR62, SS61a, WS64, Bor60, PPS60]. Proplastids [KPM61, KB64]. Prostate [BP60a, BP60b]. Protein [ALM64, BH63, BOS62, Car61a, CP64, DKG64, HS63a, HS63b, HE62, KP64, Kon63, MP64b, Mon61, RP64, RCA62, RAC62, See64, Str62, Tay63, TW61, War62a, War62b, WHF62, WRS61, ZD62, vdD63, MMB61, WD61, FP60, Hal60, Kan60]. Protein-Secreting [ZD62]. Proteins [GC63, Tel61, WLD63, KF60a]. Proteus [RP62a, RP62b, RD61, WP64b, CAH60]. Protoplasmic [BB64b]. Protozoa [RS64c]. Proximal [Mil60, Tho62]. Pseudoplatanus [WN64]. Puffs [EB62]. Pulmonary [WK64]. pulmonate [BH60a, BH60b]. Pump [Cop64]. punctulata [Kan60]. Purification [ALM64]. Purkinje [Her63]. Puromycin [SL64]. Purposes [HH61b, HHP60]. purpuratus [BM60]. Pyriformis [EKB62, EB64, McD62, SP64]. Pyrimidines [KB60]. Pyrophosphatase [MB63b, Osi64].

Quantitative [HLG63, Lou62, TG63]. Quintuple [MP62]. Quintuple-Layered [MP62].

Rabbit [BCM64, DBWK63, DB64a, DHV61, EL62, HS60, HS62, Lar62, NB63, PT62, She62, Tor63]. Rabies [Mat63]. Radial [Pet61]. Radiation [Par62b, SNE60]. Radioactive [SHSS62]. Radioautographs [WLD63]. Radioautography [Car61a, FG64, PL64, Sim64, vT61]. Radiosensitive [DR64]. Rana [KS64, MF63, PM64, Sed61b, Sed62, War62a, War62b, Sed61a]. Ranarum [Rob63b]. Ranvier [UV60]. Rapid [BC61, O’B64b]. Rat [ALS63, BZ62, BGSN60, Beh63, BM64a, BP63, BT62b, Cam61, CMRS62, CZ63, CP60, DOB63, DdI61, DG62, DS62b, EMNB62, ESP62, EJ62, Far64, FVK61, FHI63, FM64b, GK61, GRL64, GP63, Her64a, HH61a, HT64, HM63b, KSW63, KH63, Kar64, LA64, LMM61, LK61, LC63b, MF62, Mil64, MPH61, Mug64, Nag63, Nev60, NP60, NEO63, OM62b, Osi64, PM62, PHH63, RBF^{+}61, Rah62, Ros62b, RBL62, SD61, SS62a, Sip63, SWD62, Str62, SR61, TJA63, WN61, Wal63, Wat62, WSB64, vdD63, Cla60, EB60, Fra60, HHP60, Kee60b, Kee60a, Law60a, LBB60, LA60, Mer60, MRC60, Odo60, SH60, Wat60, ZM60]. Rate [CP60, Epe63]. Rates [Amo62a]. Ratios [Ive64]. Rats [CS64a, MB62, Sim64, SD64, You62]. Ray [ES63, Fin60b, Fin60c, Hal62a, HR62, Yas62a, Yas62b]. rayons [Ryt60]. Rays [MP64a, TD61]. Reabsorption [MB61, Str64b]. Reaction [FT62, Hal62a, Hal62b, Hir60, Mit60]. Reactions [MPH62]. Reaggregation [DL63]. Reappearance [BM64b]. Reared [NGD64]. Receptor [FW62a]. Receptors [FW62a]. Red [Bou62, KBB64, PF63a]. Reduced [DHV61].

Reductive [vIL64b, vIL64a]. Reeve [Yas62b]. Reference [Ack62, BK63, CD61b, CC61b, FW62a, Flo64, RNS63, Rob61a, RLF62, RM64, San62, Tor63, BH60b]. Regenerating [Bry63, Tro64]. Regeneration [SJ64, Tho64a]. Region [Aro63a, CC61b, PL64]. Regions [Aro63b, TD63b]. Regressing [Web64]. Regular [Gra60]. Regulating [Riz64]. Regulation [HS61b, HS63a]. Rejoining [SNE60]. Related [Loc61, PR61, KB60]. Relation [BH61, BT60, CC64a, DCPN64, ES63, Ste61, Str64b, Tay63, Tho64a, BM60, CC60a]. Relations [See64]. Relationship [Gib61, GE60, HL62b, Nic61, NP60, Ols62, Ros62c, Str64a, TG64, PP60]. Relationships [ESP62, Gib62b, MCC62, Wat63, Fra60]. Relative [Ive64, Ros63a, MK60b]. Release [Eps62]. Remarkable [Van61]. Removing [MHR61]. Remyelination [BBR61]. Renal [FP62, GSW64, MP64b, Yam60]. Renewing [CG63]. Repeating [FMOBG64]. Replicas [Hag61]. Replication [Gal61, Pai61, PHH63, PH64]. Resection [LC63b, LA60]. Reserpine [CZ63]. Resin [Luf61, MHR61, Stä63, Law60b]. Resins [FL62, Fin60a]. Resolution [BH62, CvT62, Car62]. Resolved [ES63]. Respiratory [Amo62b, Amo63, EH61b]. Response [Fri64]. Responses [CS64a, Flo64, Tay62]. Resting [CKB61, HR62]. Restitution [HF62b]. Results [NSD61]. Rete [LBB60]. Retention [Sto62a]. Reticular [Smi61b, Man60]. Reticulum [BK62, BM64a, EMNB62, ESP62, Fah63b, Fah64, FR61, FA63, Her64b, Ito61, Lav60, PY60, Por61, Reg61, Rev62, RP60, SJ64, PM60]. Retina [Las61, LW62, SC63a, TY60]. Retinal [DS62b, OB64c, WG61, LdR60, MR60, TY60]. Retinulae [Gol62]. Reutilization [Bry63, Rie62]. Revealed [Flo64, HH61a, San62, WLD63, Yas60, YTT60]. Reversal [Cha62]. Reversible [SKdBH63]. Rhizoctonia [BB64b]. Rhodomicrobium [BD61]. Rhodospirillum [Boa64, CBK63]. Ribonuclease [RBF^{+}61]. Ribonucleases [Rot60b, Rot60a]. Ribonucleic [Bie63, DCPN64, Eds60a, Eds60b, EGS61, EG63, EH61a, HS63b, HE62, Ive64, MK60a, Mon64, RC62]. Ribonucleoprotein [vdD63, KZ60]. Ribonukleinsäure [Bie63]. Ribosomes [CMRS62, MB62]. Rich [MMR62, PF63b, Rah62, MMB61]. RNA [DML61, JSB63, KP64, Kon63, Per61, See64]. Roach [And64]. Rodent [NHK60]. Rods [OB64c, TY60]. Roentgens [HR62]. Role [CC63a, CC63b, KK64, Las61, LFB64b]. Rooster [CM64]. Root [Amo63, CG61, SNE60]. Root-Tip [CG61]. Rootlets [Ols62]. Roots [Amo62a, Amo62b, BG64]. Rosy [RR62]. Rotifer [LL61b]. Rotifers [LL61a, Lan64]. Rough [DOB63, MB62]. Route [Tel61]. RPL [DK60]. RPL-12 [DK60]. Rubrum [Boa64, CBK63]. Rudimentary [Sor62]. Rumen [RS64c]. Rustica [WT62].

S [CP60]. S. [GH61]. S3 [Pai61]. Saccharated [Ric61]. Saccharomyces [VNL61, VL61]. Saccoglossus [CC63a, CC63b]. Saintpaulia [Ehr60]. Salamander [Her60, Mit60]. Saline [MF62]. Salivary [EB62, JS63, MK60a].

Salts [NSD61, Sto61]. Sand [SK61a, KB60]. Sap [EG63]. Sarcolemma [MA61]. Sarcolemmal [FaP64, SO62]. Sarcoma [CW64, Har64, Rie62]. Sarcomere [Aro61, Gil62]. Sarcoplasmic [Fah63b, Fah64, FR61, FA63, Por61, Reg61, Rev62]. Sarcosomes [Smi63b]. Sarcotubular [MACAvdD61]. Satellite [Mau61]. Satellites [Szo64]. Saturniid [Tel61]. Say [SKdBH63]. Scales [LM63]. Scanning [Hal62a, Yas62a, Yas62b]. Scattering [GTL61]. Schedule [BH60b]. Schiff [FT62, Hir60]. Schiff-Formaldehyde [Hir60]. Schwann [GP63, Tho64a]. Sciatic [UV60]. Scintillans [Afz63]. Sea [Amo63, BPA62, BPA64a, BPA64b, CKP61, Dir61, Epe63, Har61b, Har62, Hin62, PCK64, Sak60a, Sak60b, BM60, GPN60, Kan60]. Secondary [HN64, WN64]. Secreting [ZD62, KWCL60]. Secretion [GG64, LBDL63, San62, Wis63, Wis60b]. Secretory [DdI61, PM62]. Sectioning [Dan61, FL62, RLF62]. Sections [BC61, Boo64, Bul60, Car61a, CS62, FT62, Fel62c, Gib63, HQ61, Mar61, Mil61b, Mun61a, PD60, Par61, Per64, RNF60, Sch64, Wat63, DZ60, Hol60, Law60b]. Seen [AMFW60, LBB60]. Segment [OB64c]. Segments [BL62, LMM61, Str64b]. Segregation [FP60]. Selected [Blo60]. Selection [RG64a]. Seminal [GJ60]. Sensitive [GG62]. Sensory [FW62a, Flo64]. Separation [MHB60]. Septal [BB64b]. Serial [BC61, Gib63]. Serological [Mon61]. Serum [HW61, KK62a]. Serum-Free [KK62a]. Set [Mun60]. Several [PD60, ZD62]. Sex [ATD62, KS60]. Shadow [ODH63]. Shadow-Casting [ODH63]. Shape [NTM60, Pac63a, Pac63b, TS63b]. Shaped [Voe64]. Sheath [CHG61]. Sheathed [GKH63]. Sheaths [BBP62, Hes60, Pet60a, Pet60b, Pet61, Pet64a, RP61]. Sheep [SO62]. Shell [But62, WW61]. Shells [Wat63]. Short [HE62]. Shortening [Aro63b]. Shown [PL64]. Sialic [GB63]. side [KF60a]. side-chain [KF60a]. Siemens [Mun60]. Significance [AC62, TD63b, CC60a, JB60]. Silica [PCDMP61]. Silver [Mar61]. Simple [BC61, BH62, Kar61, MHR61, Par61]. Simplified [DZ60]. Simultaneous [PD60]. Single [DF62, Gib63, SSKD60]. Sister [MP64a]. Site [Car61a, GL64, Har61a, LCOVG64]. Sites [LF64, vIL64b, vIL64a]. Situ [MS63]. Sixth [Uch62]. Size [Aro61, Aro63a, Col63, LFB64a, Pac63a, Pac63b, TS63b]. Skeletal [FA64, KC62, Mau61, MA61, MACAvdD61, PG63, PF63a, Ton61, Cap60]. Slices [MPH61]. Slides [Lee62]. Slime [MS60]. Slow [PH62]. Small [ALS63, Far64, FS61b, MB62, Tri63]. Smooth [CMRS62, DOB63, MBH63, Sor62, Tha63]. snail [BH60a, BH60b]. Sodium [Wet61]. Solani [BB64b]. Solidissima [RS64a]. Solids [MK60b]. Solitary [MS60]. Soluble [FG64, MR63a, MSP64, Stä63]. Solutions [Kee61, KT63, PMRGG62, Sug64, Too64]. Somatic [Cam64]. Some [EKY^{+}60, Eps61, Hal62b, Har62, LK63, Mer62, NM64, Rob61a, RS61b, Sto62b, Sto63, Too64, Van61, WML60, Yam63b, Ham60a, MR60]. Sonic [SC64]. Sound [Sko61]. Source [BKS61]. Sources [DML61]. Sp [Smi61a]. Space [LW62]. Special [BK63, CD61b, CC61b, FW62a, Flo64, Rob61a, RLF62, San62, TC60, BH60b]. Specialization [You62, SL60]. Specializations [OS64]. Species [KC63, PC63, WOM63b]. Specific [BOS62, JP64, Pep61, PF61, PFH61, SS61a, WLD63]. Specimen [Dan61, Mun60]. Specimens [FM64a, Hei62]. Spectrophotometry [FT62]. Speculations [ZD62]. Sperm [SR61, Tay64, CC60a, CC60b]. Spermatid [Nag62, Nag63, NTM60]. Spermatids [Kay64]. Spermatocytes [Mon62]. Spermatogenesis [TD61, Yas60, YTT60, BH60b]. Spermatogenic [Ito60]. Spermatogonia [Mon62]. Spermatozoa [BZ64, Gib63, SBdlTL61]. Spermatozoan [SHT61]. Spermatozoon [CC61c, CC61a, CC61b]. Spermiogenesis [Baw64, Mos61a, Mos61b, BB64a]. Spherical [VR61]. Spheroplasts [Boa64]. Spherules [BHE61]. Spider [Eds60a]. Spinach [MNP62]. Spinal [BTAD60, BBR61, Ros63c, RW64, BBR60]. Spindle [Buc61, GC63, Voe64]. Spindle-Associated [GC63]. Spindle-Shaped [Voe64]. Spindles [RS64a, Mer60]. Spisula [RS64a]. Spleen [ER62, SC63b]. Splitting [EH63b]. Spontaneous [See64]. Sporangia [Hol60]. Spore [FJ60, OM62a, WOM63b, YFJ62]. Spores [JE62, RLF62, Tho64b, WOM63a, HG60a, Hol60]. Spread [Par63]. Squash [WQSW60]. stål [SL60]. Stability [Sat62, SHSS62]. Stage [KSW63]. Stages [Beh63, CKP61, KBB64, PCK64, Rei63, dW62, HPWJ60]. Stain [AK63, CH64, GB63, WA61, Rey63b, Law60b]. Stained [AMFW60, Nad64, Sto61, Sto63]. Staining [AH63, Bul60, Fel62c, HH61a, HH61b, HZ61, Mil61b, Mun61a, NSD61, PD60, Par61, Par63, SW64, Str64a, Kar61, DZ60, HHP60]. Stains [SS61a, ZB61b]. Stalked [PCB64]. Staphylococcus [Sug61, Sug64]. Starch [But62, WT62]. Starch-Free [WT62]. Starfish [EGS61]. Starved [CC64a]. Statistical [NTM60]. Stenosis [WS61b]. Stimulated [IC63, PF63b]. Stimulation [Her60]. Stolon [Ove63]. Stomach [SF61]. Stomatitis [MW63b]. Storage [CP64, SBdlTL61, SSK62]. Strain [DS61a, Dal63, DS61b, NHK60, NdTBB62, Ros62a]. Stranded [HC61]. Stratification [Bou63]. Streaming [BB64b]. Streptomyces [GH60, GH61, HG60b, HG60d]. Stretch [CKB61]. Striated [CKB61, Fah63b, Hil64, Kar60c, PH63, PH62, Smi61b, Ver61]. Striatum [Mug64]. Stroma [HL62a]. Strongylocentrotus [BM60]. Structural [AGZ61, ALS63, Bar62, Bon63, Har62, KB64, KK62a, Nag63, NB63, PH62, SD64, TB62, WA60, AGC60]. Structure [Afz61, BTAD60, Bel61, BS64b, BS62, BD61, Boa64, Bou62, BP63, BT62b, But62, Cam61, CD61b, Cho63, CF61b, CR64, CBK63, CBPK64, CM64, CC61b, DF62, Dav61, DS62a, DB64b, DG62, Dro60a, Dru63, DS61c, EW60, EW62, EW64, Ech63, EK62, EMNB62, End62, EL64, Eps61, ESP62, ES63, Far64, FMOBG64, FR62, Fin60b, Fin60c, Flo64, Fre62, FM64b, Gib61, Gil62, GH60, GH61, GKH63, Gol62, GY64, GRL64, Gré61, Hay61, HR63, Her63, Her64a, Hes60, HG60c, HA60a, IW63, KG64, Kay64, KS64, Kes62, KPM61, KBB64, Lal62, LL61a, LP63, LBB60, LH62, Mac63, MA61, MJ63, MNP62, Mer61, MBH63, MF62, MM63a, MM63b, Nad64, Nag62, NM64, OM62a, OS64, PM64, Pea61, PR61, Pet60a, Pet60b, Pet64a, PP61, Pet63, PC63]. Structure [PCB64, RNS63, Reg61, RFP63, RK63, RS61b, RP61, Ros62b, Ros63b, RL64, RS64c, SF61, SHT61, SO62, Smi61c, Smi63b, SC64, Szo64, TMT^{+}61, Tho62, TJO64, Tor63, TG64, Tro64, TC60, Uch62, VR61, Ver61, WOM63a, Yam63b, ZB61a, AP60, CR60, HG60b, BT60, BP60a, BP60b, Ehr60, EB60, GG60b, GPN60, Ham60b, Ham60a, HGCN60, Kan60, Law60a, Mer60, MR60, TY60, VO60, dL60]. Structures [FW62b, HZB63, HZ61, KT61, Loc61, Lou62, Mac64, MZ60, Nac64, Ols62, OB64c, WG61, UV60]. Struthiopteris [GA63]. Stubenfliege [Bie63]. Studied [Bel61, GP60, Ton61]. Studies [AH64, ALM64, BTAD60, BK63, Buc62, Cha61, CSC61, CBBH63, CHG61, Dro60a, EN62, EH61b, FM62, Fin60b, Fin60c, FKKH60, GG60a, GDLD63, GTL61, HC63, HB61, HGCN60, HH61b, JS63, JSB63, Kar63, KP62, KPDM62, Kay62a, KWB64, Kes63, LVN62, Lit64, LK64, LW63, LGC61, MSP63a, MSP63b, MP61, MPS61, MSBN63, Mat63, MBM62, MF62, Mon61, MW63b, PG63, PM60, PY60, Reb61, RS61a, ROD60, Rot60a, Rot60b, RD62, RCA62, RAC62, Sak60a, Sak60b, San62, Sat63, Sch64, SA64, STA64, Sha63a, Sha63b, SR60, SK62, TMT^{+}61, TKMT61, TG63, Tri63, VNL61, VL61, WK63, WW61, Wat63, WBT64, WSL64, YFJ62, Cla60, Odo60, dHF60]. Study [BZ62, BZ64, Bas62, Baw64, BM64a, BB61, BP60c, BP62, BT62a, BBR61, CC64b, Dal62, DBWK63, DML61, DB64b, EES63, EJ62, Far64, FR62, FW62a, GSW64, Gil62, GTWB63, Har61b, HL62b, HM63b, KT63, LC63a, LA64, Lav60, Mat60, Mon62, MB61, Nak64, NdTBB62, NMW60, NMWVF61, PT62, PM64, Par62b, PF63a, PM62, PCDMP61, SJ64, SP62, Tro64, WK64, WR64, BBR60, DK60, Ess60, Her60, HA60b, Kar60b, Kee60b, Kee60a, MRC60, PSH60, SP60a, SP60b]. Studying [FG64]. Styela [Kes62]. Sub [MPH61]. Subcellular [KC63]. Submicroscopic [Csi63, LD60]. Suboxydans [CR64]. Subsequent [Bou63, RG64a, SJ64]. Substitution [Bul60, Reb61]. Substrate [Hum63, NMRS60]. Subsurface [Ros62c]. Subtilis [Van61]. Subunit [Rob63a]. Succinic [LY60, RT63, WS61a, WS63]. Sucrose [EK62, KT63, SH60]. Sulfate [CD61a, Dzi62, TD63a, Yam62]. Sulfation [GL64, LCOVG64]. Sulfhydryl [BH61, Sak60a, Sak60b]. Supporting [Hay61]. Suppression [BG64]. Surf [RS64a]. Surface [BPA62, BPA64a, BPA64b, EGG62, EH63a, EH63b, Hag61, O’B64a, OS64, Pea62, TP64]. Surfaced [MB62]. Surfaces [Ros63a]. Survey [Wal63]. Survival [She62]. Suspensions [BP61, Mal63, O’B64b]. Sv40 [GTWB63]. Sweat [EM61, MB61, Mun61b]. Swelling [Har61a, KK62b, RT62]. Swim [Sko61]. Swim-Bladder [Sko61]. Swimming [Afz61]. Swimming-Plates [Afz61]. Symbiont [HC63]. Sympathetic [Yam63b]. Synapses [PP61, Rob63a, RBS63, Wol64, dL60]. Synaptic [CM64, DPRG61, FW62b, GY64]. Synchronously [EKB62]. Syneresis [AC62]. Synovial [BCM64, BNH62]. Synthese [Bie63]. Syntheses [HS63a]. Synthesis [ALM64, ATD62, BG64, Bas62, BH63, CG63, Cam64, CP64, Dzi62, ES63, Ger64, GC63, HS61b, HS63b, KP64, KMTH60, Kon63, dFB62, Luc61, McD62, MF63, Mon62, Mon64, MD63, Nak64, ODH62, Pet64b, PL64, PKC62, RP62a, RP62b, RR64, Riz64, SK61b, SP64, Tay63, WLD63, Wes64, WP64b, BB64a, MMB61, Bie63, BM60, DK60, EKY^{+}60]. Synthetic [Das62, SGW64]. Syrian [RNS63]. System [BBP62, FaP64, GP63, HG63, Kle60, MACAvdD61, NHK60, NB63, Pet60a, Pet60b, Pet64a, RR63, SGW64, SO62, Ton61]. Systems [Ito60, KC63, LH62, MPH62, Sto62b, ZD62].

T [FaP64]. Tadpole [BS64b]. Tail [SHT61, Web64]. Tarsonemid [Aro61]. Tdr [Ton61]. Teased [Aro63b]. Technical [SKF63]. Technique [BPA62, CFG62, Gon62, ODH63, PD60, Stä63]. Techniques [CHG61]. Teleost [FW62b]. Temperature [Kle60]. Tenebrio [Smi60]. Tension [Amo62a]. Terminal [MP62, RB62a, KF60a]. Terminations [Hes60]. Test [Kes62]. Testes [MZ60, ZM60]. Testicular [CF61b, Yas62b]. Testis [Mon62, Mon64]. Tetrahymena [EKB62, EB64, McD62, Sea61, SP64]. Tetranitro [RT63]. Tetranitro-Blue [RT63]. Tetrazolium [NSD61, RT63]. Tetroxide [HLG63, PMRGG62, Rob61a, Too64, Tor64]. Tetroxide-Fixed [Tor64]. Tetroxide-Lipoprotein [HLG63]. Their [ALM64, BM64a, CS62, DF62, Dal62, Das62, EH61a, GE60, HM63b, RP61, Ros62c, Szo64, TG63, CC60b, Hes60, TY60, HPWJ60, MHB60]. Theoretical [Rob61a]. there [GJ60]. Thermobia [Baw64]. Thermoelectric [FM64a]. Thermophilic [RS61b]. Theta [San62]. Thiamine [MB63b, Osi64]. Thickness [BS60, WK64, Yam63a]. Thicknesses [Mil64]. Thin [Boo64, Bul60, Car61a, Fel62c, HQ61, Mil61b, PD60, RNF60, RLF62, Sch64, Wat63, DZ60]. Thinly [Par63]. ThioglycÉRol [Lal62]. Thiosorbitol [Lal62]. Third [FP62]. Thoracic [She62]. Three [OFS^{+}63, PC63, JP60]. Thuringiensis [Mon61]. Thymidine [HQ61, MP64a, Mon62, NGD64, SK61a, TS63a, Ton61]. Thymus [ALM64, KT63, Rah62]. Thyroid [FVK61, FM64b, Wis63, WSB64, Her60, Wis60b]. Thyrotropin [KWCL60]. Tibia [Cam61]. Tigrina [Mac63]. Time [PHH63, PH64, SNE60, MHB60]. Timing [PKC62, SK61b]. Tingible [SC63b]. Tip [CG61]. Tips [Amo63, SNE60]. Tipula [TW61]. Tissue [BCM64, BB60, CS64a, CS62, Eds60b, Har64, HKA64, MHR61, MCC62, Mun61a, RNF60, RP60, Ros62a, RBL62, Wil64, Owe63, OM63, SSKD60, DK60, Law60b, NR60]. Tissues [PMRGG62, Par61, Pel64, TS63a, HHP60]. Tiv [TW61]. Toad [Cho63, GE60, Hil64, Las61, LW62, NM64, PR61, Ros63c, RW64]. Toadfish [Sko61]. Tobacco [Sha64, WR64]. Todaro [GA63]. Toe [MB61]. Tomato [Sha64]. touch [CR60]. Tpn [WS63]. Tpn-Dependent [WS63]. Trachea [CP60]. Tradescantia [CG61, NTM60]. Transfilter [KG64]. Transform [KC63]. Transformations [Pac63a, Pac63b, Ryt60]. Transient [HS61b, dW62]. Transition [DKG64, BH60b]. Transplantable [KWCL60]. Transplantation [SM64]. Transport [Bie63, Str64b]. Transverse [NB63, SO62]. Treated [BAS64, Jac64, Ric61, Sug64, SPB61, GOE60]. Treatment [MF62, SJ64]. Tree [RB62a]. Tri [EJ62]. Tri- [EJ62]. Trillium [HS63b]. Trim [Dan61]. Triphosphatase [ALS63, EH63a, Har64, SC63a]. Triphosphatase-Like [EH63a]. Triphosphatases [PG63]. Triphosphate [EL62, EH63b]. Triphosphate-Linked [EL62]. Tritiated [Mon62, NGD64, RC62, SK61a, TS63a, Ton61]. Tritium [WQSW60, vT61]. Triturus [TD63b, EG63]. Tropic [JE62]. Tropocollagen [MR63a]. Trypan [RBF^{+}61]. Trypanosoma [SN60, Ste60, Ste64]. Trypanosome [SN60]. Trypsin [HG63, LD63, SS62b]. Trypsin-Like [HG63]. Tubal [Odo60]. Tubular [NB63, SO62, TS64]. Tubule [Mil60, Tho62]. Tubules [Mil61a]. Tumor [Bas62, NdTBB62, RCA62, RAC62, Bor60, EKY^{+}60]. Tumors [Fre62, PDLP61, PBD^{+}61, RNS63, SPB61, KWCL60]. Tumour [See64]. Turnover [CD63]. Twitch [PH62]. Two [Bas62, BPA62, BS63, Dal62, EN62, LC63b, Men61, NSD61, VR61, WMG61, LA60]. Two-Emulsion [Bas62]. Two-Layer [BPA62]. Two-Wavelength [Men61]. Type [FP62, HM63a, TC60]. Types [OFS^{+}63, WMG61, Wol64, ZD62]. Tyrosine [Sim64].

Ultra [BN60]. Ultra-Violette [BN60]. Ultracentrifuged [BTAD60]. Ultramicrotome [MMWFW61]. Ultrasonic [SJ64]. Ultrastructural [BB60, BBR61, HC63, Lan63, LB61, Nak64, SD61, SGW64, Tho64b, Web64]. Ultrastructure [Bir61, BAS64, CKB61, CHG61, GE60, JE64, KC62, LK61, MS63, Mun61b, OM62b, Pea63, Rei61, Rei63, RP62c, RG64b, RBS63, SS61b, Tha63]. Ultrathin [Mar61]. Ultraviolet [BL62, GG62, JM61, HG60a]. Ultraviolet-Irradiated [BL62]. ultraviolets [Ryt60]. Undergoing [HN64]. Undifferentiated [AB62]. Unfertilised [BPA64a, BPA64b]. Unfertilized [Odo60]. Unit [FMOBG64, Rob63a, Yam63a]. Units [Eds60b]. Untersuchungen [SSKD60]. Unusual [Eps61, Hay61, Her64b, Par62a, Ros62a]. Upon [Pea62, SBdlTL61]. Uptake [BCM64, CBB^{+}63, Dal63, RW64, RP64, RCA62, RAC62, SK61a, Str62, CAH60]. Urchin [Amo63, BPA62, BPA64a, BPA64b, CKP61, Dir61, Epe63, Har61b, Har62, Hin62, PCK64, Sak60a, Sak60b, BM60, GPN60, Kan60]. Uridine [KF60b, Luc61]. Uridine-H [KF60b]. Urinary [Bay64, Cho63, PR61]. Use [AK63, CH64, FM62, Hum63, Mal63, ODH63, Pep61, PF61, PFH61, RT63, TS63b, WS63, WA61, vT61, Rey63b, Law60b]. Using [CFG62, FM64a, Mon62, Par61]. Uterine [Lar62]. Utilizing [Nak64]. Utriculi [Flo64].

V [TY60]. Vaccinia [DS61a, Dal63]. Values [Gar64]. Vanadium [CH64]. Vannielii [BD61]. Var [HS61a]. variabilis [Sha60]. Variation [DS62a]. Variations [Aro63a, Lan64]. Varied [AB62]. Various [Eds60a, FP63, Fri64, Wal63]. Varying [MP64a]. Vas [MBH63]. Ventricular [KH63, NM64]. Vesicle [GJ60]. Vesicles [CMRS62, DOB63, DPRG61, MB63a, Tor63]. Vesicular [MW63b]. Vesiculated [Tha63]. Vessels [Kar60c, PT62, Ham60a]. Vibrio [GKH63]. Vicia [CG61, LC63a, SNE60, WRS61]. Vinylpyrrolidone [FL62]. Violaceoruber [GH61]. Violette [BN60]. Viral [DF62, Dal63, MJ63]. Virginia [MR63b]. Virginiana [RL64]. viridifasciata [BB64a]. virilis [Ito60]. Virus [AH63, Bay64, DS61a, Dal63, Eps62, EH63a, GTWB63, HA60a, LW63, Mat63, MW63b, NHK60, NdTBB62, NMW60, NMWVF61, Par63, Sha64, TW61, WR64, DK60, HA60b, RGB60]. Viruses [DF62, Dal62]. Visual [BGW63, LD60]. Visualization [LF64]. Visualized [CD63]. Vitamin [GDLD63]. Vitelline [CC60b]. Vitellogenesis [And64]. Vitro [Buc62, FT62, Fri64, HKA64, Hsu64, KF60b, LM63, MSBN63, RG64a, SBdlTL61, SK61b, Str63, Wes64, Cap60, Ozz64, SH60]. Vivo [LK61, Rie62]. Volume [BOS62, PP60]. vulgare [TC60]. Vulgaris [FW62b].

Wachstum [SSKD60]. Walker [Buc61]. Wall [HN64, Mil64, TMT^{+}61, WN64, KG60, Gre60]. Wallerian [TS64, dW62]. walls [AP60, Ehr60]. Water [BG64, BH61, FL62, FG64, LH62, Stä63, Sto62b]. Water-Insoluble [Stä63]. Water-Soluble [FG64, Stä63]. Wavelength [Men61]. Well [DL63]. Wendl [Ehr60]. Which [KC63]. White [FA64, Nap63b, NP60, PF63a, Str62]. Whole [O’B64b, Rah62]. Whole-Body [Rah62]. Width [ODH63]. Winter [RM64]. Winter-Hibernating [RM64]. Within [Fel64, FS61a, Smi61b, Yam60]. Wound [RB61, RB62b, RB62c, RB64].

X [Rah62, Yas62b, ES63, Fin60b, Fin60c, Hal62a, HR62, MP64a, Ryt60, TD61, Yas62a]. X-Irradiation [Rah62]. X-Ray [Yas62b, ES63, Fin60b, Fin60c, Hal62a, HR62, Yas62a]. X-Rays [MP64a, TD61]. Xenopus [Pet60b, Web64]. Xylem [WN64]. Xylinum [ODH62].

Yeast [LVN62, MM63a, Rob61b]. Yolk [BK63, Bel61, Kar63, RP64, WK63, War62a, War62b]. Young [Cam61, You62].

Z [KC62]. Zea [JSB63]. Zellen [SSKD60]. Zonation [SD61]. Zone [Car61a, RM64].

References

Agrell:1962:CEV

[AB62] Ivar Agrell and Hans-Åke Bergqvist. Cytochemical evidence for varied DNA complexes in the nuclei of undifferentiated cells. Journal of Cell Biology, 15(3):604–??, December 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/15/3/604.

Allen:1963:NMP

[ABM63] Robert D. Allen, James Brault, and Robert D. Moore. A new method of polarization microscopic analysis. Journal of Cell Biology, 18(2):223–??, August 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/18/2/223.

Allen:1962:SAM

[AC62] Robert D. Allen and Ronald R. Cowden. Syneresis in ameboid movement: Its localization by interference microscopy and its significance. Journal of Cell Biology, 12(1):185–??, January 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/12/1/
185.

Ackerman:1960:HDD

[Ack60] G. Adolph Ackerman. Histochemical demonstration of dehydrogenase activity in the cells of normal human blood and bone marrow. Journal of Cell Biology, 8(1):61–??, September 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/8/1/61.

Ackerman:1961:HCC

[Ack61] G. Adolph Ackerman. Histochemistry of the centrioles and centrosomes of the leukemic cells from human myeloblastic leukemia. Journal of Cell Biology, 11(3):717–??, December 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/11/3/
717.

Ackerman:1962:EMB

[Ack62] G. Adolph Ackerman. Electron microscopy of the bursa of fabricius of the embryonic chick with particular reference to the lympho-epithelial nodules. Journal of Cell Biology, 13 (1):127–??, April 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/13/1/127.

Afzelius:1961:FSC

[Afz61] Björn A. Afzelius. The fine structure of the cilia from ctenophore swimming-plates. Journal of Cell Biology, 9(2):

383–??, February 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/9/2/383.

Afzelius:1963:NNS

[Afz63] Björn A. Afzelius. The nucleus of noctiluca scintillans. Journal of Cell Biology, 19(1):229–??, October 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/19/1/229.

Arcos:1960:FSA

[AGC60] Joseph C. Arcos, George W. Griffith, and Richard W. Cunningham. Fine Structural Alterations in Cell Particles During Chemical Carcinogenesis. Journal of Cell Biology, 7(1): 49–??, February 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/7/1/49.

Arcos:1961:FSA

[AGZ61] Joseph C. Arcos, Hans H. Gosch, and David Zickafoose. Fine structural alterations in cell particles during chemical carcinogenesis. Journal of Cell Biology, 10(1):23–??, May 1961. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/10/1/23.

Almeida:1963:NSM

[AH63] June D. Almeida and Allan F. Howatson. A negative staining method for cell-associated virus. Journal of Cell Biology, 16 (3):616–??, March 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/16/3/616.

Adams:1964:SGP

[AH64] Eleanor C. Adams and Arthur T. Hertig. Studies on guinea pig oocytes. Journal of Cell Biology, 21(3):397–??, June 1964. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/21/3/397.

Albersheim:1963:UBE

[AK63] Peter Albersheim and Ursula Killias. The use of bismuth as an electron stain for nucleic acids. Journal of Cell Biology, 17(1):93–??, April 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/17/1/93.

Allen:1960:CAC

[All60] Robert D. Allen. The consistency of ameba cytoplasm and its bearing on the mechanism of ameboid movement. Journal of Cell Biology, 8(2):379–??, October 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/8/2/379.

Allfrey:1964:MPT

[ALM64] V. G. Allfrey, V. C. Littau, and A. E. Mirsky. Methods for the purification of thymus nuclei and their application to studies of nuclear protein synthesis. Journal of Cell Biology, 21(2):213–??, May 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
21/2/213.

Ashworth:1963:FSL

[ALS63] C. T. Ashworth, F. J. Luibel, and S. C. Stewart. The fine structural localization of adenosine triphosphatase in the small intestine, kidney, and liver of the rat. Journal of Cell Biology, 17(1):1–??, April 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/17/1/1.

Albersheim:1960:SPS

[AMFW60] Peter Albersheim, K. Mühlethaler, and A. Frey-Wyssling. Stained pectin as seen in the electron microscope. Journal of Cell Biology, 8(2):501–??, October 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/8/2/501.

Amoore:1962:OTR

[Amo62a] J. E. Amoore. Oxygen tension and the rates of mitosis and interphase in roots. Journal of Cell Biology, 13(3):365–??, June 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/13/3/
365.

Amoore:1962:PNR

[Amo62b] J. E. Amoore. Participation of a non-respiratory ferrous complex during mitosis in roots. Journal of Cell Biology, 13(3):373– ??, June 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
13/3/373.

Amoore:1963:NIM

[Amo63] J. E. Amoore. Non-identical mechanisms of mitotic arrest by respiratory inhibitors in pea root tips and sea urchin eggs. Journal of Cell Biology, 18(3):555–??, September 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/18/3/555.

Anderson:1964:ODV

[And64] Everett Anderson. Oocyte differentiation and vitellogenesis in the roach Periplaneta Americana. Journal of Cell Biology, 20 (1):131–??, January 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/20/1/131.

Aronson:1960:MSC

[AP60] Jerome M. Aronson and R. D. Preston. The microfibrillar structure of the cell walls of the filamentous fungus, Allomyces. Journal of Cell Biology, 8(1):247–??, September 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/8/1/247.

Ashford:1962:CCH

[AP62] Thomas P. Ashford and Keith R. Porter. Cytoplasmic components in hepatic cell lysosomes. Journal of Cell Biology, 12 (1):198–??, January 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/12/1/198.

Aronson:1961:SSD

[Aro61] John Aronson. Sarcomere size in developing muscles of a tarsonemid mite. Journal of Cell Biology, 11(1):147–??, October 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/11/1/
147.

Aronson:1962:EMI

[Aro62] John Aronson. The elongation of myofibrils from the indirect flight muscle of Drosophila. Journal of Cell Biology, 13 (1):33–??, April 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/13/1/33.

Aronson:1963:OVS

[Aro63a] John Aronson. Observations on the variations in size of the a region of arthropod muscle. Journal of Cell Biology, 19(2): 359–??, November 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/19/2/359.

Aronson:1963:OBC

[Aro63b] John Aronson. Overlap of the birefringent component of adjacent a regions during the induced shortening of fibrils teased from Drosophila muscle. Journal of Cell Biology, 19(1):107–??, October 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
19/1/107.

Atkins:1962:ADS

[ATD62] Leonard Atkins, Priscilla D. Taft, and Kusum P. Dalal. Asynchronous DNA synthesis of sex chromatin in human interphase nuclei. Journal of Cell Biology, 15(2):390–??, November 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/15/2/390.

Barrnett:1962:FSL

[Bar62] Russell J. Barrnett. The fine structural localization of acetylcholinesterase at the myoneural junction. Journal of Cell Biology, 12(2):247–??, February 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/12/2/247.

Baserga:1962:SNA

[Bas62] Renato Baserga. A study of nucleic acid synthesis in ascites tumor cells by two-emulsion autoradiography. Journal of Cell Biology, 12(3):633–??, March 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/12/3/633.

Burgos:1964:UIK

[BAS64] Mario H. Burgos, Agustin Aoki, and Fabio L. Sacerdote. Ultrastructure of isolated kidney mitochondria treated with phlorizin and ATP. Journal of Cell Biology, 23(2):207–??, November 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/23/2/
207.

Batham:1961:INF

[Bat61] E. J. Batham. Infoldings of nerve fibre membranes in the opisthobranch mollusc Aplysia Californica. Journal of Cell Biology, 9(2):490–??, February 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/9/2/490.

Bawa:1964:EMS

[Baw64] S. R. Bawa. Electron microscope study of spermiogenesis in a fire-brat insect, thermobia domestica Pack. Journal of Cell Biology, 23(3):431–??, December 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/23/3/431.

Bayer:1964:EME

[Bay64] M. E. Bayer. An electron microscope examination of urinary mucoprotein and its interaction with influenza virus. Journal of Cell Biology, 21(2):265–??, May 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/21/2/265.

Barrnett:1960:MUC

[BB60] Russell J. Barrnett and Eric G. Ball. Metabolic and ultrastructural changes induced in adipose tissue by insulin. Journal of Cell Biology, 8(1):83–??, September 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/8/1/83.

Borenfreund:1961:SPM

[BB61] Ellen Borenfreund and Aaron Bendich. A study of the penetration of mammalian cells by deoxyribonucleic acids. Journal of Cell Biology, 9(1):81–??, January 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/9/1/81.

Bloch:1964:ECS

[BB64a] David P. Bloch and Sheilah D. Brack. Evidence for the cytoplasmic synthesis of nuclear histone during spermiogenesis in the grasshopper Chortophaga viridifasciata (de Geer). Journal of Cell Biology, 22(2):327–??, August 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/22/2/327.

Bracker:1964:FSP

[BB64b] C. E. Bracker and E. E. Butler. Function of the septal pore apparatus in rhizoctonia solani during protoplasmic streaming. Journal of Cell Biology, 21(1):152–??, April 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/21/1/152.

Bunge:1962:EMD

[BBP62] Mary Bartlett Bunge, Richard P. Bunge, and George D. Pappas. Electron microscopic demonstration of connections between glia and myelin sheaths in the developing mammalian central nervous system. Journal of Cell Biology, 12(2):448–??, February 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/12/2/
448.

Bunge:1960:EMS

[BBR60] Richard P. Bunge, Mary Bartlett Bunge, and Hans Ris. Electron Microscopic Study of Demyelination in an Experimentally Induced Lesion in Adult Cat Spinal Cord. Journal of Cell Biology, 7(4):685–??, July 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/7/4/685.

Bunge:1961:USR

[BBR61] Mary Bartlett Bunge, Richard P. Bunge, and Hans Ris. Ultrastructural study of remyelination in an experimental lesion in adult cat spinal cord. Journal of Cell Biology, 10(1):67–??, May 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/10/1/
67.

Barnes:1961:SRM

[BC61] Barbara G. Barnes and T. C. Chambers. A simple and rapid method for mounting serial sections for electron microscopy. Journal of Cell Biology, 9(3):724–??, March 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/9/3/724.

Ball:1964:UIR

[BCM64] J. Ball, J. A. Chapman, and K. D. Muirden. The uptake of iron in rabbit synovial tissue following intra-articular injection of iron dextran. Journal of Cell Biology, 22(2):351–??, August 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/22/2/
351.

Boatman:1961:FSP

[BD61] Edwin S. Boatman and Howard C. Douglas. Fine structure of the photosynthetic bacterium rhodomicrobium vannielii. Journal of Cell Biology, 11(2):469–??, November 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/11/2/469.

Behnke:1963:DAP

[Beh63] O. Behnke. Demonstration of acid phosphatase-containing granules and cytoplasmic bodies in the epithelium of foetal rat duodenum during certain stages of differentiation. Journal of Cell Biology, 18(2):251–??, August 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/18/2/251.

Bellairs:1961:SYH

[Bel61] Ruth Bellairs. The structure of the yolk of the hen’s egg as studied by electron microscopy. Journal of Cell Biology, 11 (1):207–??, October 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/11/1/207.

Berry:1962:MPC

[Ber62] M. N. Berry. Metabolic properties of cells isolated from adult mouse liver. Journal of Cell Biology, 15(1):1–??, October 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/15/1/1.

Bal:1964:SMM

[BG64] Arya K. Bal and Paul R. Gross. Suppression of mitosis and macromolecule synthesis in onion roots by heavy water. Journal of Cell Biology, 23(1):188–??, October 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/23/1/188.

Bensch:1964:FDC

[BGM64] Klaus Bensch, Gerald Gordon, and Leonard Miller. The fate of DNA-containing particles phagocytized by mammalian cells. Journal of Cell Biology, 21(1):105–??, April 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/21/1/105.

Becker:1960:LEA

[BGSN60] Norwin H. Becker, Sidney Goldfischer, Woo-Yung Shin, and Alex B. Novikoff. The localization of enzyme activities in the rat brain. Journal of Cell Biology, 8(3):649–??, December 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/8/3/649.

Brown:1963:VCV

[BGW63] Paul K. Brown, I. R. Gibbons, and George Wald. The visual cells and visual pigment of the mudpuppy, necturus. Journal of Cell Biology, 19(1):79–??, October 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/19/1/79.

Bloch:1960:CNH

[BH60a] David P. Bloch and Howard Y. C. Hew. Changes in nuclear histones during fertilization, and early embryonic development in the pulmonate snail, Helix aspersa. Journal of Cell Biology, 8(1):69–??, September 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/8/1/69.

Bloch:1960:SSP

[BH60b] David P. Bloch and Howard Y. C. Hew. Schedule of Spermatogenesis in the Pulmonate Snail Helix aspersa, with Special Reference to Histone Transition. Journal of Cell Biology, 7(3):515– ??, June 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/7/
3/515.

Belkin:1961:RBW

[BH61] Morris Belkin and Walter G. Hardy. Relation between water permeability and integrity of sulfhydryl groups in malignant and normal cells. Journal of Cell Biology, 9(4):733–??, April 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/9/4/733.

Beer:1962:SPG

[BH62] Michael Beer and Peter J. Highton. A simple preparation of graphite-coated grids for high resolution electron microscopy. Journal of Cell Biology, 14(3):499–??, September 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/14/3/499.

Birnstiel:1963:PSI

[BH63] Max L. Birnstiel and Beal B. Hyde. Protein synthesis by isolated pea nucleoli. Journal of Cell Biology, 18(1):41–??, July 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/18/1/
41.

Breslau:1961:EMC

[BHE61] Abraham M. Breslau, Thomas J. Hensley, and John O. Erickson. Electron microscopy of cultured spherules of coccidioides immitis. Journal of Cell Biology, 9(3):627–??, March 1961. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/9/3/627.

Bier:1963:SIT

[Bie63] Karlheinz Bier. Synthese, Interzellulärer Transport, Und Abbau Von Ribonukleinsäure Im Ovar Der Stubenfliege Musca Domestica. (German) [Synthesis, intercellular transport, and breakdown of ribonucleic acid in the ovary of the housefly Musca Domestica]. Journal of Cell Biology, 16(2):436–??, February 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/16/2/
436.

Bird:1961:UHN

[Bir61] Alan F. Bird. The ultrastructure and histochemistry of a nematode-induced giant cell. Journal of Cell Biology, 11(3): 701–??, December 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/11/3/701.

Beams:1962:IGE

[BK62] H. W. Beams and Richard Kessel. Intracisternal granules of the endoplasmic reticulum in the crayfish oöcyte. Journal of Cell Biology, 13(1):158–??, April 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/13/1/158.

Beams:1963:EMS

[BK63] H. W. Beams and R. G. Kessel. Electron microscope studies on developing crayfish oocytes with special reference to the origin of yolk. Journal of Cell Biology, 18(3):621–??, September 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/18/3/
621.

Bensch:1961:SLA

[BKS61] Klaus G. Bensch, Donald W. King, and Edward L. Socolow. The source of lipid accumulation in L cells. Journal of Cell Biology, 9(1):135–??, January 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/9/1/135.

Bloom:1962:OEM

[BL62] William Bloom and Robert J. Leider. Optical and electron microscopic changes in ultraviolet-irradiated chromosome segments. Journal of Cell Biology, 13(2):269–??, May 1962. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/13/2/269.

Bloom:1960:PSC

[Blo60] William Bloom. Preparation of a Selected Cell for Electron Microscopy. Journal of Cell Biology, 7(1):191–??, February 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/7/1/191.

Bucher:1960:DAS

[BM60] Nancy L. R. Bucher and Daniel Mazia. Deoxyribonucleic Acid Synthesis in Relation to Duplication of Centers in Dividing Eggs of the Sea Urchin, Strongylocentrotus purpuratus. Journal of Cell Biology, 7(4):651–??, July 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/7/4/651.

Behnke:1964:EMS

[BM64a] O. Behnke and H. Moe. An electron microscope study of mature and differentiating paneth cells in the rat, especially of their endoplasmic reticulum and lysosomes. Journal of Cell Biology, 22(3):633–??, September 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/22/3/633.

Bell:1964:DRM

[BM64b] Peter R. Bell and Kurt Mühlethaler. The degeneration and reappearance of mitochondria in the egg cells of a plant. Journal of Cell Biology, 20(2):235–??, February 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/20/2/235.

Brzin:1963:CDE

[BMT63] Miro Brzin and Živa Majcen-Tkačev. Cholinesterase in denervated end plates and muscle fibres. Journal of Cell Biology, 19(2):349–??, November 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/19/2/349.

Bessis:1960:IUV

[BN60] Marcel Bessis and Georges Nomarski. Irradiation ultra-violette des organites cellulaires avec observation continue en contraste De phase. Journal of Cell Biology, 8(3):777–??, December 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/8/3/777.

Barland:1962:EMH

[BNH62] Peter Barland, Alex B. Novikoff, and David Hamerman. Electron microscopy of the human synovial membrane. Journal of Cell Biology, 14(2):207–??, August 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/14/2/207.

Boatman:1964:OFS

[Boa64] E. S. Boatman. Observations on the fine structure of spheroplasts of rhodospirillum rubrum. Journal of Cell Biology, 20 (2):297–??, February 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/20/2/297.

Bonneville:1963:FSC

[Bon63] Mary A. Bonneville. Fine structural changes in the intestinal epithelium of the bullfrog during metamorphosis. Journal of Cell Biology, 18(3):579–??, September 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/18/3/579.

Boothroyd:1964:PDT

[Boo64] B. Boothroyd. The problem of demineralisation in thin sections of fully calcified bone. Journal of Cell Biology, 20(1):165–??, January 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
20/1/165.

Borst:1960:PPM

[Bor60] P. Borst. Preparation and Properties of Mitochondria from Ehrlich Ascites Tumor Cells. Journal of Cell Biology, 7(2): 381–??, April 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/7/2/381.

Brill:1962:SVD

[BOS62] A. S. Brill, J. S. Olin, and B. M. Siegel. The specific volume of dry protein. Journal of Cell Biology, 13(2):249–??, May 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/13/2/249.

Bouck:1962:CDP

[Bou62] G. Benjamin Bouck. Chromatophore development, pits, and other fine structure in the red alga, lomentaria baileyana (harv.) farlow. Journal of Cell Biology, 12(3):553–??, March 1962. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/12/3/553.

Bouck:1963:SSB

[Bou63] G. Benjamin Bouck. Stratification and subsequent behavior of plant cell organelles. Journal of Cell Biology, 18(2):441–??, August 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
18/2/441.

Brandes:1960:FSEa

[BP60a] David Brandes and Adolfo Portela. The Fine Structure of the Epithelial Cells of the Mouse Prostate. Journal of Cell Biology, 7(3):505–??, June 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/7/3/505.

Brandes:1960:FSEb

[BP60b] David Brandes and Adolfo Portela. The Fine Structure of the Epithelial Cells of the Mouse Prostate. Journal of Cell Biology, 7(3):511–??, June 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/7/3/511.

Brandt:1960:EMS

[BP60c] Philip W. Brandt and George D. Pappas. An electron microscopic study of pinocytosis in ameba. Journal of Cell Biology, 8(3):675–??, December 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/8/3/675.

Baker:1961:NEM

[BP61] R. F. Baker and H. E. Pearson. New embedding method for cell suspensions. Journal of Cell Biology, 9(1):217–??, January 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/9/1/217.

Brandt:1962:EMS

[BP62] Philip W. Brandt and George D. Pappas. An electron microscopic study of pinocytosis in ameba. Journal of Cell Biology, 15(1):55–??, October 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/15/1/55.

Brightman:1963:FSE

[BP63] M. W. Brightman and S. L. Palay. The fine structure of ependyma in the brain of the rat. Journal of Cell Biology, 19 (2):415–??, November 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/19/2/415.

Baxandall:1962:TLT

[BPA62] Jane Baxandall, P. Perlmann, and B. A. Afzelius. A two-layer technique for detecting surface antigens in the sea urchin egg with ferritin-conjugated antibody. Journal of Cell Biology, 14 (1):144–??, July 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/14/1/144.

Baxandall:1964:IEMa

[BPA64a] Jane Baxandall, P. Perlmann, and B. A. Afzelius. Immunoelectron microscope analysis of the surface layers of the unfertilised sea urchin egg. Journal of Cell Biology, 23(3):609–??, December 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
23/3/609.

Baxandall:1964:IEMb

[BPA64b] Jane Baxandall, P. Perlmann, and B. A. Afzelius. Immunoelectron microscope analysis of the surface layers of the unfertilised sea urchin egg. Journal of Cell Biology, 23(3):629–??, December 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
23/3/629.

Byers:1963:CAa

[BPG63a] Thomas J. Byers, Dorothy B. Platt, and Lester Goldstein. The cytonucleoproteins of amebae. Journal of Cell Biology, 19(3): 453–??, December 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/19/3/453.

Byers:1963:CAb

[BPG63b] Thomas J. Byers, Dorothy B. Platt, and Lester Goldstein. The cytonucleoproteins of amebae. Journal of Cell Biology, 19(3): 467–??, December 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/19/3/467.

Brown:1962:MHJ

[Bro62] Arnold L. Brown. Microvilli of the human jejunal epithelial cell. Journal of Cell Biology, 12(3):623–??, March 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/12/3/623.

Bryant:1963:RLD

[Bry63] B. J. Bryant. Reutilization of lymphocyte DNA by cells of intestinal crypts and regenerating liver. Journal of Cell Biology, 18(3):515–??, September 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/18/3/515.

Burge:1960:MMT

[BS60] R. E. Burge and N. R. Silvester. The measurement of mass, thickness, and density in the electron microscope. Journal of Cell Biology, 8(1):1–??, September 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/8/1/1.

Berry:1962:FSC

[BS62] M. N. Berry and F. O. Simpson. Fine structure of cells isolated from adult mouse liver. Journal of Cell Biology, 15(1):9–??, October 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
15/1/9.

Brown:1963:CET

[BS63] A. D. Brown and C. D. Shorey. The cell envelopes of two extremely halophilic bacteria. Journal of Cell Biology, 18(3): 681–??, September 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/18/3/681.

Barker:1964:LCN

[BS64a] Douglas Cameron Barker and George Svihla. Localization of cytoplasmic nucleic acid during growth and encystment of entamoeba invadens. Journal of Cell Biology, 20(3):389–??, March 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/20/3/
389.

Berrill:1964:FSC

[BS64b] N. J. Berrill and H. Sheldon. The fine structure of the connections between muscle cells in ascidian tadpole larva. Journal of Cell Biology, 23(3):664–??, December 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/23/3/664.

Beattie:1963:BM

[BSB63] Diana S. Beattie, Howard R. Sloan, and R. E. Basford. Brain mitochondria. Journal of Cell Biology, 19(2):309–??, November 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/19/2/
309.

Blondel:1960:RBB

[BT60] Benigna Blondel and Gilbert Turian. Relation Between Basophilia and Fine Structure of Cytoplasm in the Fungus Allomyces macrogynus Em. Journal of Cell Biology, 7(1):127– ??, February 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/7/1/127.

Buck:1962:EMS

[BT62a] Robert C. Buck and James M. Tisdale. An electron microscopic study of the development of the cleavage furrow in mammalian cells. Journal of Cell Biology, 13(1):117–??, April 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/13/1/117.

Buck:1962:FSM

[BT62b] Robert C. Buck and James M. Tisdale. The fine structure of the mid-body of the rat erythroblast. Journal of Cell Biology, 13(1):109–??, April 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/13/1/109.

Bell:1963:DFK

[BT63] Eugene Bell and Y. T. Thathachari. Development of feather keratin during embryogenesis of the chick. Journal of Cell Biology, 16(2):215–??, February 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/16/2/215.

Beams:1960:SFS

[BTAD60] H. W. Beams, T. N. Tahmisian, Everett Anderson, and Rosemarie Devine. Studies on the fine structure of ultracentrifuged spinal ganglion cells. Journal of Cell Biology, 8(3):793–??, December 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/8/
3/793.

Buck:1961:LSM

[Buc61] Robert C. Buck. Lamellae in the spindle of mitotic cells of Walker 256 carcinoma. Journal of Cell Biology, 11(1):227–??, October 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
11/1/227.

Buckley:1962:CIV

[Buc62] I. K. Buckley. Cellular injury in vitro: Phase contrast studies on injured cytoplasm. Journal of Cell Biology, 14(3):401– ??, September 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/14/3/401.

Bullivant:1960:STS

[Bul60] Stanley Bullivant. The staining of thin sections of mouse pancreas prepared by the FernÁNdez–MorÁN helium ii freezesubstitution method. Journal of Cell Biology, 8(3):639–??, December 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/8/
3/639.

Buttrose:1962:IES

[But62] M. S. Buttrose. The influence of environment on the Shell structure of starch granules. Journal of Cell Biology, 14(2):159–??, August 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
14/2/159.

Bonneville:1963:NCC

[BV63] Mary A. Bonneville and Bruce R. Voeller. A new cytoplasmic component of plant cells. Journal of Cell Biology, 18(3): 703–??, September 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/18/3/703.

Bahr:1962:SMR

[BZ62] Gunter F. Bahr and Elmar Zeitler. Study of mitochondria in rat liver. Journal of Cell Biology, 15(3):489–??, December 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/15/3/489.

Bahr:1964:SBS

[BZ64] Gunter F. Bahr and Elmar Zeitler. Study of bull spermatozoa. Journal of Cell Biology, 21(2):175–??, May 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/21/2/175.

Chapman-Andresen:1960:UFA

[CAH60] Cicily Chapman-Andresen and Howard Holtzer. The uptake of fluorescent albumin by pinocytosis in Amoeba proteus. Journal of Cell Biology, 8(1):288–??, September 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/8/1/288.

Cameron:1961:FSO

[Cam61] D. A. Cameron. The fine structure of osteoblasts in the metaphysis of the tibia of the Young rat. Journal of Cell Biology, 9 (3):583–??, March 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/9/3/583.

Cameron:1964:DDS

[Cam64] I. L. Cameron. Is the duration of DNA synthesis in somatic cells of mammals and birds a constant? Journal of Cell Biology, 20(1):185–??, January 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/20/1/185.

Capers:1960:MSM

[Cap60] Charles R. Capers. Multinucleation of Skeletal Muscle in vitro. Journal of Cell Biology, 7(3):559–??, June 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/7/3/559.

Caro:1961:EMR

[Car61a] Lucien G. Caro. Electron microscopic radioautography of thin sections: the Golgi zone as a site of protein concentration in pancreatic acinar cells. Journal of Cell Biology, 10(1):37–??, May 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
10/1/37.

Caro:1961:LMEa

[Car61b] Lucien G. Caro. Localization of macromolecules in Escherichia coli. Journal of Cell Biology, 9(3):539–??, March 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/9/3/539.

Caro:1962:HRAb

[Car62] Lucien G. Caro. High-resolution autoradiogaphy. Journal of Cell Biology, 15(2):189–??, November 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/15/2/189.

Chorazy:1963:UMC

[CBB^{+}63] M. Choraży, A. Bendich, E. Borenfreund, O. L. Ittensohn, and D. J. Hutchison. Uptake of mammalian chromosomes by mammalian cells. Journal of Cell Biology, 19(1):71–??, October 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/19/1/71.

Chorazy:1963:SIM

[CBBH63] M. Choraży, A. Bendich, E. Borenfreund, and D. J. Hutchison. Studies on the isolation of metaphase chromosomes. Journal of Cell Biology, 19(1):59–??, October 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/19/1/59.

Cohen-Bazire:1963:FSR

[CBK63] Germaine Cohen-Bazire and Riyo Kunisawa. The fine structure of rhodospirillum rubrum. Journal of Cell Biology, 16(2): 401–??, February 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/16/2/401.

Cohen-Bazire:1964:FSG

[CBPK64] Germaine Cohen-Bazire, Norbert Pfennig, and Riyo Kunisawa. The fine structure of green bacteria. Journal of Cell Biology, 22(1):207–??, July 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/22/1/207.

Colwin:1960:EML

[CC60a] Arthur L. Colwin and Laura Hunter Colwin. Egg Membrane Lytic Activity of Sperm Extract and its Significance in Relation to Sperm Entry in Hydroides hexagonus (Annelida). Journal of Cell Biology, 7(2):321–??, April 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/7/2/321.

Colwin:1960:FSE

[CC60b] Laura Hunter Colwin and Arthur L. Colwin. Formation of Sperm Entry Holes in the Vitelline Membrane of Hydroides hexagonus (Annelida) and Evidence of their Lytic Origin. Journal of Cell Biology, 7(2):315–??, April 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/7/2/315.

Colwin:1961:CSDb

[CC61a] Arthur L. Colwin and Laura Hunter Colwin. Changes in the spermatozoon during fertilization in hydroides hexagonus (annelida). Journal of Cell Biology, 10(2):255–??, June 1961. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/10/2/255.

Colwin:1961:FSS

[CC61b] Arthur L. Colwin and Laura Hunter Colwin. Fine structure of the spermatozoon of hydroides hexagonus (annelida), with special reference to the acrosomal region. Journal of Cell Biology, 10(2):211–??, June 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/10/2/211.

Colwin:1961:CSDa

[CC61c] Laura Hunter Colwin and Arthur L. Colwin. Changes in the spermatozoon during fertilization in hydroides hexagonus (annelida). Journal of Cell Biology, 10(2):231–??, June 1961. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/10/2/231.

Colwin:1963:RGMa

[CC63a] Arthur L. Colwin and Laura Hunter Colwin. Role of the gamete membranes in fertilization in saccoglossus kowalevskii (enteropneusta). Journal of Cell Biology, 19(3):477–??, December 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/19/3/477.

Colwin:1963:RGMb

[CC63b] Laura Hunter Colwin and Arthur L. Colwin. Role of the gamete membranes in fertilization in saccoglossus kowalevskii (enteropneusta). Journal of Cell Biology, 19(3):501–??, December 1963.

CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/19/3/501.

Cameron:1964:IMR

[CC64a] Ivan L. Cameron and Günter Cleffmann. Initiation of mitosis in relation to the cell cycle following feeding of starved chickens. Journal of Cell Biology, 21(2):169–??, May 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/21/2/169.

Cohen:1964:IAF

[CC64b] Alan S. Cohen and Evan Calkins. The isolation of amyloid fibrils and a study of the effect of collagenase and hyaluronidase. Journal of Cell Biology, 21(3):481–??, June 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/21/3/481.

Campo:1961:CPC

[CD61a] Robert D. Campo and Dominic D. Dziewiatkowski. A consideration of the permeability of cartilage to inorganic sulfate. Journal of Cell Biology, 9(2):401–??, February 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/9/2/401.

Chapman:1961:FSL

[CD61b] George B. Chapman and Alden B. Dawson. Fine structure of the larval anuran epidermis, with special reference to the figures of eberth. Journal of Cell Biology, 10(3):425–??, July 1961. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/10/3/425.

Campo:1963:TOM

[CD63] Robert D. Campo and Dominic D. Dziewiatkowski. Turnover of the organic matrix of cartilage and bone as visualized by autoradiography. Journal of Cell Biology, 18(1):19–??, July 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/18/1/19.

Caro:1961:LMEb

[CF61a] Lucien G. Caro and Frederick Forro. Localization of macromolecules in Escherichia coli. Journal of Cell Biology, 9(3): 555–??, March 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/9/3/555.

Christensen:1961:NFS

[CF61b] A. Kent Christensen and Don W. Fawcett. The normal fine structure of opossum testicular interstitial cells. Journal of Cell Biology, 9(3):653–??, March 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/9/3/653.

Craig:1962:ETE

[CFG62] Elson L. Craig, Walter J. Frajola, and Marie H. Greider. An embedding technique for electron microscopy using epon 812. Journal of Cell Biology, 12(1):190–??, January 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/12/1/190.

Callaghan:1961:ICL

[CG61] John J. Callaghan and Paul Grun. Incorporation of C^{14}-labeled maleic hydrazide into the root-tip cells of Allium Cernuum, Vicia Faba, and Tradescantia Paludosa. Journal of Cell Biology, 10(4):567–??, August 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/10/4/567.

Cameron:1963:EEC

[CG63] Ivan L. Cameron and Richard C. Greulich. Evidence for an essentially constant duration of DNA synthesis in renewing epithelia of the adult mouse. Journal of Cell Biology, 18(1):31–??, July 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
18/1/31.

Callahan:1964:UVS

[CH64] William P. Callahan and Jack A. Horner. The use of vanadium as a stain for electron microscopy. Journal of Cell Biology, 20(2):350–??, February 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/20/2/350.

Chapman:1961:MCS

[Cha61] J. A. Chapman. Morphological and chemical studies of collagen formation. Journal of Cell Biology, 9(3):639–??, March 1961.

CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/9/3/639.

Chandra:1962:RMM

[Cha62] Satish Chandra. The reversal of mitochondrial membrane. Journal of Cell Biology, 12(3):503–??, March 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/12/3/503.

Condie:1961:SPP

[CHG61] Richard M. Condie, A. Ervin Howell, and Robert A. Good. Studies on the problem of preservation of myelin sheath ultrastructure: Evaluation of fixation, dehydration, and embedding techniques. Journal of Cell Biology, 9(2):429–??, February 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/9/2/429.

Choi:1963:FSU

[Cho63] Jae Kwon Choi. The fine structure of the urinary bladder of the toad, bufo marinus. Journal of Cell Biology, 16(1):53–??, January 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
16/1/53.

Carlsen:1961:URC

[CKB61] F. Carlsen, G. G. Knappeis, and F. Buchthal. Ultrastructure of the resting and contracted striated muscle fiber at different degrees of stretch. Journal of Cell Biology, 11(1):95–??, October 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/11/1/
95.

Couffer-Kaltenbach:1961:AED

[CKP61] Jane Couffer-Kaltenbach and Peter Perlmann. Antigens in eggs and developmental stages of the sea urchin. Journal of Cell Biology, 9(1):93–??, January 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/9/1/93.

Clark:1960:EMS

[Cla60] Wallace H. Clark. Electron Microscope Studies of Nuclear Extrusions in Pancreatic Acinar Cells of the Rat. Journal of Cell Biology, 7(2):345–??, April 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/7/2/345.

Coleman:1964:DFS

[CM64] James R. Coleman and Montrose J. Moses. DNA and the fine structure of synaptic chromosomes in the domestic rooster (Gallus Domesticus). Journal of Cell Biology, 23(1):63–??, October 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/23/1/
63.

Chauveau:1962:ISV

[CMRS62] J. Chauveau, Y. Moulé, C. Rouiller, and J. Schneebeli. Isolation of smooth vesicles and free ribosomes from rat liver microsomes. Journal of Cell Biology, 12(1):17–??, January 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/12/1/17.

Colvin:1963:SCM

[Col63] J. Ross Colvin. The size of the cellulose microfibril. Journal of Cell Biology, 17(1):105–??, April 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/17/1/105.

Copeland:1964:MPC

[Cop64] Eugene Copeland. A mitochondrial pump in the cells of the anal papillae of mosquito larvae. Journal of Cell Biology, 23(2): 253–??, November 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/23/2/253.

Cosmos:1964:IDC

[Cos64] Ethel Cosmos. Intracellular distribution of calcium in developing breast muscle of normal and dystrophic chickens. Journal of Cell Biology, 23(2):241–??, November 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/23/2/241.

Crocker:1960:IRL

[CP60] T. Timothy Crocker and S. R. Pelc. Incorporation and rate of loss of S^{35}-methionine by adult rat trachea in organ cultures and in living animals. Journal of Cell Biology, 7(2):283–??, April 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/7/
2/283.

Caro:1964:PSS

[CP64] Lucien G. Caro and George E. Palade. Protein synthesis, storage, and discharge in the pancreatic exocrine cell. Journal of Cell Biology, 20(3):473–??, March 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/20/3/473.

Cauna:1960:FSM

[CR60] Nikolajs Cauna and Leonard L. Ross. The fine structure of Meissner’s touch corpuscles of human fingers. Journal of Cell Biology, 8(2):467–??, October 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/8/2/467.

Claus:1964:FSG

[CR64] G. W. Claus and L. E. Roth. Fine structure of the Gramnegative bacterium acetobacter suboxydans. Journal of Cell Biology, 20(2):217–??, February 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/20/2/217.

Casley-Smith:1962:ICL

[CS62] J. R. Casley-Smith. The identification of chylomicra and lipoproteins in tissue sections and their passage into jejunal lacteals. Journal of Cell Biology, 15(2):259–??, November 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/15/2/259.

Cameron:1964:CRB

[CS64a] Ivan L. Cameron and Robert E. Smith. Cytological responses of Brown fat tissue in cold-exposed rats. Journal of Cell Biology, 23(1):89–??, October 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/23/1/89.

Cohen:1964:EMC

[CS64b] Maimon M. Cohen and Margery W. Shaw. Effects of mitomycin C on human chromosomes. Journal of Cell Biology, 23 (2):386–??, November 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/23/2/386.

Chatterjee:1961:EMS

[CSC61] S. N. Chatterjee, P. Sadhukhan, and J. B. Chatterjea. Electron microscope studies on the haemoglobin molecules. Journal of Cell Biology, 10(1):113–??, May 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/10/1/113.

Csillik:1963:SOP

[Csi63] Bertalan Csillik. Submicroscopic organization of the postsynaptic membrane in the myoneural junction. Journal of Cell Biology, 17(3):571–??, June 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/17/3/571.

Curtis:1964:MAC

[Cur64] A. S. G. Curtis. The mechanism of adhesion of cells to glass. Journal of Cell Biology, 20(2):199–??, February 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/20/2/199.

Caro:1962:HRAa

[CvT62] Lucien G. Caro and Robert P. van Tubergen. High-resolution autoradiography. Journal of Cell Biology, 15(2):173–??, November 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/15/2/
173.

Chambers:1964:ALS

[CW64] Velma C. Chambers and Russell S. Weiser. Annulate lamellae in sarcoma I cells. Journal of Cell Biology, 21(1):133–??, April 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/21/1/
133.

Clementi:1963:MPE

[CZ63] Francesco Clementi and Gian Paolo Zocche. Morphological and pharmacological effects of reserpine, given alone or after iproniazid, on the catechol amines of the adrenal glands of the rat. Journal of Cell Biology, 17(3):587–??, June 1963. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/17/3/587.

Dales:1962:EMS

[Dal62] Samuel Dales. An electron microscope study of the early association between two mammalian viruses and their hosts. Journal of Cell Biology, 13(2):303–??, May 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/13/2/303.

Dales:1963:UDV

[Dal63] Samuel Dales. The uptake and development of vaccinia virus in strain L cells followed with labeled viral deoxyribonucleic acid. Journal of Cell Biology, 18(1):51–??, July 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/18/1/51.

Danon:1961:ITP

[Dan61] David Danon. An instrument to trim plastic specimen blocks for electron microscopy prior to sectioning. Journal of Cell Biology, 9(3):726–??, March 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/9/3/726.

Das:1962:SCC

[Das62] Nirmal K. Das. Synthetic capacities of chromosome fragments correlated with their ability to maintain nucleolar material. Journal of Cell Biology, 15(1):121–??, October 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/15/1/121.

Davies:1961:SNE

[Dav61] Howard G. Davies. Structure in nucleated erythrocytes. Journal of Cell Biology, 9(3):671–??, March 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/9/3/671.

Dannenberg:1964:HER

[DB64a] Arthur M. Dannenberg and William E. Bennett. Hydrolytic enzymes of rabbit mononuclear exudate cells. Journal of Cell Biology, 21(1):1–??, April 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/21/1/1.

Dewey:1964:SSD

[DB64b] M. M. Dewey and L. Barr. A study of the structure and distribution of the nexus. Journal of Cell Biology, 23(3):553–??, December 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/23/3/
553.

deBeyer:1962:AAI

[dBdMP62] J. M. de Beyer, J. C. H. de Man, and J-P. Persijn. Atpase activity on the intercalated disc and C_{z} bands of mouse heart muscle. Journal of Cell Biology, 13(3):452–??, June 1962. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/13/3/452.

Dannenberg:1963:HSP

[DBWK63] Arthur M. Dannenberg, Marvin S. Burstone, Paul C. Walter, and June W. Kinsley. A histochemical study of phagocytic and enzymatic functions of rabbit mononuclear and polymorphonuclear exudate cells and alveolar macrophages. Journal of Cell Biology, 17(3):465–??, June 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/17/3/465.

Duck-Chong:1964:RBI

[DCPN64] Coral Duck-Chong, J. K. Pollak, and R. J. North. The relation between the intracellular ribonucleic acid distribution and amino acid incorporation in the liver of the developing chick embryo. Journal of Cell Biology, 20(1):25–??, January 1964. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/20/1/25.

DeRobertis:1961:PSP

[DdI61] Eduardo De Robertis and Amanda Pellegrino de Iraldi. Plurivesicular secretory processes and nerve endings in the pineal gland of the rat. Journal of Cell Biology, 10(3):361–??, July 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
10/3/361.

Dales:1962:CCF

[DF62] Samuel Dales and Richard M. Franklin. A comparison of the changes in fine structure of L cells during single cycles of viral multiplication, following their infection with the viruses of mengo and encephalomyocarditis. Journal of Cell Biology, 14 (2):281–??, August 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/14/2/281.

Lima-de-Faria:1962:PDS

[dFB62] A. Lima de Faria and Kirstine Borum. The period of DNA synthesis prior to meiosis in the mouse. Journal of Cell Biology, 14(3):381–??, September 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/14/3/381.

Lima-de-Faria:1963:HHM

[dFR63] A. Lima de Faria and J. Reitalu. Heterochromatin in human male leukocytes. Journal of Cell Biology, 16(2):315–??, February 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/16/2/
315.

Dowling:1962:FSP

[DG62] John E. Dowling and I. R. Gibbons. The fine structure of the pigment epithelium in the albino rat. Journal of Cell Biology, 14(3):459–??, September 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/14/3/459.

deHarven:1960:FEM

[dHF60] Etienne de Harven and Charlotte Friend. Further Electron Microscope Studies of a Mouse Leukemia Induced by Cell–Free Filtrates. Journal of Cell Biology, 7(4):747–??, July 1960. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/7/4/747.

Deshpande:1961:MIR

[DHV61] P. D. Deshpande, D. D. Hickman, and R. W. Von Korff. Morphology of isolated rabbit heart muscle mitochondria and the oxidation of extramitochondrial reduced diphosphopyridine nucleotide. Journal of Cell Biology, 11(1):77–??, October 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/11/1/77.

Dirksen:1961:PCA

[Dir61] E. Roter Dirksen. The presence of centrioles in artificially activated sea urchin eggs. Journal of Cell Biology, 11(1):244–??, October 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
11/1/244.

Defendi:1960:HSD

[DK60] V. Defendi and D. Kritchevsky. A Histoautoradiographic Study of Deoxyribonucleic Acid Synthesis in Tissue Cultures of Chicken Embryo Liver Cells Infected with RPL-12 Lymphomatosis Virus. Journal of Cell Biology, 7(1):201–??, February 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
7/1/201.

Das:1964:HPT

[DKG64] C. C. Das, B. P. Kaufmann, and Helen Gay. Histone protein transition in Drosophila Melanogaster. Journal of Cell Biology, 23(3):423–??, December 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/23/3/423.

deLorenzo:1960:FSS

[dL60] A. J. de Lorenzo. The Fine Structure of Synapses in the Ciliary Ganglion of the Chick. Journal of Cell Biology, 7(1): 31–??, February 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/7/1/31.

Dickson:1963:RMD

[DL63] J. A. Dickson and I. Leslie. The reaggregation and metabolism of dissociated adult guinea pig cells in filter-well cultures. Journal of Cell Biology, 19(2):449–??, November 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/19/2/449.

Danon:1961:CEM

[DML61] D. Danon, Y. Marikovsky, and U. Z. Littauer. A comparative electron microscopical study of RNA from different sources. Journal of Cell Biology, 9(2):253–??, February 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/9/2/253.

Dallner:1963:IPR

[DOB63] G. Dallner, S. Orrenius, and A. Bergstrand. Isolation and properties of rough and smooth vesicles from rat liver. Journal of Cell Biology, 16(2):426–??, February 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/16/2/426.

DeRobertis:1961:INE

[DPRG61] E. De Robertis, A. Pellegrino De Iraldi, Georgina Rodriguez, and C. J. Gomez. On the isolation of nerve endings and synaptic vesicles. Journal of Cell Biology, 9(1):229–??, January 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/9/1/229.

Daniels:1964:EMM

[DR64] E. W. Daniels and L. E. Roth. Electron microscopy of mitosis in a radiosensitive giant Amoeba. Journal of Cell Biology, 20(1):75–??, January 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/20/1/75.

Drochmans:1960:EMS

[Dro60a] P. Drochmans. Electron microscope studies of epidermal melanocytes, and the fine structure of melanin granules. Journal of Cell Biology, 8(1):165–??, September 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/8/1/165.

Drochmans:1960:MED

[Dro60b] P. Drochmans. Mise en Évidence du GlycogÉNe dans la cellule HÉPatique Par microscopie Électronique. Journal of Cell Biology, 8(2):553–??, October 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/8/2/553.

Drum:1963:CFS

[Dru63] Ryan W. Drum. The cytoplasmic fine structure of the diatom, nitzschia palea. Journal of Cell Biology, 18(2):429–??, August 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/18/2/

Dales:1961:DVV

[DS61a] S. Dales and L. Siminovitch. The development of vaccinia virus in Earle’s L strain cells as examined by electron microscopy. Journal of Cell Biology, 10(4):475–??, August 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/10/4/475.

Danes:1961:IID

[DS61b] B. Shannon Danes and Mary Struthers. The influence of irondextran complex on the strain L fibroblast. Journal of Cell Biology, 10(2):289–??, June 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/10/2/289.

Dudley:1961:FSB

[DS61c] H. Robert Dudley and David Spiro. The fine structure of bone cells. Journal of Cell Biology, 11(3):627–??, December 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/11/3/627.

Davies:1962:VSE

[DS62a] Howard G. Davies and Michael Spencer. The variation in the structure of erythrocyte nuclei with fixation. Journal of Cell Biology, 14(3):445–??, September 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/14/3/445.

Dowling:1962:IRD

[DS62b] John E. Dowling and Richard L. Sidman. Inherited retinal dystrophy in the rat. Journal of Cell Biology, 14(1):73–??, July 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/14/1/
73.

de-The:1964:CMD

[dT64] Guy de Thé. Cytoplasmic microtubules in different animal cells. Journal of Cell Biology, 23(2):265–??, November 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/23/2/265.

deFWebster:1962:TFA

[dW62] Henry deF. Webster. Transient, focal accumulation of axonal mitochondria during the early stages of Wallerian degeneration.

Journal of Cell Biology, 12(2):361–??, February 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/12/2/361.

Dalton:1960:SMS

[DZ60] Albert J. Dalton and Robert F. Zeigel. A Simplified Method of Staining Thin Sections of Biological Material with Lead Hydroxide for Electron Microscopy. Journal of Cell Biology, 7 (2):409–??, April 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/7/2/409.

Dziewiatkowski:1962:ISC

[Dzi62] Dominic D. Dziewiatkowski. Intracellular synthesis of chondroitin sulfate. Journal of Cell Biology, 13(3):359–??, June 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/13/3/
359.

Emmelot:1960:CFS

[EB60] P. Emmelot and E. L. Benedetti. Changes in the Fine Structure of Rat Liver Cells Brought about by Dimethylnitrosamine. Journal of Cell Biology, 7(2):393–??, April 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/7/2/393.

Edstrom:1962:BCN

[EB62] Jan-Erik Edström and Wolfgang Beermann. The base composition of nucleic acids in chromosomes, puffs, nucleoli, and cytoplasm of chironomus salivary gland cells. Journal of Cell Biology, 14(3):371–??, September 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/14/3/371.

Elliott:1964:FMD

[EB64] Alfred M. Elliott and Il Jin Bak. The fate of mitochondria during aging in Tetrahymena Pyriformis. Journal of Cell Biology, 20(1):113–??, January 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/20/1/113.

Echlin:1963:FSC

[Ech63] Patrick Echlin. The fine structure of the cell boundary of the blue-green alga anacystis Montana. Journal of Cell Biology, 17 (1):212–??, April 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/17/1/212.

Edstrom:1960:CRA

[Eds60a] Jan-Erik Edström. Composition of ribonucleic acid from various parts of spider oocytes. Journal of Cell Biology, 8(1): 47–??, September 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/8/1/47.

Edstrom:1960:EHE

[Eds60b] Jan-Erik Edström. Extraction, hydrolysis, and electrophoretic analysis of ribonucleic acid from microscopic tissue units (microphoresis). Journal of Cell Biology, 8(1):39–??, September 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/8/1/
39.

Enders:1963:EMS

[EES63] A. C. Enders, R. K. Enders, and S. Schlafke. An electron microscope study of the gland cells of the mink endometrium. Journal of Cell Biology, 18(2):405–??, August 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/18/2/405.

Edstrom:1963:BCR

[EG63] J.-E. Edström and J. G. Gall. The base composition of ribonucleic acid in lampbrush chromosomes, nucleoli, nuclear sap, and cytoplasm of Triturus oocytes. Journal of Cell Biology, 19(2): 279–??, November 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/19/2/279.

Easton:1962:DSA

[EGG62] John M. Easton, Burton Goldberg, and Howard Green. Demonstration of surface antigens and pinocytosis in mammalian cells with ferritin-antibody conjugates. Journal of Cell Biology, 12 (2):437–??, February 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/12/2/437.

Edstrom:1961:IDH

[EGS61] Jan-Erik Edström, Wolfgang Grampp, and Norberto Schor. The intracellular distribution and heterogeneity of ribonucleic acid in starfish oocytes. Journal of Cell Biology, 11(3):549–??, December 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
11/3/549.

Egyhazi:1961:EIC

[EH61a] Endre Egyházi and Holger Hydén. Experimentally induced changes in the base composition of the ribonucleic acids of isolated nerve cells and their oligodendroglial cells. Journal of Cell Biology, 10(3):403–??, July 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/10/3/403.

Estabrook:1961:SCO

[EH61b] Ronald W. Estabrook and Andrew Holowinsky. Studies on the content and organization of the respiratory enzymes of mitochondria. Journal of Cell Biology, 9(1):19–??, January 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/9/1/19.

Epstein:1963:EMO

[EH63a] M. A. Epstein and S. J. Holt. Electron microscope observations on the surface adenosine triphosphatase-like enzymes of HeLa cells infected with herpes virus. Journal of Cell Biology, 19(2):337–??, November 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/19/2/337.

Epstein:1963:LEM

[EH63b] M. A. Epstein and S. J. Holt. The localization by electron microscopy of HeLa cell surface enzymes splitting adenosine triphosphate. Journal of Cell Biology, 19(2):325–??, November 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/19/2/

Ehrlich:1960:EMD

[Ehr60] H. G. Ehrlich. Electron Microscopy of the Degeneration of Fine Structure in Saintpaulia ionantha Wendl. Pollen Walls. Journal of Cell Biology, 7(1):199–??, February 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/7/1/199.

Ernster:1962:SNT

[EJ62] Lars Ernster and Lois C. Jones. A study of the nucleoside triand Diphosphate activities of rat liver microsomes. Journal of Cell Biology, 15(3):563–??, December 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/15/3/563.

Edstrom:1961:MDA

[EK61] Jan-Erik Edström and Jerzy Kawiak. Microchemical deoxyribonucleic acid determination in individual cells. Journal of Cell Biology, 9(3):619–??, March 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/9/3/619.

Eilam:1962:ELI

[EK62] Yael Eilam and S. Klein. The effect of light intensity and sucrose feeding on the fine structure in chloroplasts and on the chlorophyll content of etiolated leaves. Journal of Cell Biology, 14(2):169–??, August 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/14/2/169.

Elliott:1962:MES

[EKB62] Alfred M. Elliott, John R. Kennedy, and Il Jin Bak. Macronuclear events in synchronously dividing Tetrahymena Pyriformis. Journal of Cell Biology, 12(3):515–??, March 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/12/3/515.

Edwards:1960:SCD

[EKY^{+}60] Joshua L. Edwards, Arthur L. Koch, Pauline Youcis, Herbert L. Freese, Melville B. Laite, and J. Thomas Donalson. Some Characteristics of DNA Synthesis and the Mitotic Cycle in Ehrlich Ascites Tumor Cells. Journal of Cell Biology, 7(2):273–??, April 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/7/
2/273.

Ebashi:1962:ATL

[EL62] Setsuro Ebashi and Fritz Lipmann. Adenosine triphosphatelinked concentration of calcium ions in a particulate fraction of rabbit muscle. Journal of Cell Biology, 14(3):389–??, September 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/14/3/
389.

Enders:1964:OFS

[EL64] Allen C. Enders and W. R. Lyons. Observations on the fine structure of lutein cells. Journal of Cell Biology, 22(1):127–??, July 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
22/1/127.

Ellis:1961:EMD

[EM61] Richard A. Ellis and William Montagna. Electron microscopy of the Duct, and especially the “cuticular border” of the eccrine sweat glands in macaca mulatta. Journal of Cell Biology, 9(1):238–??, January 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/9/1/238.

Emmelot:1962:CFS

[EMNB62] P. Emmelot, I. J. Mizrahi, R. Naccarato, and E. L. Benedetti. Changes in function and structure of the endoplasmic reticulum of rat liver cells after administration of cysteine. Journal of Cell Biology, 12(1):177–??, January 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/12/1/177.

Essner:1961:LAP

[EN61] E. Essner and Alex B. Novikoff. Localization of acid phosphatase activity in hepatic lysosomes by means of electron microscopy. Journal of Cell Biology, 9(4):773–??, April 1961. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/9/4/773.

Essner:1962:CST

[EN62] Edward Essner and Alex B. Novikoff. Cytological studies on two functional hepatomas. Journal of Cell Biology, 15(2): 289–??, November 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/15/2/289.

Enders:1962:OFS

[End62] Allen C. Enders. Observations on the fine structure of lutein cells. Journal of Cell Biology, 12(1):101–??, January 1962. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/12/1/101.

Epel:1963:ECM

[Epe63] David Epel. The effects of carbon monoxide inhibition on ATP level and the rate of mitosis in the sea urchin egg. Journal of Cell Biology, 17(2):315–??, May 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/17/2/315.

Epstein:1961:SUF

[Eps61] M. A. Epstein. Some unusual features of fine structure observed in HeLa cells. Journal of Cell Biology, 10(2):153–??, June 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/10/2/153.

Epstein:1962:OMR

[Eps62] M. A. Epstein. Observations on the mode of release of herpes virus from infected HeLa cells. Journal of Cell Biology, 12 (3):589–??, March 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/12/3/589.

Eichel:1962:ILE

[ER62] Herbert J. Eichel and Jay S. Roth. Intracellular localization of enzymes in spleen. Journal of Cell Biology, 12(2):263–??, February 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/12/2/
263.

Erlandson:1964:NMD

[Erl64] Robert A. Erlandson. A new Maraglas, D.E.R.(R) 732, embedment for electron microscopy. Journal of Cell Biology, 22(3):

704–??, September 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/22/3/704.

Evans:1963:RBD

[ES63] H. J. Evans and J. R. K. Savage. The relation between DNA synthesis and chromosome structure as resolved by X-ray damage. Journal of Cell Biology, 18(3):525–??, September 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/18/3/525.

Ernster:1962:ESR

[ESP62] Lars Ernster, Philip Siekevitz, and George E. Palade. Enzymestructure relationships in the endoplasmic reticulum of rat liver. Journal of Cell Biology, 15(3):541–??, December 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/15/3/541.

Essner:1960:EMS

[Ess60] Edward Essner. An Electron Microscopic Study of Erythrophagocytosis. Journal of Cell Biology, 7(2):329–??, April 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/7/2/
329.

Eakin:1960:FOF

[EW60] Richard M. Eakin and Jane A. Westfall. Further observations on the fine structure of the parietal eye of lizards. Journal of Cell Biology, 8(2):483–??, October 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/8/2/483.

Eakin:1962:FSN

[EW62] Richard M. Eakin and Jane A. Westfall. Fine structure of the notochord of amphioxus. Journal of Cell Biology, 12(3):646–??, March 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
12/3/646.

Eakin:1964:FSE

[EW64] Richard M. Eakin and Jane A. Westfall. Fine structure of the eye of a chaetognath. Journal of Cell Biology, 21(1):115–??, April 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
21/1/115.

Franzini-Armstrong:1963:PSR

[FA63] Clara Franzini-Armstrong. Pores in the sarcoplasmic reticulum. Journal of Cell Biology, 19(3):637–??, December 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/19/3/637.

Fahimi:1964:CLL

[FA64] H. Dariush Fahimi and Chandra Raj Amarasingham. Cytochemical localization of lactic dehydrogenase in white skeletal muscle. Journal of Cell Biology, 22(1):29–??, July 1964. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/22/1/29.

Fahrenbach:1963:CGK

[Fah63a] Wolf H. Fahrenbach. A contribution to glass knife breaking. Journal of Cell Biology, 18(2):475–??, August 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/18/2/475.

Fahrenbach:1963:SRS

[Fah63b] Wolf H. Fahrenbach. The sarcoplasmic reticulum of striated muscle of a cyclopoid copepod. Journal of Cell Biology, 17 (3):629–??, June 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/17/3/629.

Fahrenbach:1964:NCS

[Fah64] Wolf H. Fahrenbach. A new configuration of the sarcoplasmic reticulum. Journal of Cell Biology, 22(2):477–??, August 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/22/2/477.

Franzini-armstrong:1964:SIC

[FaP64] Clara Franzini-armstrong and Keith R. Porter. Sarcolemmal invaginations constituting the T system in fish muscle fibers. Journal of Cell Biology, 22(3):675–??, September 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/22/3/675.

Farbman:1964:EMS

[Far64] Albert I. Farbman. Electron microscope study of a small cytoplasmic structure in rat oral epithelium. Journal of Cell Biology, 21(3):491–??, June 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/21/3/491.

Feldherr:1962:IDF

[Fel62a] Carl M. Feldherr. The intracellular distribution of ferritin following microinjection. Journal of Cell Biology, 12(1):159–??, January 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
12/1/159.

Feldherr:1962:NAP

[Fel62b] Carl M. Feldherr. The nuclear annuli as pathways for nucleocytoplasmic exchanges. Journal of Cell Biology, 14(1):65–??, July 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/14/1/
65.

Feldman:1962:MST

[Fel62c] Dorothy Gelber Feldman. A method of staining thin sections with lead hydroxide for precipitate-free sections. Journal of Cell Biology, 15(3):592–??, December 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/15/3/592.

Feldherr:1964:BWN

[Fel64] Carl M. Feldherr. Binding within the nuclear annuli and its possible effect on nucleocytoplasmic exchanges. Journal of Cell Biology, 20(1):188–??, January 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/20/1/188.

Fischman:1964:MSI

[FG64] Donald A. Fischman and Michael D. Gershon. A method for studying intracellular movement of water-soluble isotopes prior to radioautography. Journal of Cell Biology, 21(1):139–??, April 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/21/1/

Fisher:1963:DGI

[FHI63] Richard F. Fisher, David J. Holbrook, and J. Logan Irvin. Density gradient isolation of rat liver nuclei with high DNA content. Journal of Cell Biology, 17(2):231–??, May 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/17/2/231.

Finck:1960:ERE

[Fin60a] Henry Finck. Epoxy Resins in Electron Microscopy. Journal of Cell Biology, 7(1):27–??, February 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/7/1/27.

Finean:1960:EMXa

[Fin60b] J. B. Finean. Electron microscope and X-ray diffraction studies of the effects of dehydration on the structure of nerve myelin. Journal of Cell Biology, 8(1):13–??, September 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/8/1/13.

Finean:1960:EMXb

[Fin60c] J. B. Finean. Electron microscope and X-ray diffraction studies of the effects of dehydration on the structure of nerve myelin. Journal of Cell Biology, 8(1):31–??, September 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/8/1/31.

Fitz-James:1960:PCM

[FJ60] Philip C. Fitz-James. Participation of the cytoplasmic membrane in the growth and spore formation of bacilli. Journal of Cell Biology, 8(2):507–??, October 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/8/2/507.

Finger:1960:ISI

[FKKH60] Irving Finger, Michael Kaback, Philip Kittner, and Carol Heller. Immunological studies of isolated particulates of paramecium aurelia. Journal of Cell Biology, 8(3):591–??, December 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/8/
3/591.

Faberge:1962:VWC

[FL62] A. C. Fabergé and C. Willard Lewis. N-vinylpyrrolidone as a water compatible constituent of embedding resins for sectioning in electron microscopy. Journal of Cell Biology, 15(3):579–??, December 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
15/3/579.

Flock:1964:SMU

[Flo64] Åke Flock. Structure of the macula utriculi with special reference to directional interplay of sensory responses as revealed by morphological polarization. Journal of Cell Biology, 22(2): 413–??, August 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/22/2/413.

Feldherr:1962:UCG

[FM62] Carl M. Feldherr and J. M. Marshall. The use of colloidal gold for studies of intracellular exchanges in the ameba chaos chaos. Journal of Cell Biology, 12(3):640–??, March 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/12/3/640.

Frist:1964:FDB

[FM64a] Ramsey H. Frist and Morton D. Maser. Freeze-drying of biological specimens for electron microscopy using thermoelectric cooling. Journal of Cell Biology, 22(1):293–??, July 1964. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/22/1/293.

Fujita:1964:FSI

[FM64b] Hisao Fujita and Mitsuo Machino. Fine structure of intramitochondrial crystals in rat thyroid follicular cell. Journal of Cell Biology, 23(2):383–??, November 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/23/2/383.

Fernandez-Moran:1964:MRU

[FMOBG64] H. Fernández-Morán, T. Oda, P. V. Blair, and D. E. Green. A macromolecular repeating unit of mitochondrial structure and function. Journal of Cell Biology, 22(1):63–??, July 1964. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/22/1/63.

Farquhar:1960:SFG

[FP60] Marilyn G. Farquhar and George E. Palade. Segregation of Ferritin in Glomerular Protein Absorption Droplets. Journal of Cell Biology, 7(2):297–??, April 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/7/2/297.

Farquhar:1962:FEE

[FP62] Marilyn G. Farquhar and George E. Palade. Functional evidence for the existence of a third cell type in the renal glomerulus. Journal of Cell Biology, 13(1):55–??, April 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/13/1/55.

Farquhar:1963:JCV

[FP63] Marilyn G. Farquhar and George E. Palade. Junctional complexes in various epithelia. Journal of Cell Biology, 17(2):375– ??, May 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
17/2/375.

Fawcett:1961:SRF

[FR61] Don W. Fawcett and Jean Paul Revel. The sarcoplasmic reticulum of a fast-acting fish muscle. Journal of Cell Biology, 10 (4):89–??, August 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/10/4/89.

Filshie:1962:EMS

[FR62] B. K. Filshie and G. E. Rogers. An electron microscope study of the fine structure of feather keratin. Journal of Cell Biology, 13(1):1–??, April 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/13/1/1.

Franchi:1960:EMO

[Fra60] Leslie L. Franchi. Electron Microscopy of Oocyte–Follicle Cell Relationships in the Rat Ovary. Journal of Cell Biology, 7 (2):397–??, April 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/7/2/397.

Franklin:1962:CDM

[Fra62] Richard M. Franklin. A cytochemical description of the multiplication of mengovirus in L-929 cells. Journal of Cell Biology, 12(1):1–??, January 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/12/1/1.

Frei:1962:FSB

[Fre62] J. V. Frei. The fine structure of the basement membrane in epidermal tumors. Journal of Cell Biology, 15(2):335–??, November 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/15/2/
335.

Friede:1964:ERA

[Fri64] Reinhard L. Friede. The enzymatic response of astrocytes to various ions in vitro. Journal of Cell Biology, 20(1):5–??, January 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
20/1/5.

Frei:1961:CIC

[FS61a] J. V. Frei and H. Sheldon. Corpus intra cristam: a dense body within mitochondria of cells in hyperplastic mouse epidermis. Journal of Cell Biology, 11(3):724–??, December 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/11/3/724.

Frei:1961:SGC

[FS61b] J. V. Frei and H. Sheldon. A small granular component of the cytoplasm of keratinizing epithelia. Journal of Cell Biology, 11(3):719–??, December 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/11/3/719.

Freeman:1962:NEE

[FS62] James A. Freeman and Ben O. Spurlock. A new epoxy embedment for electron microscopy. Journal of Cell Biology, 13 (3):437–??, June 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/13/3/437.

Fand:1964:PAM

[FS64] Sally B. Fand and Richard P. Spencer. Off-peak absorption measurements in Feulgen cytophotometry. Journal of Cell Biology, 22(3):515–??, September 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/22/3/515.

Fand:1962:SPA

[FT62] Sally B. Fand and Bo Thorell. Spectrophotometry of the periodic acid-Schiff reaction with pituitary hormones in vitro and in histological sections. Journal of Cell Biology, 13(2):239–??, May 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/13/2/
239.

Feldman:1961:MRF

[FVK61] Joseph D. Feldman, Jacinto J. Vazquez, and Stanley M. Kurtz. Maturation of the rat fetal thyroid. Journal of Cell Biology, 11 (2):365–??, November 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/11/2/365.

Flock:1962:SOS

[FW62a] Åke Flock and Jan Wersäll. A study of the orientation of the sensory hairs of the receptor cells in the lateral line organ of fish, with special reference to the function of the receptors. Journal of Cell Biology, 15(1):19–??, October 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/15/1/19.

Flock:1962:SSL

[FW62b] Åke Flock and Jan Wersäll. Synaptic structures in the lateral line canal organ of the teleost fish lota vulgaris. Journal of Cell Biology, 13(2):337–??, May 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/13/2/337.

Gantt:1963:CDG

[GA63] Elisabeth Gantt and Howard J. Arnott. Chloroplast division in the gametophyte of the fern matteuccia struthiopteris (l.) todaro. Journal of Cell Biology, 19(2):446–??, November 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/19/2/446.

Gall:1961:CR

[Gal61] Joseph G. Gall. Centriole replication. Journal of Cell Biology, 10(2):163–??, June 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/10/2/163.

Galton:1962:DCP

[Gal62] Michael Galton. DNA content of placental nuclei. Journal of Cell Biology, 13(2):183–??, May 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/13/2/183.

Garcia:1964:FDV

[Gar64] Alfredo Mariano Garcia. Feulgen-DNA values in megakaryocytes. Journal of Cell Biology, 20(2):342–??, February 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/20/2/342.

Gasic:1963:HSS

[GB63] Gabriel Gasic and Leonard Berwick. Hale stain for sialic acidcontaining mucins. Journal of Cell Biology, 19(1):223–??, October 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/19/1/
223.

Gross:1963:SSA

[GC63] Paul R. Gross and Gilles H. Cousineau. Synthesis of spindleassociated proteins in early cleavage. Journal of Cell Biology, 19 (1):260–??, October 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/19/1/260.

Glauert:1963:SMA

[GDLD63] Audrey M. Glauert, Mary R. Daniel, J. A. Lucy, and J. T. Dingle. Studies on the mode of action of excess of vitamin A. Journal of Cell Biology, 17(1):111–??, April 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/17/1/111.

Grimley:1960:UCD

[GE60] Philip M. Grimley and George A. Edwards. The ultrastructure of cardiac desmosomes in the toad and their relationship to the intercalated disc. Journal of Cell Biology, 8(2):305–??, October 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/8/
2/305.

German:1964:PDS

[Ger64] James German. The pattern of DNA synthesis in the chromosomes of human blood cells. Journal of Cell Biology, 20 (1):37–??, January 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/20/1/37.

Gamble:1960:SH

[GG60a] Charles N. Gamble and David Glick. Studies in histochemistry. Journal of Cell Biology, 8(1):53–??, September 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/8/1/53.

Gibbons:1960:FSC

[GG60b] I. R. Gibbons and A. V. Grimstone. On Flagellar Structure in Certain Flagellates. Journal of Cell Biology, 7(4):697–??, July 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/7/
4/697.

Goldberg:1960:IC

[GG60c] Burton Goldberg and Howard Green. Immune Cytolysis. Journal of Cell Biology, 7(4):645–??, July 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/7/4/645.

Gibor:1962:USF

[GG62] Aharon Gibor and Sam Granick. Ultraviolet sensitive factors in the cytoplasm that affect the differentiation of euglena plastids. Journal of Cell Biology, 15(3):599–??, December 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/15/3/599.

Goldberg:1964:ACS

[GG64] Burton Goldberg and Howard Green. An analysis of collagen secretion by established mouse fibroblast lines. Journal of Cell Biology, 22(1):227–??, July 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/22/1/227.

Glauert:1960:FSS

[GH60] Audrey M. Glauert and David A. Hopwood. The fine structure of Streptomyces coelicolor. Journal of Cell Biology, 7(3):479–??, June 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/7/
3/479.

Glauert:1961:FSS

[GH61] Audrey M. Glauert and David A. Hopwood. The fine structure of Streptomyces Violaceoruber (S. Coelicolor). Journal of Cell Biology, 10(4):505–??, August 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/10/4/505.

Gibbons:1961:RBF

[Gib61] I. R. Gibbons. The relationship between the fine structure and direction of beat in Gill cilia of a lamellibranch mollusc. Journal of Cell Biology, 11(1):179–??, October 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/11/1/179.

Gibbs:1962:CDO

[Gib62a] Sarah P. Gibbs. Chloroplast development in ochromonas danica. Journal of Cell Biology, 15(2):343–??, November 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/15/2/343.

Gibbs:1962:NEC

[Gib62b] Sarah P. Gibbs. Nuclear envelope-chloroplast relationships in algae. Journal of Cell Biology, 14(3):433–??, September 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/14/3/433.

Gibbons:1963:MOS

[Gib63] I. R. Gibbons. A method for obtaining serial sections of known orientation from single spermatozoa. Journal of Cell Biology, 16(3):626–??, March 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/16/3/626.

Gilev:1962:SMS

[Gil62] Vladimir P. Gilëv. A study of myofibril sarcomere structure during contraction. Journal of Cell Biology, 12(1):135–??, January 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/12/1/
135.

Gall:1960:TMD

[GJ60] Joseph G. Gall and William W. Johnson. Is there “Metabolic” DNA in the Mouse Seminal Vesicle? Journal of Cell Biology, 7(4):657–??, July 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/7/4/657.

Gonzales:1961:EMO

[GK61] Federico Gonzales and Morris J. Karnovsky. Electron microscopy of osteoclasts in healing fractures of rat bone. Journal of Cell Biology, 9(2):299–??, February 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/9/2/299.

Glauert:1963:FSM

[GKH63] Audrey M. Glauert, D. Kerridge, and R. W. Horne. The fine structure and mode of attachment of the sheathed flagellum of vibrio metchnikovii. Journal of Cell Biology, 18(2):327–??, August 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
18/2/327.

Godman:1964:SSC

[GL64] Gabriel C. Godman and Nathan Lane. On the site of sulfation in the chondrocyte. Journal of Cell Biology, 21(3):353–??, June 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/21/3/
353.

Greenblatt:1960:AME

[GOE60] C. L. Greenblatt, R. A. Olson, and E. K. Engel. Absorption Microscopy of Enzymatically Treated Cell–Free Chloroplasts. Journal of Cell Biology, 7(2):235–??, April 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/7/2/235.

Goldsmith:1962:FSR

[Gol62] Timothy H. Goldsmith. Fine structure of the retinulae in the compound eye of the honey-bee. Journal of Cell Biology, 14(3): 489–??, September 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/14/3/489.

Gonzales:1962:MTC

[Gon62] Federico Gonzales. A masking technique for contrast control in electron micrographs. Journal of Cell Biology, 15(1):146–??, October 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
15/1/146.

Godman:1960:CSE

[GP60] Gabriel C. Godman and Keith R. Porter. Chondrogenesis, studied with the electron microscope. Journal of Cell Biology, 8(3): 719–??, December 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/8/3/719.

Grillo:1963:CSC

[GP63] Mary A. Grillo and Sanford L. Palay. Ciliated Schwann cells in the autonomic nervous system of the adult rat. Journal of Cell Biology, 16(2):430–??, February 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/16/2/430.

Gross:1960:EMC

[GPN60] Paul R. Gross, Delbert E. Philpott, and Sylvan Nass. Electron Microscopy of the Centrifuged Sea Urchin Egg, with a Note on the Structure of the Ground Cytoplasm. Journal of Cell Biology, 7(1):135–??, February 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/7/1/135.

Gilmour:1964:CGM

[GR64] D. Gilmour and P. M. Robinson. Contraction in glycerinated myofibrils of an insect (orthoptera, acrididae). Journal of Cell Biology, 21(3):385–??, June 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/21/3/385.

Gray:1960:ROM

[Gra60] E. G. Gray. Regular organisation of material in certain mitochondria of neuroglia of lizard brain. Journal of Cell Biology, 8(1):282–??, September 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/8/1/282.

Green:1960:MGC

[Gre60] Paul B. Green. Multinet Growth in the Cell Wall of Nitella. Journal of Cell Biology, 7(2):289–??, April 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/7/2/289.

Gregoire:1961:SCC

[Gré61] Charles Grégoire. Structure of the conchiolin cases of the prisms in mytilus edulis linne. Journal of Cell Biology, 9(2): 395–??, February 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/9/2/395.

Griffin:1960:ICI

[Gri60] Joe L. Griffin. The Isolation, Characterization, and Identification of the Crystalline Inclusions of the Large Free–Living Amebae. Journal of Cell Biology, 7(2):227–??, April 1960. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/7/2/227.

Greenawalt:1964:EAA

[GRL64] John W. Greenawalt, Carlo S. Rossi, and Albert L. Lehninger. Effect of active accumulation of calcium and phosphate ions on the structure of rat liver mitochondria. Journal of Cell Biology, 23(1):21–??, October 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/23/1/21.

Grasso:1962:ODE

[GSA62] Joseph A. Grasso, Hewson Swift, and G. Adolph Ackerman. Observations on the development of erythrocytes in mammalian fetal liver. Journal of Cell Biology, 14(2):235–??, August 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/14/2/235.

Giacomelli:1964:SMR

[GSW64] Filiberto Giacomelli, David Spiro, and Joseph Wiener. A study of metastatic renal calcification at the cellular level. Journal of Cell Biology, 22(1):189–??, July 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/22/1/189.

Gotterer:1961:ALS

[GTL61] Gerald S. Gotterer, Thomas E. Thompson, and Albert L. Lehninger. Angular light-scattering studies on isolated mitochondria. Journal of Cell Biology, 10(1):15–??, May 1961. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/10/1/15.

Granboulan:1963:EMS

[GTWB63] Nicole Granboulan, P. Tournier, R. Wicker, and W. Bernhard. An electron microscope study of the development of Sv40 virus. Journal of Cell Biology, 17(2):423–??, May 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/17/2/423.

Gray:1964:EMS

[GY64] E. G. Gray and J. Z. Young. Electron microscopy of synaptic structure of octopus brain. Journal of Cell Biology, 21(1):87–??, April 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
21/1/87.

Howatson:1960:OFS

[HA60a] A. F. Howatson and J. D. Almeida. Observations on the fine structure of polyoma virus. Journal of Cell Biology, 8(3): 828–??, December 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/8/3/828.

Howatson:1960:EMS

[HA60b] Allan F. Howatson and June D. Almeida. An Electron Microscope Study of Polyoma Virus in Hamster Kidney. Journal of Cell Biology, 7(4):753–??, July 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/7/4/753.

Haggis:1961:EMR

[Hag61] G. H. Haggis. Electron microscope replicas from the surface of a fracture through frozen cells. Journal of Cell Biology, 9 (4):841–??, April 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/9/4/841.

Hall:1960:MGP

[Hal60] Cecil E. Hall. Measurement of Globular Protein Molecules by Electron Microscopy. Journal of Cell Biology, 7(4):613–??, July 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/7/
4/613.

Hale:1961:HIA

[Hal61] Arthur J. Hale. Historadiographic identification of alkaline phosphatase. Journal of Cell Biology, 11(2):488–??, November 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/11/2/
488.

Hale:1962:ICR

[Hal62a] Arthur J. Hale. Identification of cytochemical reaction products by scanning X-ray emission microanalysis. Journal of Cell Biology, 15(3):427–??, December 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/15/3/427.

Hale:1962:ISC

[Hal62b] Arthur J. Hale. Identification of some cytochemical reaction products by historadiography. Journal of Cell Biology, 15(3): 417–??, December 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/15/3/417.

Hama:1960:FSS

[Ham60a] Kiyoshi Hama. The Fine Structure of Some Blood Vessels of the Earthworm, Eisenia foetida. Journal of Cell Biology, 7(4):717– ??, July 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/7/
4/717.

Hama:1960:FSD

[Ham60b] Kiyoshi Hama. The Fine Structure of the Desmosomes in Frog Mesothelium. Journal of Cell Biology, 7(3):575–??, June 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/7/3/575.

Hannay:1961:FBP

[Han61] Christopher L. Hannay. Fowler’s bacillus and its parasporal body. Journal of Cell Biology, 9(2):285–??, February 1961. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/9/2/285.

Harris:1961:SSM

[Har61a] E. J. Harris. The site of swelling in muscle. Journal of Cell Biology, 9(2):502–??, February 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/9/2/502.

Harris:1961:EMS

[Har61b] Patricia Harris. Electron microscope study of mitosis in sea urchin blastomeres. Journal of Cell Biology, 11(2):419– ??, November 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/11/2/419.

Harris:1962:SSF

[Har62] Patricia Harris. Some structural and functional aspects of the mitotic apparatus in sea urchin embryos. Journal of Cell Biology, 14(3):475–??, September 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/14/3/475.

Hartmann:1964:CLA

[Har64] John F. Hartmann. Cytochemical localization of adenosine triphosphatase in the mitotic apparatus of HeLa and sarcoma 180 tissue culture cells. Journal of Cell Biology, 23(2): 363–??, November 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/23/2/363.

Hay:1961:FSU

[Hay61] Elizabeth D. Hay. Fine structure of an unusual intracellular supporting network in the leydig cells of amblystoma epidermis. Journal of Cell Biology, 10(3):457–??, July 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/10/3/457.

Hancox:1961:MPE

[HB61] N. M. Hancox and B. Boothroyd. Motion picture and electron microscope studies on the embryonic avian osteoclast. Journal of Cell Biology, 11(3):651–??, December 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/11/3/651.

Hall:1961:FSD

[HC61] Cecil E. Hall and Liebe F. Cavalieri. Four-stranded DNA as determined by electron microscopy. Journal of Cell Biology, 10(3):347–??, July 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/10/3/347.

Hall:1963:USB

[HC63] William T. Hall and George Claus. Ultrastructural studies on the blue-green algal symbiont in cyanophora paradoxa korschikoff. Journal of Cell Biology, 19(3):551–??, December 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/19/3/551.

Hess:1960:CLD

[HD60] R. Hess and F. M. Dietrich. Cytochemical localization of dehydrogenase and diaphorase activity in bacillus megaterium. Journal of Cell Biology, 8(2):546–??, October 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/8/2/546.

Hyden:1962:CBC

[HE62] Holger Hydén and Endre Egyházi. Changes in the base composition of nuclear ribonucleic acid of neurons during a short period of enhanced protein production. Journal of Cell Biology, 15(1):37–??, October 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/15/1/37.

Heide:1962:EMO

[Hei62] Hans Gunther Heide. Electron microscopic observation of specimens under controlled gas pressure. Journal of Cell Biology, 13(1):147–??, April 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/13/1/147.

Herman:1960:EMS

[Her60] Lawrence Herman. An Electron Microscope Study of the Salamander Thyroid during Hormonal Stimulation. Journal of Cell Biology, 7(1):143–??, February 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/7/1/143.

Herndon:1963:FSP

[Her63] Robert M. Herndon. The fine structure of the Purkinje cell. Journal of Cell Biology, 18(1):167–??, July 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/18/1/167.

Herndon:1964:FSR

[Her64a] Robert M. Herndon. The fine structure of the rat cerebellum. Journal of Cell Biology, 23(2):277–??, November 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/23/2/277.

Herndon:1964:LBU

[Her64b] Robert M. Herndon. Lamellar bodies, an unusual arrangement of the granular endoplasmic reticulum. Journal of Cell Biology, 20(2):338–??, February 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/20/2/338.

Hess:1960:FSD

[Hes60] Arthur Hess. The fine structure of degenerating nerve fibers, their sheaths, and their terminations in the central nerve cord of the cockroach (Periplaneta americana). Journal of Cell Biology, 7(2):339–??, April 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/7/2/339.

Herman:1962:DCP

[HF62a] Lawrence Herman and Patrick J. Fitzgerald. The degenerative changes in pancreatic acinar cells caused by dl-ethionine. Journal of Cell Biology, 12(2):277–??, February 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/12/2/277.

Herman:1962:RPA

[HF62b] Lawrence Herman and Patrick J. Fitzgerald. Restitution of pancreatic acinar cells following ethionine. Journal of Cell Biology, 12(2):297–??, February 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/12/2/297.

Hashimoto:1960:MUM

[HG60a] Tadayo Hashimoto and Philipp Gerhardt. Monochromatic Ultraviolet Microscopy of Microorganisms: Preliminary Observations on Bacterial Spores. Journal of Cell Biology, 7(1): 195–??, February 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/7/1/195.

Hopwood:1960:FSS

[HG60b] David A. Hopwood and Audrey M. Glauert. The fine structure of Streptomyces coelicolor. Journal of Cell Biology, 8(1): 267–??, September 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/8/1/267.

Hopwood:1960:FSN

[HG60c] David A. Hopwood and Audrey M. Glauert. The fine structure of the nuclear material of a blue-green alga, anabaena cylindrica lemm. Journal of Cell Biology, 8(3):813–??, December 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/8/3/813.

Hopwood:1960:OCB

[HG60d] David A. Hopwood and Audrey M. Glauert. Observations on the chromatinic bodies of Streptomyces coelicolor. Journal of Cell Biology, 8(1):257–??, September 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/8/1/257.

Hopsu:1963:HEK

[HG63] Väinö K. Hopsu and George G. Glenner. A histochemical enzyme kinetic system applied to the trypsin-like amidase and esterase activity in human mast cells. Journal of Cell Biology, 17(3):503–??, June 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/17/3/503.

Hashimoto:1960:SFS

[HGCN60] Tadayo Hashimoto, Philipp Gerhardt, S. F. Conti, and H. B. Naylor. Studies on the Fine Structure of Microorganisms. Journal of Cell Biology, 7(2):305–??, April 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/7/2/305.

Holt:1961:LAP

[HH61a] S. J. Holt and R. Marian Hicks. The localization of acid phosphatase in rat liver cells as revealed by combined cytochemical staining and electron microscopy. Journal of Cell Biology, 11 (1):47–??, October 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/11/1/47.

Holt:1961:SFF

[HH61b] S. J. Holt and R. Marian Hicks. Studies on formalin fixation for electron microscopy and cytochemical staining purposes. Journal of Cell Biology, 11(1):31–??, October 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/11/1/31.

Hamberger:1963:IEC

[HH63] Anders Hamberger and Holger Hydén. Inverse enzymatic changes in neurons and glia during increased function and hypoxia. Journal of Cell Biology, 16(3):521–??, March 1963. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/16/3/521.

Hamashima:1964:LAF

[HHC64] Yoshihiro Hamashima, John G. Harter, and Albert H. Coons. The localization of albumin and fibrinogen in human liver cells. Journal of Cell Biology, 20(2):271–??, February 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/20/2/271.

Holt:1960:PIR

[HHP60] S. J. Holt, Eluned E. Hobbiger, and G. L. S. Pawan. Preservation of Integrity of Rat Tissues for Cytochemical Staining Purposes. Journal of Cell Biology, 7(2):383–??, April 1960. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/7/2/383.

Hill:1964:LAN

[Hil64] D. K. Hill. The location of adenine nucleotide in the striated muscle of the toad. Journal of Cell Biology, 20(3):435–??, March 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/20/3/
435.

Hinegardner:1962:INE

[Hin62] Ralph T. Hinegardner. The isolation of nuclei from eggs and embryos of the sea urchin. Journal of Cell Biology, 15(3): 503–??, December 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/15/3/503.

Hiraoka:1960:EAC

[Hir60] Tosisuke Hiraoka. Electrophoretic analysis of colored product produced by Schiff-formaldehyde reaction. Journal of Cell Biology, 8(1):286–??, September 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/8/1/286.

Himmelhoch:1961:HDG

[HK61a] S. Ralph Himmelhoch and Morris J. Karnovsky. The histochemical demonstration of glyceraldehyde-3-phosphate dehydrogenase activity. Journal of Cell Biology, 9(3):573–??, March 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/9/3/573.

Himmelhoch:1961:OHE

[HK61b] S. Ralph Himmelhoch and Morris J. Karnovsky. Oxidative and hydrolytic enzymes in the nephron of necturus maculosus. Journal of Cell Biology, 9(4):893–??, April 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/9/4/893.

Herrmann:1964:VCM

[HKA64] Heinz Herrmann, N. W. Klein, and G. Albers. In vitro culture of muscle tissue from normal, heterozygous, and dystrophic chicks. Journal of Cell Biology, 22(2):391–??, August 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/22/2/391.

Herrmann:1962:ALC

[HL62a] Heinz Herrmann and Patricia L. Lebeau. ATP level, cell injury, and apparent epithelium-stroma interaction in the cornea. Journal of Cell Biology, 13(3):465–??, June 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/13/3/465.

Hyden:1962:KSN

[HL62b] Holger Hydén and Paul W. Lange. A kinetic study of the neuron-glia relationship. Journal of Cell Biology, 13(2):233–??, May 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
13/2/233.

Hayes:1963:QDO

[HLG63] Thomas L. Hayes, Frank T. Lindgren, and John W. Gofman. A quantitative determination of the osmium tetroxide-lipoprotein interaction. Journal of Cell Biology, 19(1):251–??, October 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/19/1/
251.

Hymer:1962:ICP

[HM62] W. C. Hymer and W. H. McShan. Isolation of cytoplasmic pituitary granules by column chromatography. Journal of Cell Biology, 13(2):350–??, May 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/13/2/350.

Hess:1963:CDL

[HM63a] Oswald Hess and Günther F. Meyer. Chromosomal differentiations of the lampbrush type formed by the Y chromosome in Drosophila Hydei and Drosophila Neohydei. Journal of Cell Biology, 16(3):527–??, March 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/16/3/527.

Hymer:1963:IRP

[HM63b] W. C. Hymer and W. H. McShan. Isolation of rat pituitary granules and the study of their biochemical properties and hormonal activities. Journal of Cell Biology, 17(1):67–??, April 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/17/1/
67.

Hartley:1960:ICP

[HMR60] Marshall W. Hartley, W. H. McShan, and Hans Ris. Isolation of Cytoplasmic Pituitary Granules with Gonadotropic Activity. Journal of Cell Biology, 7(2):209–??, April 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/7/2/209.

Hepler:1964:MFC

[HN64] Peter K. Hepler and Eldon H. Newcomb. Microtubules and fibrils in the cytoplasm of coleus cells undergoing secondary wall deposition. Journal of Cell Biology, 20(3):529–??, March 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/20/3/
529.

Holbert:1960:EMP

[Hol60] Pauline E. Holbert. An Effective Method of Preparing Sections of Bacillus polymyxa Sporangia and Spores for Electron Microscopy. Journal of Cell Biology, 7(2):373–??, April 1960. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/7/2/373.

Huff:1960:MBL

[HPWJ60] Clay G. Huff, Alan C. Pipkin, A. B. Weathersby, and Dinniemaud V. Jensen. The Morphology and Behavior of Living Exoerythrocytic Stages of Plasmodium gallinaceum and P. fallax and Their Host Cells. Journal of Cell Biology, 7(1): 93–??, February 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/7/1/93.

Hampton:1961:CAE

[HQ61] James C. Hampton and Henry Quastler. Combined autoradiography and electron microscopy of thin sections of intestinal epithelial cells of the mouse labeled with H^{3}-thymidine. Journal of Cell Biology, 10(1):140–??, May 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/10/1/140.

Horst:1962:EMD

[HR62] Antoni Horst and Tadeusz Rudnicki. The effect of a medium dose (430 roentgens) of X-ray irradiation on resting cells of the liver. Journal of Cell Biology, 13(2):261–??, May 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/13/2/261.

Hay:1963:FSD

[HR63] Elizabeth D. Hay and J. P. Revel. The fine structure of the DNP component of the nucleus. Journal of Cell Biology, 16(1):29–??, January 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
16/1/29.

Hadek:1960:CIR

[HS60] Robert Hadek and Hewson Swift. A crystalloid inclusion in the rabbit blastocyst. Journal of Cell Biology, 8(3):836–??, December 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/8/
3/836.

Hotta:1961:DDM

[HS61a] Yasuo Hotta and Herbert Stern. Deamination of deoxycytidine and 5-methyldeoxycytidine in developing anthers of lilium longiflorum (var. croft). Journal of Cell Biology, 9(2):279–??, February 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
9/2/279.

Hotta:1961:TPD

[HS61b] Yasuo Hotta and Herbert Stern. Transient phosphorylation of deoxyribosides and regulation of deoxyribonucleic acid synthesis. Journal of Cell Biology, 11(2):311–??, November 1961. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/11/2/311.

Hadek:1962:NEI

[HS62] Robert Hadek and Hewson Swift. Nuclear extrusion and intracisternal inclusions in the rabbit blastocyst. Journal of Cell Biology, 13(3):445–??, June 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/13/3/445.

Hotta:1963:IPS

[HS63a] Yasuo Hotta and Herbert Stern. Inhibition of protein syntheses during meiosis and its bearing on intracellular regulation. Journal of Cell Biology, 16(2):259–??, February 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/16/2/259.

Hotta:1963:SML

[HS63b] Yasuo Hotta and Herbert Stern. Synthesis of messenger-like ribonucleic acid and protein during meiosis in isolated cells of trillium erectum. Journal of Cell Biology, 19(1):45–??, October 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/19/1/
45.

Hsu:1964:MCV

[Hsu64] T. C. Hsu. Mammalian chromosomes in vitro. Journal of Cell Biology, 23(1):53–??, October 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/23/1/53.

Hsu:1964:LER

[HT64] Lichu Hsu and A. L. Tappel. Lysosomal enzymes of rat intestinal mucosa. Journal of Cell Biology, 23(2):233–??, November 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/23/2/
233.

Humphreys:1963:PCF

[Hum63] Walter J. Humphreys. Preparation of carbon films by use of a glycerin substrate. Journal of Cell Biology, 19(3):634–??, December 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
19/3/634.

Holmes:1961:SFE

[HW61] Richard Holmes and Sonya W. Wolfe. Serum fractionation and the effects of bovine serum fractions on human cells grown in a chemically defined medium. Journal of Cell Biology, 10(3):389– ??, July 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
10/3/389.

Huxley:1961:PSN

[HZ61] H. E. Huxley and G. Zubay. Preferential staining of nucleic acid-containing structures for electron microscopy. Journal of Cell Biology, 11(2):273–??, November 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/11/2/273.

Hendee:1963:ASC

[HZB63] William R. Hendee, William Zebrun, and Frederick J. Bonte. Anomalous structures in the cytoplasm of HeLa cells cultured in the presence of 5-bromodeoxyuridine. Journal of Cell Biology, 17(3):675–??, June 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/17/3/675.

Inman:1963:EMH

[IC63] D. R. Inman and E. H. Cooper. Electron microscopy of human lymphocytes stimulated by phytohaemagglutinin. Journal of Cell Biology, 19(2):441–??, November 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/19/2/441.

Ito:1960:LSC

[Ito60] Susumu Ito. The lamellar systems of cytoplasmic membranes in dividing spermatogenic cells of Drosophila virilis. Journal of Cell Biology, 7(3):433–??, June 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/7/3/433.

Ito:1961:ERG

[Ito61] Susumu Ito. The endoplasmic reticulum of gastric parietal cells. Journal of Cell Biology, 11(2):333–??, November 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/11/2/333.

Iverson:1964:CRB

[Ive64] Ray M. Iverson. The content and relative base ratios of ribonucleic acid in Amoeba. Journal of Cell Biology, 20(1):1–??, January 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
20/1/1.

Ito:1963:FSG

[IW63] Susumu Ito and Robert J. Winchester. The fine structure of the gastric mucosa in the bat. Journal of Cell Biology, 16(3):541–??, March 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
16/3/541.

Jackson:1964:CAC

[Jac64] John F. Jackson. Chromosome aberrations in cultured human leukocytes treated with 8-ethoxycaffeine. Journal of Cell Biology, 22(1):291–??, July 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/22/1/291.

Jackson:1960:SEC

[JB60] David S. Jackson and John P. Bentley. On the Significance of the Extractable Collagens. Journal of Cell Biology, 7(1): 37–??, February 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/7/1/37.

Jaffe:1962:OLT

[JE62] Lionel Jaffe and Helmut Etzold. Orientation and locus of tropic photoreceptor molecules in spores of botrytis and osmunda. Journal of Cell Biology, 13(1):13–??, April 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/13/1/13.

Jones:1964:EAD

[JE64] K. W. Jones and T. R. Elsdale. The effects of actinomycin D on the ultrastructure of the nucleus of the amphibian embryonic cell. Journal of Cell Biology, 21(2):245–??, May 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/21/2/245.

Jensen:1964:OFN

[Jen64] William A. Jensen. Observations on the fusion of nuclei in plants. Journal of Cell Biology, 23(3):669–??, December 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/23/3/669.

Jansen:1961:UMP

[JM61] M. T. Jansen and I. Molenaar. Ultraviolet microscopy of pancreatic acinar nuclei fixed in potassium permanganate. Journal of Cell Biology, 9(3):716–??, March 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/9/3/716.

Johansen:1960:EMO

[JP60] Erling Johansen and Harold F. Parks. Electron Microscopic Observations on the Three–Dimensional Morphology of Apatite Crystallites of Human Dentine and Bone’. Journal of Cell Biology, 7(4):743–??, July 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/7/4/743.

Jamieson:1964:SGA

[JP64] J. D. Jamieson and G. E. Palade. Specific granules in atrial muscle cells. Journal of Cell Biology, 23(1):151–??, October 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/23/1/
151.

Jacob:1963:EMS

[JS63] J. Jacob and J. L. Sirlin. Electron microscope studies on salivary gland cells. Journal of Cell Biology, 17(1):153–??, April 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/17/1/
153.

Jacobson:1963:CSC

[JSB63] Ann B. Jacobson, Hewson Swift, and Lawrence Bogorad. Cytochemical studies concerning the occurrence and distribution of RNA in plastids of Zea Mays. Journal of Cell Biology, 17 (3):557–??, June 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/17/3/557.

Kane:1960:EPP

[Kan60] R. E. Kane. The Effect of Partial Protein Extraction on the Structure of the Eggs of the Sea Urchin, Arbacia punctulata. Journal of Cell Biology, 7(1):21–??, February 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/7/1/21.

Kane:1962:MA

[Kan62a] R. E. Kane. The mitotic apparatus. Journal of Cell Biology, 15(2):279–??, November 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/15/2/279.

Kane:1962:MAI

[Kan62b] R. E. Kane. The mitotic apparatus: isolation by controlled pH. Journal of Cell Biology, 12(1):47–??, January 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/12/1/47.

Karrer:1960:CIN

[Kar60a] H. E. Karrer. Cell Interconnections in Normal Human Cervical Epithelium. Journal of Cell Biology, 7(1):181–??, February 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
7/1/181.

Karrer:1960:EMS

[Kar60b] H. E. Karrer. Electron Microscopic Study of the Phagocytosis Process in Lung. Journal of Cell Biology, 7(2):357–??, April 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/7/
2/357.

Karrer:1960:SMB

[Kar60c] H. E. Karrer. The striated musculature of blood vessels. Journal of Cell Biology, 8(1):135–??, September 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/8/1/135.

Karnovsky:1961:SMS

[Kar61] Morris J. Karnovsky. Simple methods for “staining with lead” at high pH in electron microscopy. Journal of Cell Biology, 11 (3):729–??, December 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/11/3/729.

Karasaki:1963:SAY

[Kar63] Shuichi Karasaki. Studies on amphibian yolk. Journal of Cell Biology, 18(1):135–??, July 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/18/1/135.

Karnovsky:1964:LCA

[Kar64] Morris J. Karnovsky. The localization of cholinesterase activity in rat cardiac muscle by electron microscopy. Journal of Cell Biology, 23(2):217–??, November 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/23/2/217.

Kaye:1962:SCc

[Kay62a] Gordon I. Kaye. Studies on the cornea. Journal of Cell Biology, 15(2):241–??, November 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/15/2/241.

Kaye:1962:AFH

[Kay62b] Jerome S. Kaye. Acrosome formation in the house Cricket. Journal of Cell Biology, 12(2):411–??, February 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/12/2/411.

Kaye:1964:FSF

[Kay64] Jerome S. Kaye. The fine structure of flagella in spermatids of the house Cricket. Journal of Cell Biology, 22(3):710– ??, September 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/22/3/710.

Karnofsky:1960:EFR

[KB60] David A. Karnofsky and Ross S. Basch. Effects of 5-Fluorodeoxyuridine and Related Halogenated Pyrimidines on the Sand–Dollar Embryo. Journal of Cell Biology, 7(1):61– ??, February 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/7/1/61.

Klein:1964:FSC

[KB64] Shimon Klein and Lawrence Bogorad. Fine structural changes in proplastids during photodestruction of pigments. Journal of Cell Biology, 22(2):443–??, August 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/22/2/443.

Klein:1964:ESD

[KBB64] Shimon Klein, G. Bryan, and Lawrence Bogorad. Early stages in the development of plastid fine structure in red and far-red light. Journal of Cell Biology, 22(2):433–??, August 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/22/2/433.

Knappeis:1962:UZD

[KC62] G. G. Knappeis and F. Carlsen. The ultrastructure of the Z disc in skeletal muscle. Journal of Cell Biology, 13(2):323–??, May 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/13/2/
323.

Keck:1963:ACS

[KC63] Konrad Keck and Elizabeth A. Choules. An analysis of cellular and subcellular systems which transform the species character of acid phosphatase in acetabularia. Journal of Cell Biology, 18(2):459–??, August 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/18/2/459.

Keech:1960:EMSb

[Kee60a] M. K. Keech. Electron Microscope Study of the Lathyritic Rat Aorta. Journal of Cell Biology, 7(3):539–??, June 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/7/3/539.

Keech:1960:EMSa

[Kee60b] M. K. Keech. Electron Microscope Study of the Normal Rat Aorta. Journal of Cell Biology, 7(3):533–??, June 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/7/3/533.

Keech:1961:FFC

[Kee61] M. K. Keech. The formation of fibrils from collagen solutions. Journal of Cell Biology, 9(1):193–??, January 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/9/1/193.

Kessel:1962:FSP

[Kes62] Richard G. Kessel. Fine structure of pigment inclusions in the test cells of the ovary of styela. Journal of Cell Biology, 12 (3):637–??, March 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/12/3/637.

Kessel:1963:EMS

[Kes63] Richard G. Kessel. Electron microscope studies on the origin of annulate lamellae in oocytes of necturus. Journal of Cell Biology, 19(2):391–??, November 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/19/2/391.

Karnovsky:1960:HMD

[KF60a] Morris J. Karnovsky and Gerald D. Fasman. A histochemical method for distinguishing between side-chain and terminal (α-acylamido) carboxyl groups of proteins. Journal of Cell Biology, 8(2):319–??, October 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/8/2/319.

King:1960:VIU

[KF60b] R. C. King and Gretchen J. Falk. In vitro incorporation of uridine-H^{3} into developing fruit fly oocytes. Journal of Cell Biology, 8(2):550–??, October 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/8/2/550.

Klein:1960:EMI

[KG60] Shimon Klein and B. Ginzburg. An electron microscopic investigation into the effect of EDTA on plant cell wall. Journal of Cell Biology, 7(2):335–??, April 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/7/2/335.

Kallman:1964:FSD

[KG64] Frances Kallman and Clifford Grobstein. Fine structure of differentiating mouse pancreatic exocrine cells in transfilter culture. Journal of Cell Biology, 20(3):399–??, March 1964. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/20/3/399.

Karnovsky:1963:NMB

[KH63] Morris J. Karnovsky and Katharina Hug. The nature of the M band enzyme in rat ventricular muscle. Journal of Cell Biology, 19(1):255–??, October 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/19/1/255.

Kojima:1962:SCH

[KK62a] Kiyohide Kojima and Sadao Kozuka. Structural changes in HeLa cells cultivated in serum-free medium. Journal of Cell Biology, 14(1):141–??, July 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/14/1/141.

Kroll:1962:PPI

[KK62b] Arnold J. Kroll and Toichiro Kuwabara. Prevention of phosphate-induced mitochondrial swelling. Journal of Cell Biology, 15(1):29–??, October 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/15/1/29.

Koulish:1964:RN

[KK64] Sasha Koulish and Ruth G. Kleinfeld. The role of the nucleolus. Journal of Cell Biology, 23(1):39–??, October 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/23/1/39.

Kishimoto:1964:NMC

[KL64] Susumu Kishimoto and Irving Lieberman. A nuclear membrane change after partial hepatectomy. Journal of Cell Biology, 23(3):511–??, December 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/23/3/511.

Klein:1960:ELT

[Kle60] Shimon Klein. The effect of low temperature on the development of the lamellar system in chloroplasts. Journal of Cell Biology, 8(2):529–??, October 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/8/2/529.

Kluss:1962:EMM

[Klu62] B. C. Kluss. Electron microscopy of the macronucleus of euplotes eurystomus. Journal of Cell Biology, 13(3):462–??, June 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/13/3/
462.

Koehler:1963:EMA

[KMF63] J. K. Koehler, K. Mühlethaler, and A. Freywyssling. Electron microscopic autoradiography. Journal of Cell Biology, 16 (1):73–??, January 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/16/1/73.

Konigsberg:1960:DDA

[KMTH60] Irwin R. Konigsberg, Norma McElvain, Martha Tootle, and Heinz Herrmann. The dissociability of deoxyribonucleic acid synthesis from the development of multinuclearity of muscle cells in culture. Journal of Cell Biology, 8(2):333–??, October 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
8/2/333.

Konrad:1963:PSR

[Kon63] Carol G. Konrad. Protein synthesis and RNA synthesis during mitosis in animal cells. Journal of Cell Biology, 19(2): 267–??, November 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/19/2/267.

Kaye:1962:SCa

[KP62] Gordon I. Kaye and George D. Pappas. Studies on the cornea. Journal of Cell Biology, 12(3):457–??, March 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/12/3/457.

Kimball:1964:RPS

[KP64] R. F. Kimball and D. M. Prescott. RNA and protein synthesis in amacronucleate Paramecium Aurelia. Journal of Cell Biology, 21(3):496–??, June 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/21/3/496.

Kaye:1962:SCb

[KPDM62] Gordon I. Kaye, George D. Pappas, Anthony Donn, and Nancy Mallett. Studies on the cornea. Journal of Cell Biology, 12 (3):481–??, March 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/12/3/481.

Klein:1961:FSP

[KPM61] Shimon Klein and A. Poljakoff-Mayber. Fine structure and pigment conversion in isolated etiolated proplastids. Journal of Cell Biology, 11(2):433–??, November 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/11/2/433.

Klinger:1960:SCH

[KS60] Harold P. Klinger and Hans G. Schwarzacher. The sex chromatin and heterochromatic bodies in human diploid and polyploid nuclei. Journal of Cell Biology, 8(2):345–??, October 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/8/2/345.

Kelly:1964:FSP

[KS64] Douglas E. Kelly and Stuart W. Smith. Fine structure of the pineal organs of the adult frog, rana pipiens. Journal of Cell Biology, 22(3):653–??, September 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/22/3/653.

Kallenbach:1963:GAA

[KSW63] E. Kallenbach, E. Sandborn, and H. Warshawsky. The Golgi apparatus of the ameloblast of the rat at the stage of enamel matrix formation. Journal of Cell Biology, 16(3):629–??, March 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/16/3/
629.

Koike:1961:FSI

[KT61] Masaatsu Koike and Kenji Takeya. Fine structures of intracytoplasmic organelles of mycobacteria. Journal of Cell Biology, 9 (3):597–??, March 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/9/3/597.

Kodama:1963:EMS

[KT63] Robert M. Kodama and Henry Tedeschi. An electron microscope study of calf thymus nuclear preparations isolated in sucrose solutions. Journal of Cell Biology, 18(3):541–??, September 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/18/3/
541.

Keiser:1964:SL

[KWB64] Harold Keiser, Gerald Weissmann, and Alan W. Bernheimer. Studies on lysosomes. Journal of Cell Biology, 22(1):101–??, July 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
22/1/101.

Kamat:1960:ILH

[KWCL60] Virendra B. Kamat, Donald F. Hoelzl Wallach, John F. Crigler, and Aaron J. Ladman. The Intracellular Localization of Hormonal Activity in Transplantable Thyrotropin–Secreting Pituitary Tumors in Mice. Journal of Cell Biology, 7(2):219–??, April 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/7/
2/219.

Kuff:1960:CRC

[KZ60] Edward L. Kuff and Robert F. Zeigel. Cytoplasmic Ribonucleoprotein Components of the Novikoff Hepatoma. Journal of Cell Biology, 7(3):465–??, June 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/7/3/465.

Loran:1960:CPI

[LA60] M. R. Loran and T. L. Althausen. Cellular Proliferation of Intestinal Epithelia in the Rat Two Months after Partial Resection of the Ileum. Journal of Cell Biology, 7(4):667–??, July 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/7/4/667.

Lafontaine:1964:LEM

[LA64] J. G. Lafontaine and C. Allard. A light and electron microscope study of the morphological changes induced in rat liver cells by the azo dye 2-me-dab. Journal of Cell Biology, 22(1):143–??, July 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
22/1/143.

Lallier:1962:EDT

[Lal62] R. Lallier. Les effets du thiosorbitol et du ThioglycÉRol sur la structure De l’appareil mitotique De l’oeuf De l’oursin paracentrotus lividus. Journal of Cell Biology, 15(2):382–??, November 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/15/2/
382.

Lang:1963:AUC

[Lan63] Norma J. Lang. An additional ultrastructural component of flagella. Journal of Cell Biology, 19(3):631–??, December 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/19/3/631.

Lansing:1964:AVC

[Lan64] Albert I. Lansing. Age variations in cortical membranes of rotifers. Journal of Cell Biology, 23(3):403–??, December 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/23/3/403.

Larsen:1962:EMU

[Lar62] Jørgen Falck Larsen. Electron microscopy of the uterine epithelium in the rabbit. Journal of Cell Biology, 14(1):49–??, July 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/14/1/
49.

Lasansky:1961:MBN

[Las61] Arnaldo Lasansky. Morphological bases for a nursing role of glia in the toad retina. electron microscope observations. Journal of Cell Biology, 11(1):237–??, October 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/11/1/237.

Lavallard:1960:EAM

[Lav60] Par Roger Lavallard. Étude au microscope Électronique du reticulum endoplasmique dans des fibres musculaires du crabe bleu. (French) [Electroscopic study of endoplasmic reticulum in muscle fibers of the blue crab]. Journal of Cell Biology, 7 (2):399–??, April 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/7/2/399.

Lawn:1960:OFS

[Law60a] A. M. Lawn. Observations on the Fine Structure of the Gastric Parietal Cell of the Rat. Journal of Cell Biology, 7(1): 161–??, February 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/7/1/161.

Lawn:1960:UPP

[Law60b] A. M. Lawn. The Use of Potassium Permanganate as an Electron–Dense Stain for Sections of Tissue Embedded in Epoxy Resin. Journal of Cell Biology, 7(1):197–??, February 1960. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/7/1/197.

Leduc:1961:UC

[LB61] Elizabeth H. Leduc and Wilhelm Bernhard. Ultrastructural cytochemistry. Journal of Cell Biology, 10(3):437–??, July 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/10/3/437.

Longley:1960:SRM

[LBB60] J. B. Longley, W. G. Banfield, and D. C. Brindley. Structure of the Rete Mirabile in the Kidney of the Rat as Seen with the Electron Microscope. Journal of Cell Biology, 7(1): 103–??, February 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/7/1/103.

Lindall:1963:IIS

[LBDL63] Arnold W. Lindall, G. Eric Bauer, Padmaker K. Dixit, and Arnold Lazarow. Isolation of an insulin secretion granule fraction. Journal of Cell Biology, 19(2):317–??, November 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/19/2/317.

Low:1962:PME

[LC62] Frank N. Low and Max R. Clevenger. Polyester-methacrylate embedments for electron microscopy. Journal of Cell Biology, 12(3):615–??, March 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/12/3/615.

Lafontaine:1963:CLE

[LC63a] J. G. Lafontaine and L. A. Chouinard. A correlated light and electron microscope study of the nucleolar material during mitosis in Vicia Faba. Journal of Cell Biology, 17(1):167–??, April 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/17/1/
167.

Loran:1963:PDI

[LC63b] M. R. Loran and T. T. Crocker. Population dynamics of intestinal epithelia in the rat two months after partial resection of the ileum. Journal of Cell Biology, 19(2):285–??, November 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/19/2/
285.

Lane:1964:SSC

[LCOVG64] Nathan Lane, Lucien Caro, Luis R. Otero-Vilardebó, and Gabriel C. Godman. On the site of sulfation in colonic goblet cells. Journal of Cell Biology, 21(3):339–??, June 1964. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/21/3/339.

Lasansky:1960:SAG

[LD60] Arnaldo Lasansky and Eduardo De Robertis. Submicroscopic Analysis of the Genetic Distrophy of Visual Cells in C3H Mice. Journal of Cell Biology, 7(4):679–??, July 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/7/4/679.

Lust:1963:ETL

[LD63] J. Lust and P. Drochmans. Effect of trypsin on liver microsomes. Journal of Cell Biology, 16(1):81–??, January 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/16/1/81.

Lasansky:1960:EMR

[LdR60] Arnaldo Lasansky and Eduardo de Robertis. Electron Microscopy of Retinal Photoreceptors. Journal of Cell Biology, 7(3):493–??, June 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/7/3/493.

Leeson:1962:ADC

[Lee62] C. Roland Leeson. The application of diazotype color films to the production of projection slides of electron microscopic material. Journal of Cell Biology, 15(2):394–??, November 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/15/2/394.

Lee:1964:VAS

[LF64] Robert E. Lee and Joseph D. Feldman. Visualization of antigenic sites of human erythrocytes with ferritin-antibody conjugates. Journal of Cell Biology, 23(2):396–??, November 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/23/2/396.

Lynn:1964:OMA

[LFB64a] William S. Lynn, Sydney Fortney, and Rose H. Brown. Osmotic and metabolic alterations of mitochondrial size. Journal of Cell Biology, 23(1):1–??, October 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/23/1/1.

Lynn:1964:REM

[LFB64b] William S. Lynn, Sydney Fortney, and Rose H. Brown. Role of EDTA and metals in mitochondrial contraction. Journal of Cell Biology, 23(1):9–??, October 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/23/1/9.

Lowther:1961:MCS

[LGC61] D. A. Lowther, N. M. Green, and J. A. Chapman. Morphological and chemical studies of collagen formation. Journal of Cell Biology, 10(3):373–??, July 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/10/3/373.

Luzzati:1962:SLC

[LH62] V. Luzzati and F. Husson. The structure of the liquid-crystalline phases of lipid-water systems. Journal of Cell Biology, 12(2): 207–??, February 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/12/2/207.

Lillibridge:1961:MHP

[Lil61] Clinton B. Lillibridge. Membranes of the human pepsinogen granule. Journal of Cell Biology, 10(1):145–??, May 1961. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/10/1/145.

Litt:1964:SEE

[Lit64] Mortimer Litt. Studies in experimental eosinophilia. Journal of Cell Biology, 23(2):355–??, November 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/23/2/355.

Leeson:1961:EVD

[LK61] T. S. Leeson and H. Kalant. Effects of in vivo decalcification on ultrastructure of adult rat liver. Journal of Cell Biology, 10(1):95–??, May 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/10/1/95.

Loewenstein:1963:ECP

[LK63] Werner R. Loewenstein and Yoshinobu Kanno. The electrical conductance and potential across the membrane of some cell nuclei. Journal of Cell Biology, 16(2):421–??, February 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/16/2/421.

Loewenstein:1964:SEG

[LK64] Werner R. Loewenstein and Yoshinobu Kanno. Studies on an epithelial (gland) cell junction. Journal of Cell Biology, 22(3): 565–??, September 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/22/3/565.

Lansing:1961:FSC

[LL61a] Albert I. Lansing and Francois Lamy. Fine structure of the cilia of rotifers. Journal of Cell Biology, 9(4):799–??, April 1961. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/9/4/799.

Lansing:1961:LAA

[LL61b] Albert I. Lansing and François Lamy. Localization of ATP ase in rotifer cilia. Journal of Cell Biology, 11(2):498–??, November 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/11/2/
498.

Loud:1963:IMG

[LM63] Alden V. Loud and Yutaka Mishima. The induction of melanization in goldfish scales with acth, in vitro. Journal of Cell Biology, 18(1):181–??, July 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/18/1/181.

Latta:1961:CDS

[LMM61] Harrison Latta, Arvid B. Maunsbach, and Sidney C. Madden. Cilia in different segments of the rat nephron. Journal of Cell Biology, 11(1):248–??, October 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/11/1/248.

Locke:1961:PCR

[Loc61] Michael Locke. Pore canals and related structures in insect cuticle. Journal of Cell Biology, 10(4):589–??, August 1961. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/10/4/589.

Loud:1962:MQE

[Lou62] Alden V. Loud. A method for the quantitative estimation of cytoplasmic structures. Journal of Cell Biology, 15(3):481–??, December 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
15/3/481.

Ledbetter:1963:MPC

[LP63] M. C. Ledbetter and K. R. Porter. A “microtubule” in plant cell fine structure. Journal of Cell Biology, 19(1):239–??, October 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/19/1/
239.

Luck:1961:GSU

[Luc61] David J. L. Luck. Glycogen synthesis from uridine Diphosphate glucose. Journal of Cell Biology, 10(2):195–??, June 1961. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/10/2/195.

Luck:1963:FMN

[Luc63] David J. L. Luck. Formation of mitochondria in neurospora crassa. Journal of Cell Biology, 16(3):483–??, March 1963. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/16/3/483.

Luft:1961:IER

[Luf61] John H. Luft. Improvements in epoxy resin embedding methods. Journal of Cell Biology, 9(2):409–??, February 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/9/2/409.

Linnane:1962:SOY

[LVN62] Anthony W. Linnane, Eberhards Vitols, and Patricia G. Nowland. Studies on the origin of yeast mitochondria. Journal of Cell Biology, 13(2):345–??, May 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/13/2/345.

Lasansky:1962:EST

[LW62] Arnaldo Lasansky and Flora Wald. The extracellular space in the toad retina as defined by the distribution of ferrocyanide. Journal of Cell Biology, 15(3):463–??, December 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/15/3/463.

Love:1963:CSN

[LW63] Robert Love and Peter Wildy. Cytochemical studies of the nucleoproteins of HeLa cells infected with herpes virus. Journal of Cell Biology, 17(2):237–??, May 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/17/2/237.

Leslie:1960:SDC

[LY60] I. Leslie and Margaret Yarnell. Succinic Dehydrogenase and Cytochrome Oxidase Activities in Cell Cultures. Journal of Cell Biology, 7(2):265–??, April 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/7/2/265.

Mauro:1961:SSF

[MA61] Alexander Mauro and W. Robert Adams. The structure of the sarcolemma of the frog skeletal muscle fiber. Journal of Cell Biology, 10(4):177–??, August 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/10/4/177.

MacRae:1963:OFS

[Mac63] Edith Krugelis MacRae. Observations on the fine structure of pharyngeal muscle in the planarian dugesia tigrina. Journal of Cell Biology, 18(3):651–??, September 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/18/3/651.

MacRae:1964:OSN

[Mac64] Edith Krugelis MacRae. Organized structures in the nucleolus. Journal of Cell Biology, 23(1):195–??, October 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/23/1/195.

Muscatello:1961:SSF

[MACAvdD61] U. Muscatello, Ebba Andersson-Cedergren, G. F. Azzone, and Alexandra von der Decken. The sarcotubular system of frog skeletal muscle. Journal of Cell Biology, 10(4):201–??, August 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/10/4/
201.

Malamed:1963:UMP

[Mal63] Sasha Malamed. Use of a microcentrifuge for preparation of isolated mitochondria and cell suspensions for electron microscopy. Journal of Cell Biology, 18(3):696–??, September 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/18/3/696.

Manton:1960:RDG

[Man60] Irene Manton. On a reticular derivative from Golgi bodies in the meristem of Anthoceros. Journal of Cell Biology, 8(1): 221–??, September 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/8/1/221.

Marinozzi:1961:SIU

[Mar61] Vittorio Marinozzi. Silver impregnation of ultrathin sections for electron microscopy. Journal of Cell Biology, 9(1):121–??, January 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/9/
1/121.

Maturana:1960:FAO

[Mat60] Humberto R. Maturana. The fine anatomy of the optic nerve of Anurans — an electron microscope study. Journal of Cell Biology, 7(1):107–??, February 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/7/1/107.

Matsumoto:1963:EMS

[Mat63] Seiichi Matsumoto. Electron microscope studies of rabies virus in mouse brain. Journal of Cell Biology, 19(3):565–??, December 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/19/3/
565.

Mauro:1961:SCS

[Mau61] Alexander Mauro. Satellite cell of skeletal muscle fibers. Journal of Cell Biology, 9(2):493–??, February 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/9/2/493.

Munger:1961:EMS

[MB61] Bryce L. Munger and Saul W. Brusilow. An electron microscopic study of eccrine sweat glands of the cat foot and toe pads — evidence for ductal reabsorption in the human. Journal of Cell Biology, 11(2):403–??, November 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/11/2/403.

Moe:1962:CBC

[MB62] H. Moe and O. Behnke. Cytoplasmic bodies containing mitochondria, ribosomes, and rough surfaced endoplasmic membranes in the epithelium of the small intestine of newborn rats. Journal of Cell Biology, 13(1):168–??, April 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/13/1/168.

Marchesi:1963:DEA

[MB63a] V. T. Marchesi and R. J. Barrnett. The demonstration of enzymatic activity in pinocytic vesicles of blood capillaries with the electron microscope. Journal of Cell Biology, 17(3):547–??, June 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
17/3/547.

Meek:1963:LTP

[MB63b] G. A. Meek and S. Bradbury. Localization of thiamine pyrophosphatase activity in the Golgi apparatus of a mollusc, Helix Aspersa. Journal of Cell Biology, 18(1):73–??, July 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/18/1/73.

Merrillees:1963:CFS

[MBH63] Neil C. R. Merrillees, Geoffrey Burnstock, and Mollie E. Holman. Correlation of fine structure and physiology of the innervation of smooth muscle in the guinea pig Vas Deferens. Journal of Cell Biology, 19(3):529–??, December 1963. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/19/3/529.

Mercer:1962:SCC

[MBM62] Frank V. Mercer, Lawrence Bogorad, and Rosemary Mullens. Studies with cyanidium caldarium. Journal of Cell Biology, 13(3):393–??, June 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/13/3/393.

Micou:1962:NCR

[MCC62] Julie Micou, Carter C. Collins, and T. Timothy Crocker. Nuclear-cytoplasmic relationships in human cells in tissue culture. Journal of Cell Biology, 12(1):195–??, January 1962. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/12/1/195.

McDonald:1962:SDA

[McD62] Barbara B. McDonald. Synthesis of deoxyribonucleic acid by micro- and macronuclei of Tetrahymena Pyriformis. Journal of Cell Biology, 13(2):193–??, May 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/13/2/193.

McAlear:1961:MLA

[MCT61] James H. McAlear, George Camougis, and Lloyd F. Thibodeau. Mapping of large areas with the electron microscope. Journal of Cell Biology, 10(1):133–??, May 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/10/1/133.

Moorhead:1963:ADS

[MD63] P. S. Moorhead and Vittorio Defendi. Asynchrony of DNA synthesis in chromosomes of human diploid cells. Journal of Cell Biology, 16(1):202–??, January 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/16/1/202.

Molbert:1960:DAP

[MDvD60] Elisabeth R. G. Mölbert, Franz Duspiva, and Otto H. von Deimlimg. The Demonstration of Alkaline Phosphatase in the Electron Microscope. Journal of Cell Biology, 7(2):387–??, April 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/7/2/
387.

Mendelsohn:1961:TWM

[Men61] Mortimer L. Mendelsohn. The two-wavelength method of microspectrophotometry. Journal of Cell Biology, 11(3):509– ??, December 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/11/3/509.

Merrillees:1960:FSM

[Mer60] Neil C. R. Merrillees. The Fine Structure of Muscle Spindles in the Lumbrical Muscles of the Rat. Journal of Cell Biology, 7(4):725–??, July 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/7/4/725.

Merriam:1961:FSC

[Mer61] R. W. Merriam. On the fine structure and composition of the nuclear envelope. Journal of Cell Biology, 11(3):559–??, December 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/11/3/

Merriam:1962:SDA

[Mer62] R. W. Merriam. Some dynamic aspects of the nuclear envelope. Journal of Cell Biology, 12(1):79–??, January 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/12/1/79.

Millington:1962:EMS

[MF62] P. F. Millington and J. B. Finean. Electron microscope studies of the structure of the microvilli on principal epithelial cells of rat jejunum after treatment in hypo- and hypertonic saline. Journal of Cell Biology, 14(1):125–??, July 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/14/1/125.

Mezger-Freed:1963:ESD

[MF63] Liselotte Mezger-Freed. Evidence for synthesis of deoxyribonucleic acid in enucleated rana pipiens eggs. Journal of Cell Biology, 18(2):471–??, August 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/18/2/471.

Mazia:1960:MMC

[MHB60] Daniel Mazia, Patricia J. Harris, and Thomas Bibring. The Multiplicity of the Mitotic Centers and the Time–Course of Their Duplication and Separation. Journal of Cell Biology, 7 (1):1–??, February 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/7/1/1.

Mayor:1961:SMR

[MHR61] Heather Donald Mayor, James C. Hampton, and Benjamin Rosario. A simple method for removing the resin from epoxyembedded tissue. Journal of Cell Biology, 9(4):909–??, April 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/9/4/
909.

Miller:1960:HAC

[Mil60] Fritz Miller. Hemoglobin absorption by the cells of the proximal convoluted tubule in mouse kidney. Journal of Cell Biology, 8(3):689–??, December 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/8/3/689.

Miller:1961:LGC

[Mil61a] Fritz Miller. Lipoprotein granules in the cortical collecting tubules of mouse kidney. Journal of Cell Biology, 9(1):157–??, January 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/9/
1/157.

Millonig:1961:MPL

[Mil61b] Giuseppe Millonig. A modified procedure for lead staining of thin sections. Journal of Cell Biology, 11(3):736–??, December 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/11/3/
736.

Millington:1964:CTL

[Mil64] P. F. Millington. Comparison of the thicknesses of the lateral wall membrane and the microvillus membrane of intestinal epithelial cells from rat and mouse. Journal of Cell Biology, 20 (3):514–??, March 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/20/3/514.

Mitsui:1960:AEM

[Mit60] Tadao Mitsui. Application of the Electron Microscope to the Cytochemical Peroxidase Reaction in Salamander Leukocytes. Journal of Cell Biology, 7(2):251–??, April 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/7/2/251.

Mayor:1963:AFS

[MJ63] Heather Donald Mayor and Liane E. Jordan. Acrolein — a fine structure fixative for viral cytochemistry. Journal of Cell Biology, 18(1):207–??, July 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/18/1/207.

McMaster-Kaye:1960:MCN

[MK60a] Rachel McMaster-Kaye. The metabolic characteristics of nucleolar, chromosomal, and cytoplasmic ribonucleic acid of Drosophila salivary glands. Journal of Cell Biology, 8(2):365–??, October 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/8/
2/365.

Merriam:1960:RCS

[MK60b] R. W. Merriam and William E. Koch. The Relative Concentration of Solids in the Nucleolus, Nucleus, and Cytoplasm of the Developing Nerve Cell of the Chick. Journal of Cell Biology, 7(1):151–??, February 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/7/1/151.

Meek:1961:MIM

[MM61] G. A. Meek and M. J. Moses. Microtubulation of the inner membrane of the nuclear envelope. Journal of Cell Biology, 10(1):121–??, May 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/10/1/121.

Moor:1963:FSF

[MM63a] H. Moor and K. Mühlethaler. Fine structure in frozen-etched yeast cells. Journal of Cell Biology, 17(3):609–??, June 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/17/3/609.

Moore:1963:FSM

[MM63b] Royall T. Moore and James H. McAlear. Fine structure of mycota. Journal of Cell Biology, 16(1):131–??, January 1963. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/16/1/131.

Morre:1964:IGA

[MM64] D. James Morré and Hilton H. Mollenhauer. Isolation of the Golgi apparatus from plant cells. Journal of Cell Biology, 23 (2):295–??, November 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/23/2/295.

Mackenzie:1961:EPG

[MMB61] Cosmo G. Mackenzie, Julia B. Mackenzie, and Paul Beck. The effect of pH on growth, protein synthesis, and lipid-rich particles of cultured mammalian cells. Journal of Cell Biology, 9 (1):141–??, January 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/9/1/141.

Mazia:1961:DIM

[MMMH61] Daniel Mazia, J. M. Mitchison, Heitor Medina, and Patricia Harris. The direct isolation of the mitotic apparatus. Journal of Cell Biology, 10(4):467–??, August 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/10/4/467.

MacKenzie:1962:IPC

[MMR62] Cosmo G. MacKenzie, Julia B. MacKenzie, and Oscar K. Reiss. Isolation and properties of a cell from liver characterized by lipid-rich particles. Journal of Cell Biology, 14(2):269–??, August 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
14/2/269.

Moor:1961:NFU

[MMWFW61] H. Moor, K. Mühlethaler, H. Waldner, and A. Frey-Wyssling. A new freezing-ultramicrotome. Journal of Cell Biology, 10(1):1– ??, May 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
10/1/1.

Mercer:1962:EMD

[MNP62] F. V. Mercer, Maret Nittim, and J. V. Possingham. The effect of manganese deficiency on the structure of spinach chloroplasts. Journal of Cell Biology, 15(2):379–??, November 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/15/2/379.

Monro:1961:SSF

[Mon61] R. E. Monro. Serological studies on the formation of protein parasporal inclusions in bacillus thuringiensis. Journal of Cell Biology, 11(2):321–??, November 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/11/2/321.

Monesi:1962:ASD

[Mon62] Valerio Monesi. Autoradiographic study of DNA synthesis and the cell cycle in spermatogonia and spermatocytes of mouse testis using tritiated thymidine. Journal of Cell Biology, 14 (1):1–??, July 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/14/1/1.

Monesi:1964:RAS

[Mon64] Valerio Monesi. Ribonucleic acid synthesis during mitosis and meiosis in the mouse testis. Journal of Cell Biology, 22(3): 521–??, September 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/22/3/521.

Moses:1961:SCPa

[Mos61a] Montrose J. Moses. Spermiogenesis in the crayfish (Procambarus Clarkii). Journal of Cell Biology, 9(1):222–??, January 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/9/1/222.

Moses:1961:SCPb

[Mos61b] Montrose J. Moses. Spermiogenesis in the crayfish (Procambarus Clarkii). Journal of Cell Biology, 10(3):301–??, July 1961. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/10/3/301.

Majno:1961:SIa

[MP61] G. Majno and G. E. Palade. Studies on inflammation. Journal of Cell Biology, 11(3):571–??, December 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/11/3/571.

Muir:1962:QLM

[MP62] A. R. Muir and A. Peters. Quintuple-layered membrane junctions at terminal bars between endothelial cells. Journal of Cell Biology, 12(2):443–??, February 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/12/2/443.

Marin:1964:FSC

[MP64a] G. Marin and D. M. Prescott. The frequency of sister chromatid exchanges following exposure to varying doses of H^{3}-thymidine or X-rays. Journal of Cell Biology, 21(2):159–??, May 1964. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/21/2/159.

Miller:1964:LAR

[MP64b] Fritz Miller and George E. Palade. Lytic activities in renal protein absorption droplets. Journal of Cell Biology, 23(3):

519–??, December 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/23/3/519.

Morgan:1961:AAA

[MPH61] W. S. Morgan, P. Perlmann, and T. Hultin. Autoradiographic analysis on agar plates of antigens from sub cellular fractions of rat liver slices. Journal of Cell Biology, 10(3):411–??, July 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/10/3/
411.

Morgan:1962:CPA

[MPH62] W. S. Morgan, P. Perlmann, and T. Hultin. Characterization of the pattern of amino acid incorporation in cell-free liver systems by autoradiography of precipitin reactions. Journal of Cell Biology, 12(3):628–??, March 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/12/3/628.

Majno:1961:SIb

[MPS61] G. Majno, G. E. Palade, and Gutta I. Schoefl. Studies on inflammation. Journal of Cell Biology, 11(3):607–??, December 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/11/3/
607.

Moody:1960:FSS

[MR60] M. F. Moody and J. D. Robertson. The Fine Structure of Some Retinal Photoreceptors. Journal of Cell Biology, 7(1): 87–??, February 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/7/1/87.

Maser:1963:SEC

[MR63a] Morton D. Maser and Robert V. Rice. Soluble earthworm cuticle collagen: a possible dimer of tropocollagen. Journal of Cell Biology, 18(3):569–??, September 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/18/3/569.

Munger:1963:CNP

[MR63b] Bryce L. Munger and Sanford I. Roth. The cytology of the normal parathyroid glands of man and Virginia deer. Journal of Cell Biology, 16(2):379–??, February 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/16/2/379.

Moule:1960:BMS

[MRC60] Y. Moulé, C. Rouiller, and J. Chauveau. A Biochemical and Morphological Study of Rat Liver Microsomes. Journal of Cell Biology, 7(3):547–??, June 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/7/3/547.

Mercer:1960:EMS

[MS60] E. H. Mercer and B. M. Shaffer. Electron Microscopy of Solitary and Aggregated Slime Mould Cells. Journal of Cell Biology, 7 (2):353–??, April 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/7/2/353.

Malkoff:1963:UIS

[MS63] Donald B. Malkoff and Bernard L. Strehler. The ultrastructure of isolated and in situ human cardiac age pigment. Journal of Cell Biology, 16(3):611–??, March 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/16/3/611.

McLean:1964:CLP

[MS64a] J. D. McLean and S. J. Singer. Cross-linked polyampholytes. Journal of Cell Biology, 20(3):518–??, March 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/20/3/518.

McNabb:1964:FMB

[MS64b] J. D. McNabb and E. Sandborn. Filaments in the microvillous border of intestinal cells. Journal of Cell Biology, 22(3): 701–??, September 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/22/3/701.

Martin:1963:CMS

[MSBN63] G. R. Martin, E. Schiffmann, H. A. Bladen, and Marie Nylen. Chemical and morphological studies on the in vitro calcification of aorta. Journal of Cell Biology, 16(2):243–??, February 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/16/2/
243.

Maggio:1963:SINa

[MSP63a] Rachele Maggio, Philip Siekevitz, and George E. Palade. Studies on isolated nuclei. Journal of Cell Biology, 18(2):267–??, August 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
18/2/267.

Maggio:1963:SINb

[MSP63b] Rachele Maggio, Philip Siekevitz, and George E. Palade. Studies on isolated nuclei. Journal of Cell Biology, 18(2):293–??, August 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
18/2/293.

Miller:1964:ASM

[MSP64] O. L. Miller, G. E. Stone, and D. M. Prescott. Autoradiography of soluble materials. Journal of Cell Biology, 23(3):654–??, December 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
23/3/654.

Mugnaini:1964:HFA

[Mug64] Enrico Mugnaini. Helical filaments in astrocytic mitochondria of the corpus striatum in the rat. Journal of Cell Biology, 23 (1):173–??, October 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/23/1/173.

Munden:1960:SCM

[Mun60] H. R. Munden. A set of collets for manipulating the specimen holder of the Siemens electron microscope. Journal of Cell Biology, 8(2):541–??, October 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/8/2/541.

Munger:1961:SMA

[Mun61a] Bryce L. Munger. Staining methods applicable to sections of osmium-fixed tissue for light microscopy. Journal of Cell Biology, 11(2):502–??, November 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/11/2/502.

Munger:1961:UHH

[Mun61b] Bryce L. Munger. The ultrastructure and histophysiology of human eccrine sweat glands. Journal of Cell Biology, 11(2): 385–??, November 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/11/2/385.

Mollenhauer:1963:OFG

[MW63a] Hilton H. Mollenhauer and W. Gordon Whaley. An observation on the functioning of the Golgi apparatus. Journal of Cell Biology, 17(1):222–??, April 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/17/1/222.

Mussgay:1963:EMS

[MW63b] M. Mussgay and J. Weibel. Electron microscopic studies on the development of vesicular stomatitis virus in Kb cells. Journal of Cell Biology, 16(1):119–??, January 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/16/1/119.

Mollenhauer:1960:PFG

[MZ60] Hilton H. Mollenhauer and William Zebrun. Permanganate fixation of the Golgi complex and other cytoplasmic structures of mammalian testes. Journal of Cell Biology, 8(3):761–??, December 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/8/
3/761.

Nachmias:1964:FSC

[Nac64] Vivianne T. Nachmias. Fibrillar structures in the cytoplasm of chaos chaos. Journal of Cell Biology, 23(1):183–??, October 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/23/1/
183.

Nadakavukaren:1964:FSN

[Nad64] Mathew J. Nadakavukaren. Fine structure of negatively stained plant mitochondria. Journal of Cell Biology, 23(1):193–??, October 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
23/1/193.

Nagano:1962:OFS

[Nag62] Toshio Nagano. Observations on the fine structure of the developing spermatid in the domestic chicken. Journal of Cell Biology, 14(2):193–??, August 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/14/2/193.

Nagano:1963:FSC

[Nag63] Toshio Nagano. Fine structural changes in the flagellum of the spermatid in experimental cryptorchidism of the rat. Journal of Cell Biology, 18(2):337–??, August 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/18/2/337.

Nakai:1964:SUL

[Nak64] Takashi Nakai. A study of the ultrastructural localization of hair keratin synthesis utilizing electron microscopic autoradiography in a magnetic field. Journal of Cell Biology, 21(1):63–??, April 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/21/1/
63.

Napolitano:1963:CMA

[Nap63a] Leonard Napolitano. Cytolysomes in metabolically active cells. Journal of Cell Biology, 18(2):478–??, August 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/18/2/478.

Napolitano:1963:DWA

[Nap63b] Leonard Napolitano. The differentiation of white adipose cells. Journal of Cell Biology, 18(3):663–??, September 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/18/3/663.

Nelson:1963:SCT

[NB63] Douglas A. Nelson and Ellis S. Benson. On the structural continuities of the transverse tubular system of rabbit and human myocardial cells. Journal of Cell Biology, 16(2):297–??, February 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/16/2/

Novikoff:1962:EMS

[NdTBB62] Alex B. Novikoff, Guy de Thé, D. Beard, and J. W. Beard. Electron microscopic study of the ATP ase activity of the bai strain a (myeloblastosis) avian tumor virus. Journal of Cell Biology, 15(3):451–??, December 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/15/3/451.

Novikoff:1962:CMD

[NE62] Alex B. Novikoff and Edward Essner. Cytolysomes and mitochondrial degeneration. Journal of Cell Biology, 15(1):140–??, October 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
15/1/140.

Nylen:1963:CGR

[NEO63] M. U. Nylen, E. D. Eanes, and K.-Å. Omnell. Crystal growth in rat enamel. Journal of Cell Biology, 18(1):109–??, July 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/18/1/109.

Neville:1960:ICM

[Nev60] David M. Neville. The isolation of a cell membrane fraction from rat liver. Journal of Cell Biology, 8(2):413–??, October 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/8/2/413.

Nonnenmacher-Godet:1964:ITT

[NGD64] Jacqueline Nonnenmacher-Godet and Ellsworth C. Dougherty. Incorporation of tritiated thymidine in the cells of Caenorhabditis Briggsae (Nematoda) reared in axenic culture. Journal of Cell Biology, 22(1):281–??, July 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/22/1/281.

Nelson:1960:IVC

[NHK60] Erland Nelson, Hermann Hager, and Ernest Kovács. Intracellular virus crystals in the central nervous system of mice following rodent polioencephalitis (mm strain) infection. Journal of Cell Biology, 8(3):825–??, December 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/8/3/825.

Nicklas:1961:RBD

[Nic61] R. Bruce Nicklas. The relationship between DNA content and alternative meiotic patterns in certain discocephalinids (pentatomidae; heteroptera). Journal of Cell Biology, 9(2): 486–??, February 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/9/2/486.

Nayler:1964:SOF

[NM64] Winifred G. Nayler and Neil C. R. Merrillees. Some observations on the fine structure and metabolic activity of normal and glycerinated ventricular muscle of toad. Journal of Cell Biology, 22(3):533–??, September 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/22/3/533.

Nachlas:1960:IHL

[NMRS60] Marvin M. Nachlas, Benito Monis, David Rosenblatt, and Arnold M. Seligman. Improvement in the Histochemical Localization of Leucine Aminopeptidase with a New Substrate, L -Leucyl-4-Methoxy-2-Naphthylamide. Journal of Cell Biology, 7(2):261–??, April 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/7/2/261.

Nunez-Montiel:1960:EMS

[NMW60] O. Núñez-Montiel and J. Weibel. Electron microscope study of echo 19 virus infection in monkey kidney cells. Journal of Cell Biology, 8(1):291–??, September 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/8/1/291.

Nunez-Montiel:1961:EMS

[NMWVF61] O. Núñez-Montiel, J. Weibel, and J. Vitelli-Flores. Electron microscopic study of the cytopathology of echo virus infection in cultivated cells. Journal of Cell Biology, 11(2):457–??, November 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/11/2/
457.

Nass:1963:IFDa

[NN63a] Margit M. K. Nass and Sylvan Nass. Intramitochondrial fibers with DNA characteristics. Journal of Cell Biology, 19(3):

593–??, December 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/19/3/593.

Nass:1963:IFDb

[NN63b] Sylvan Nass and Margit M. K. Nass. Intramitochondrial fibers with DNA characteristics. Journal of Cell Biology, 19(3): 613–??, December 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/19/3/613.

Niemi:1960:RME

[NP60] Mikko Niemi and A. G. E. Pearse. The relationship of the mitochondria to egg white absorption droplets in rat kidney. Journal of Cell Biology, 8(1):279–??, September 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/8/1/279.

Nishiura:1960:DET

[NR60] Mitsugu Nishiura and S. R. S. Rangan. A Device for Embedding Tissue Culture Preparations Grown on Coverslips. Journal of Cell Biology, 7(2):411–??, April 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/7/2/411.

Novikoff:1961:MLO

[NSD61] Alex B. Novikoff, Woo-Yung Shin, and Joan Drucker. Mitochondrial localization of oxidative enzymes: Staining results with two tetrazolium salts. Journal of Cell Biology, 9(1):47–??, January 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/9/
1/47.

Nebel:1960:NSD

[NTM60] Bernard R. Nebel, Sylvanus A. Tyler, and Carol J. Murphy. A note on the statistical determination of shape of chromatin elements in human spermatid and Tradescantia microsporocyte. Journal of Cell Biology, 7(2):377–??, April 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/7/2/377.

OBrien:1964:PCA

[O’B64a] B. R. A. O’Brien. The partial cytolysis of the amphibian erythrocyte and liver parencyhma cell by a non-ionogenic surface active agent. Journal of Cell Biology, 20(3):521–??, March 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/20/3/521.

OBrien:1964:RMI

[O’B64b] B. R. A. O’Brien. A rapid method for the isolation and collection of nuclei from whole cell suspensions. Journal of Cell Biology, 20(3):525–??, March 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/20/3/525.

Orzalesi:1964:FSI

[OB64c] N. Orzalesi and A. Bairati. Filamentous structures in the inner segment of human retinal rods. Journal of Cell Biology, 20 (3):509–??, March 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/20/3/509.

Ohad:1964:DCM

[OD64] I. Ohad and D. Danon. On the dimensions of cellulose microfibrils. Journal of Cell Biology, 22(1):302–??, July 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/22/1/302.

Ohad:1962:SCA

[ODH62] I. Ohad, D. Danon, and S. Hestrin. Synthesis of cellulose by acetobacter xylinum. Journal of Cell Biology, 12(1):31–??, January 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/12/1/
31.

Ohad:1963:USC

[ODH63] I. Ohad, D. Danon, and S. Hestrin. The use of shadow-casting technique for measurement of the width of elongated particles. Journal of Cell Biology, 17(2):321–??, May 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/17/2/321.

Odor:1960:EMS

[Odo60] D. Louise Odor. Electron Microscopic Studies on Ovarian Oocytes and Unfertilized Tubal Ova in the Rat. Journal of Cell Biology, 7(3):567–??, June 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/7/3/567.

Oren:1963:MPT

[OFS^{+}63] Rachel Oren, Ann E. Farnham, Kazuhisa Saito, Eva Milofsky, and Manfred L. Karnovsky. Metabolic patterns in three types of phagocytizing cells. Journal of Cell Biology, 17(3):487–??, June 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/17/3/
487.

Olsson:1962:RBC

[Ols62] Ragnar Olsson. The relationship between ciliary rootlets and other cell structures. Journal of Cell Biology, 15(3):596–??, December 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
15/3/596.

Ohye:1962:FSS

[OM62a] D. F. Ohye and W. G. Murrell. Formation and structure of the spore of bacillus coagulans. Journal of Cell Biology, 14(1):111– ??, July 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
14/1/111.

Orth:1962:EIA

[OM62b] David N. Orth and Howard E. Morgan. The effect of insulin, alloxan diabetes, and anoxia on the ultrastructure of the rat heart. Journal of Cell Biology, 15(3):509–??, December 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/15/3/509.

Owen:1963:CPKb

[OM63] Maureen Owen and Sheila MacPherson. Cell population kinetics of an osteogenic tissue • II. Journal of Cell Biology, 19 (1):33–??, October 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/19/1/33.

Overton:1964:FSC

[OS64] Jane Overton and Jane Shoup. Fine structure of cell surface specializations in the maturing duodenal mucosa of the chick. Journal of Cell Biology, 21(1):75–??, April 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/21/1/75.

Osinchak:1964:EML

[Osi64] Joseph Osinchak. Electron microscopic localization of acid phosphatase and thiamine pyrophosphatase activity in hypothalamic neurosecretory cells of the rat. Journal of Cell Biology, 21(1):35–??, April 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/21/1/35.

Overton:1963:ICO

[Ove63] Jane Overton. Intercellular connections in the outgrowing stolon of cordylophora. Journal of Cell Biology, 17(3):661–??, June 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
17/3/661.

Otero-Vilardebo:1963:DMA

[OVLG63] Luis R. Otero-Vilardebó, Nathan Lane, and Gabriel C. Godman. Demonstration of mitochondrial ATP ase activity in formalin-fixed colonic epithelial cells. Journal of Cell Biology, 19(3):647–??, December 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/19/3/647.

Otero-Vilardebo:1964:LPA

[OVLG64] Luis R. Otero-Vilardebó, Nathan Lane, and Gabriel C. Godman. Localization of phosphatase activities in colonic goblet and absorptive cells. Journal of Cell Biology, 21(3):486–??, June 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/21/3/
486.

Owen:1963:CPKa

[Owe63] Maureen Owen. Cell population kinetics of an osteogenic tissue • I. Journal of Cell Biology, 19(1):19–??, October 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/19/1/19.

Ozzello:1964:EEG

[Ozz64] Luciano Ozzello. Effects of 17-β-estradiol and of growth hormone on the production of acid mucopolysaccharides by human embryonal fibroblasts in vitro. Journal of Cell Biology, 21(2):283–??, May 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/21/2/283.

Packer:1963:SSTa

[Pac63a] Lester Packer. Size and shape transformations correlated with oxidative phosphorylation in mitochondria. Journal of Cell Biology, 18(3):487–??, September 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/18/3/487.

Packer:1963:SSTb

[Pac63b] Lester Packer. Size and shape transformations correlated with oxidative phosphorylation in mitochondria. Journal of Cell Biology, 18(3):495–??, September 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/18/3/495.

Painter:1961:ARH

[Pai61] Robert B. Painter. Asynchronous replication of HeLa S3 chromosomal deoxyribonucleic acid. Journal of Cell Biology, 11(2): 485–??, November 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/11/2/485.

Palay:1960:AAF

[Pal60] Sanford L. Palay. On the Appearance of Absorbed Fat Droplets in the Nuclear Envelope. Journal of Cell Biology, 7(2):391–??, April 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/7/
2/391.

Pannese:1962:DNP

[Pan62] Ennio Pannese. Detection of neurofilaments in the perikaryon of hypertrophic nerve cells. Journal of Cell Biology, 13(3):457–??, June 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
13/3/457.

Parsons:1961:SMO

[Par61] D. F. Parsons. A simple method for obtaining increased contrast in araldite sections by using postfixation staining of tissues with potassium permanganate. Journal of Cell Biology, 11(2): 492–??, November 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/11/2/492.

Parks:1962:UFE

[Par62a] Harold F. Parks. Unusual formations of ergastoplasm in parotid acinous cells of mice. Journal of Cell Biology, 14(2):221–??, August 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
14/2/221.

Parsons:1962:EMS

[Par62b] Donald F. Parsons. An electron microscope study of radiation damage in the mouse oocyte. Journal of Cell Biology, 14(1):31– ??, July 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
14/1/31.

Parsons:1963:NST

[Par63] Donald F. Parsons. Negative staining of thinly spread cells and associated virus. Journal of Cell Biology, 16(3):620–??, March 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/16/3/
620.

Parsons:1961:EMPb

[PBD^{+}61] D. F. Parsons, M. A. Bender, E. B. Darden, Guthrie T. Pratt, and D. L. Lindsley. Electron microscopy of plasma-cell tumors of the mouse. Journal of Cell Biology, 9(2):369–??, February 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
9/2/369.

Prescott:1962:DPC

[PBK62] D. M. Prescott, F. J. Bollum, and B. C. Kluss. Is DNA polymerase a cytoplasmic enzyme? Journal of Cell Biology, 13 (1):172–??, April 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/13/1/172.

Philpott:1963:FSC

[PC63] C. W. Philpott and D. E. Copeland. Fine structure of chloride cells from three species of fundulus. Journal of Cell Biology, 18(2):389–??, August 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/18/2/389.

Poindexter:1964:FSS

[PCB64] Jeanne L. Stove Poindexter and Germaine Cohen-Bazire. The fine structure of stalked bacteria belonging to the family caulobacteraceae. Journal of Cell Biology, 23(3):587–??, December 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
23/3/587.

Policard:1961:DSM

[PCDMP61] A. Policard, A. Collet, H. Daniel-Moussard, and S. Pregermain. Deposition of silica in mitochondria: an electron microscopic study. Journal of Cell Biology, 9(1):236–??, January 1961. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/9/1/236.

Perlmann:1964:AED

[PCK64] Peter Perlmann and Jane Couffer-Kaltenbach. Antigens in eggs and developmental stages of the sea urchin. Journal of Cell Biology, 22(2):307–??, August 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/22/2/307.

Parsons:1960:TSD

[PD60] D. F. Parsons and E. B. Darden. A technique for the simultaneous dirt-free lead staining of several electron microscope grids of thin sections. Journal of Cell Biology, 8(3):834–??, December 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/8/
3/834.

Parsons:1961:EMPa

[PDLP61] D. F. Parsons, E. B. Darden, D. L. Lindsley, and Guthrie T. Pratt. Electron microscopy of plasma-cell tumors of the mouse. Journal of Cell Biology, 9(2):353–??, February 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/9/2/353.

Peachey:1961:SLB

[Pea61] Lee D. Peachey. Structure of the longitudinal body muscles of amphioxus. Journal of Cell Biology, 10(4):159–??, August 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/10/4/
159.

Pease:1962:DHO

[Pea62] Daniel C. Pease. Demonstration of a highly ordered pattern upon a mitochondrial surface. Journal of Cell Biology, 15(2): 385–??, November 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/15/2/385.

Pease:1963:UFF

[Pea63] Daniel C. Pease. The ultrastructure of flagellar fibrils. Journal of Cell Biology, 18(2):313–??, August 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/18/2/313.

Peachey:1964:EMO

[Pea64] Lee D. Peachey. Electron microscopic observations on the accumulation of divalent cations in intramitochondrial granules. Journal of Cell Biology, 20(1):95–??, January 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/20/1/95.

Pelc:1964:LDC

[Pel64] S. R. Pelc. Labelling of DNA and cell division in so called non-dividing tissues. Journal of Cell Biology, 22(1):21–??, July 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/22/1/
21.

Pepe:1961:USAa

[Pep61] Frank A. Pepe. The use of specific antibody in electron microscopy. Journal of Cell Biology, 11(3):515–??, December 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/11/3/515.

Perry:1961:KNI

[Per61] R. P. Perry. Kinetics of nucleoside incorporation into nuclear and cytoplasmic RNA. Journal of Cell Biology, 11(1):1–??, October 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
11/1/1.

Perl:1964:CPD

[Per64] William Perl. Correction of polyribosome distributions as observed in cell sections by electron microscopy. Journal of Cell Biology, 22(3):613–??, September 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/22/3/613.

Peters:1960:FSM

[Pet60a] A. Peters. The formation and structure of myelin sheaths in the central nervous system. Journal of Cell Biology, 8(2):431–??, October 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/8/
2/431.

Peters:1960:SMS

[Pet60b] A. Peters. The structure of myelin sheaths in the central nervous system of Xenopus laevis (Daudin). Journal of Cell Biology, 7(1):121–??, February 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/7/1/121.

Peters:1961:RCC

[Pet61] A. Peters. A radial component of central myelin sheaths. Journal of Cell Biology, 11(3):733–??, December 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/11/3/733.

Peterson:1963:NSC

[Pet63] R. Price Peterson. A note on the structure of crayfish myofilaments. Journal of Cell Biology, 18(1):213–??, July 1963. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/18/1/213.

Peters:1964:FOS

[Pet64a] A. Peters. Further observations on the structure of myelin sheaths in the central nervous system. Journal of Cell Biology, 20(2):281–??, February 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/20/2/281.

Petersen:1964:DSC

[Pet64b] Arnold J. Petersen. DNA synthesis and chromosomal asynchrony. Journal of Cell Biology, 23(3):651–??, December 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/23/3/651.

Pepe:1961:USAb

[PF61] Frank A. Pepe and H. Finck. The use of specific antibody in electron microscopy. Journal of Cell Biology, 11(3):521–??, December 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
11/3/521.

Pellegrino:1963:EMS

[PF63a] Claudio Pellegrino and Clara Franzini. An electron microscope study of denervation atrophy in red and white skeletal muscle fibers. Journal of Cell Biology, 17(2):327–??, May 1963. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/17/2/327.

Plaut:1963:BGR

[PF63b] Andrew G. Plaut and William H. Fishman. Beta glucuronidaserich cytoplasmic particles in androgen-stimulated mouse kidney. Journal of Cell Biology, 16(2):253–??, February 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/16/2/253.

Pepe:1961:USAc

[PFH61] Frank A. Pepe, H. Finck, and H. Holtzer. The use of specific antibody in electron microscopy. Journal of Cell Biology, 11(3):533–??, December 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/11/3/533.

Padykula:1963:CSA

[PG63] Helen A. Padykula and Geraldine F. Gauthier. Cytochemical studies of adenosine triphosphatases in skeletal muscle fibers. Journal of Cell Biology, 18(1):87–??, July 1963. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/18/1/87.

Peachey:1962:SIT

[PH62] Lee D. Peachey and Andrew F. Huxley. Structural identification of twitch and slow striated muscle fibers of the frog. Journal of Cell Biology, 13(1):177–??, April 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/13/1/177.

Page:1963:FLS

[PH63] Sally G. Page and H. E. Huxley. Filament lengths in striated muscle. Journal of Cell Biology, 19(2):369–??, November 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/19/2/369.

Post:1964:RTP

[PH64] Joseph Post and Joseph Hoffman. The replication time and pattern of carcinogen-induced hepatoma cells. Journal of Cell Biology, 22(2):341–??, August 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/22/2/341.

Post:1963:RTP

[PHH63] Joseph Post, Cheng-Ya Huang, and Joseph Hoffman. The replication time and pattern of the liver cell in the growing rat. Journal of Cell Biology, 18(1):1–??, July 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/18/1/1.

Prescott:1962:CBT

[PKC62] D. M. Prescott, R. F. Kimball, and R. F. Carrier. Comparison between the timing of micronuclear and macronuclear DNA synthesis in euplotes eurystomus. Journal of Cell Biology, 13 (1):175–??, April 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/13/1/175.

Peterson:1964:SCC

[PL64] Marian Peterson and C. P. Leblond. Synthesis of complex carbohydrates in the Golgi region, as shown by radioautography after injection of labeled glucose. Journal of Cell Biology, 21 (1):143–??, April 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/21/1/143.

Porter:1960:SERa

[PM60] Keith R. Porter and Raul D. Machado. Studies on the Endoplasmic Reticulum. Journal of Cell Biology, 7(1):167–??, February 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
7/1/167.

Perdue:1962:IBS

[PM62] J. F. Perdue and W. H. McShan. Isolation and biochemical study of secretory granules from rat pituitary glands. Journal of Cell Biology, 15(2):159–??, November 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/15/2/159.

Parakkal:1964:SFS

[PM64] Paul F. Parakkal and A. Gedeon Matoltsy. A study of the fine structure of the epidermis of rana pipiens. Journal of Cell Biology, 20(1):85–??, January 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/20/1/85.

Palay:1962:FNT

[PMRGG62] Sanford L. Palay, S. M. McGee-Russell, Spencer Gordon, and Mary A. Grillo. Fixation of neural tissues for electron microscopy by perfusion with solutions of osmium tetroxide. Journal of Cell Biology, 12(2):385–??, February 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/12/2/385.

Porter:1961:SR

[Por61] Keith R. Porter. The sarcoplasmic reticulum. Journal of Cell Biology, 10(4):219–??, August 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/10/4/219.

Pauly:1960:RIC

[PP60] Helmut Pauly and Lester Packer. The Relationship of Internal Conductance and Membrane Capacity to Mitochondrial Volume. Journal of Cell Biology, 7(4):603–??, July 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/7/4/603.

Peterson:1961:FSI

[PP61] R. Price Peterson and Frank A. Pepe. The fine structure of inhibitory synapses in the crayfish. Journal of Cell Biology, 11 (1):157–??, October 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/11/1/157.

Pogo:1964:DDP

[PP64] Beatríz G. T. Pogo and A. O. Pogo. DNA dependence of plastid differentiation inhibition by actinomycin D. Journal of Cell Biology, 22(1):296–??, July 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/22/1/296.

Pauly:1960:EPM

[PPS60] Helmut Pauly, Lester Packer, and H. P. Schwan. Electrical Properties of Mitochondrial Membranes. Journal of Cell Biology, 7(4):589–??, July 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/7/4/589.

Peachey:1961:STU

[PR61] Lee D. Peachey and Howard Rasmussen. Structure of the toad’s urinary bladder as related to its physiology. Journal of Cell Biology, 10(4):529–??, August 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/10/4/529.

Pearse:1960:CSD

[PSH60] A. G. E. Pearse, D. G. Scarpelli, and R. Hess. A Cytochemical Study of the Dehydrogenases of Mitochondria and Mitochondrial Particulates by a Monotetrazolium–Cobalt Chelation Method. Journal of Cell Biology, 7(2):243–??, April 1960. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/7/2/243.

Pappas:1962:EMS

[PT62] George D. Pappas and Virginia M. Tennyson. An electron microscopic study of the passage of colloidal particles from the blood vessels of the ciliary processes and choroid Plexus of the rabbit. Journal of Cell Biology, 15(2):227–??, November 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/15/2/227.

Porter:1960:SERb

[PY60] Keith R. Porter and Eichi Yamada. Studies on the endoplasmic reticulum. Journal of Cell Biology, 8(1):181–??, September 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
8/1/181.

Ryser:1962:SPUb

[RAC62] H. Ryser, J. C. Aub, and J. B. Caulfield. Studies on protein uptake by isolated tumor cells. Journal of Cell Biology, 15(3): 437–??, December 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/15/3/437.

Rahman:1962:EML

[Rah62] Y. E. Rahman. Electron microscopy of lysosome-rich fractions from rat thymus isolated by density-gradient centrifugation before and after whole-body x-irradiation. Journal of Cell Biology, 13(2):253–??, May 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/13/2/253.

Ross:1961:WHC

[RB61] Russell Ross and Earl P. Benditt. Wound healing and collagen formation. Journal of Cell Biology, 11(3):677–??, December 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/11/3/
677.

Romanul:1962:LAH

[RB62a] Flaviu C. A. Romanul and Roger G. Bannister. Localized areas of high alkaline phosphatase activity in the terminal arterial tree. Journal of Cell Biology, 15(1):73–??, October 1962. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/15/1/73.

Ross:1962:WHCa

[RB62b] Russell Ross and Earl P. Benditt. Wound healing and collagen formation. Journal of Cell Biology, 12(3):533–??, March 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/12/3/533.

Ross:1962:WHCb

[RB62c] Russell Ross and Earl P. Benditt. Wound healing and collagen formation. Journal of Cell Biology, 15(1):99–??, October 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/15/1/99.

Ross:1964:WHC

[RB64] Russell Ross and Earl P. Benditt. Wound healing and collagen formation. Journal of Cell Biology, 22(2):365–??, August 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/22/2/
365.

Rabinovitch:1961:ARA

[RBF^{+}61] M. Rabinovitch, R. Brentani, S. Ferreira, N. Fausto, and T. Maack. Alkaline ribonuclease activity increase in rat kidney cortex and liver after trypan blue and other azo dyes administration. Journal of Cell Biology, 10(1):105–??, May 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/10/1/105.

Ross:1962:EMO

[RBL62] Leonard L. Ross, Murray B. Bornstein, and Gerard M. Lehrer. Electron microscopic observations of rat and mouse cerebellum in tissue culture. Journal of Cell Biology, 14(1):19–??, July 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/14/1/
19.

Robertson:1963:UMC

[RBS63] J. David Robertson, Thomas S. Bodenheimer, and David E. Stage. The ultrastructure of mauthner cell synapses and nodes in goldfish brains. Journal of Cell Biology, 19(1):159–??, October 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/19/1/
159.

Rho:1962:ITC

[RC62] Joon H. Rho and Margaret I. Chipchase. Incorporation of tritiated cytidine into ribonucleic acid by isolated pea nuclei. Journal of Cell Biology, 14(2):183–??, August 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/14/2/183.

Ryser:1962:SPUa

[RCA62] H. Ryser, J. B. Caulfield, and J. C. Aub. Studies on protein uptake by isolated tumor cells. Journal of Cell Biology, 14(2):255–??, August 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/14/2/255.

Roth:1961:IOC

[RD61] L. E. Roth and E. W. Daniels. Infective organisms in the cytoplasm of Amoeba Proteus. Journal of Cell Biology, 9(2): 317–??, February 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/9/2/317.

Roth:1962:EMS

[RD62] L. E. Roth and E. W. Daniels. Electron microscopic studies of mitosis in amebae. Journal of Cell Biology, 12(1):57–??, January 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
12/1/57.

Rifkind:1964:APD

[RDM64] Richard A. Rifkind, David Danon, and Paul A. Marks. Alterations in polyribosomes during erythroid cell maturation. Journal of Cell Biology, 22(3):599–??, September 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/22/3/599.

Rebhun:1961:AFS

[Reb61] Lionel I. Rebhun. Applications of freeze-substitution to electron microscope studies of invertebrate oocytes. Journal of Cell Biology, 9(4):785–??, April 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/9/4/785.

Reger:1961:FSN

[Reg61] James F. Reger. The fine structure of neuromuscular junctions and the sarcoplasmic reticulum of extrinsic eye muscles of fundulus heteroclitus. Journal of Cell Biology, 10(4):111–??, August 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
10/4/111.

Reith:1961:UAD

[Rei61] Edward J. Reith. The ultrastructure of ameloblasts during matrix formation and the maturation of enamel. Journal of Cell Biology, 9(4):825–??, April 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/9/4/825.

Reith:1963:UAD

[Rei63] Edward J. Reith. The ultrastructure of ameloblasts during early stages of maturation of enamel. Journal of Cell Biology, 18(3):691–??, September 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/18/3/691.

Revel:1962:SRB

[Rev62] J. P. Revel. The sarcoplasmic reticulum of the bat cricothyroid muscle. Journal of Cell Biology, 12(3):571–??, March 1962. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/12/3/571.

Reynolds:1963:LPC

[Rey63a] Edward S. Reynolds. Liver parenchymal cell injury. Journal of Cell Biology, 19(1):139–??, October 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/19/1/139.

Reynolds:1963:ULC

[Rey63b] Edward S. Reynolds. The use of lead citrate at high pH as an electron-opaque stain in electron microscopy. Journal of Cell Biology, 17(1):208–??, April 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/17/1/208.

Revel:1963:OMS

[RFP63] J. P. Revel, D. W. Fawcett, and C. W. Philpott. Observations on mitochondrial structure. Journal of Cell Biology, 16(1):187–??, January 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
16/1/187.

Rudkin:1962:PON

[RG62] George T. Rudkin and Helen A. Griech. On the persistence of oocyte nuclei from fetus to maturity in the laboratory mouse.

Journal of Cell Biology, 12(1):169–??, January 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/12/1/169.

Robbins:1964:VSM

[RG64a] Elliott Robbins and Nicholas K. Gonatas. In vitro selection of the mitotic cell for subsequent electron microscopy. Journal of Cell Biology, 20(2):356–??, February 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/20/2/356.

Robbins:1964:UMC

[RG64b] Elliott Robbins and Nicholas K. Gonatas. The ultrastructure of a mammalian cell during the mitotic cycle. Journal of Cell Biology, 21(3):429–??, June 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/21/3/429.

Rapp:1960:OMV

[RGB60] Fred Rapp, Irving Gordon, and Richard F. Baker. Observations of Measles Virus Infection of Cultured Human Cells. Journal of Cell Biology, 7(1):43–??, February 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/7/1/43.

Richter:1961:IAF

[Ric61] Goetz W. Richter. Intranuclear aggregates of ferritin in liver cells of mice treated with saccharated iron oxide. Journal of Cell Biology, 9(2):263–??, February 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/9/2/263.

Rieke:1962:VRL

[Rie62] William O. Rieke. The in vivo reutilization of lymphocytic and sarcoma DNA by cells growing in the peritoneal cavity. Journal of Cell Biology, 13(2):205–??, May 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/13/2/205.

Rizki:1961:ILK

[Riz61] M. T. M. Rizki. Intracellular localization of kynurenine in the fatbody of Drosophila. Journal of Cell Biology, 9(3):567–??, March 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/9/
3/567.

Rizki:1963:GCC

[Riz63] T. M. Rizki. Genetic control of cytodifferentiation. Journal of Cell Biology, 16(3):513–??, March 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/16/3/513.

Rizki:1964:MGR

[Riz64] T. M. Rizki. Mutant genes regulating the inducibility of kynurenine synthesis. Journal of Cell Biology, 21(2):203–??, May 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/21/2/203.

Richter:1963:HHH

[RK63] G. W. Richter and A. Kellner. Hypertrophy of the human heart at the level of fine structure. Journal of Cell Biology, 18(1):195– ??, July 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
18/1/195.

Roth:1964:FSN

[RL64] Lawrence M. Roth and Sarah A. Luse. Fine structure of the neurohypophysis of the opossum (Didelphis Virginiana). Journal of Cell Biology, 20(3):459–??, March 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/20/3/459.

Rode:1962:EMS

[RLF62] L. J. Rode, C. Willard Lewis, and J. W. Foster. Electron microscopy of spores of bacillus megaterium with special reference to the effects of fixation and thin sectioning. Journal of Cell Biology, 13(3):423–??, June 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/13/3/423.

Robbins:1963:DAO

[RM63] Elliott Robbins and Philip I. Marcus. Dynamics of acridine orange-cell interaction. Journal of Cell Biology, 18(2):237–??, August 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
18/2/237.

Rosenbaum:1964:CDK

[RM64] Robert M. Rosenbaum and Arnold Melman. Cytochemical differences in kidneys from winter-hibernating and aroused bats (Myotis Lucifugus), with particular reference to the Golgi zone. Journal of Cell Biology, 21(3):325–??, June 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/21/3/325.

Robbins:1964:DAO

[RMG64] Elliott Robbins, Philip I. Marcus, and Nicholas K. Gonatas. Dynamics of acridine orange-cell interaction. Journal of Cell Biology, 21(1):49–??, April 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/21/1/49.

Revel:1960:IGE

[RNF60] Jean Paul Revel, Leonard Napolitano, and Don W. Fawcett. Identification of glycogen in electron micrographs of thin tissue sections. Journal of Cell Biology, 8(3):575–??, December 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/8/3/575.

Rappaport:1963:FSN

[RNS63] Henry Rappaport, Takashi Nakai, and Hewson Swift. The fine structure of normal and neoplastic melanocytes in the Syrian hamster, with particular reference to carcinogen-induced melanotic tumors. Journal of Cell Biology, 16(1):171–??, January 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/16/1/
171.

Robbins:1961:STA

[Rob61a] Elliott Robbins. Some theoretical aspects of osmium tetroxide fixation with special reference to the metaphase chromosomes of cell cultures. Journal of Cell Biology, 11(2):449–??, November 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/11/2/

Robinow:1961:MYL

[Rob61b] C. F. Robinow. Mitosis in the yeast lipomyces lipofer. Journal of Cell Biology, 9(4):879–??, April 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/9/4/879.

Robertson:1963:OSP

[Rob63a] J. David Robertson. The occurrence of a subunit pattern in the unit membranes of club endings in mauthner cell synapses in goldfish brains. Journal of Cell Biology, 19(1):201–??, October 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/19/1/
201.

Robinow:1963:OCG

[Rob63b] C. F. Robinow. Observations on cell growth, mitosis, and division in the fungus basidiobolus ranarum. Journal of Cell Biology, 17(1):123–??, April 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/17/1/123.

Roth:1960:EMS

[ROD60] L. E. Roth, S. W. Obetz, and E. W. Daniels. Electron microscopic studies of mitosis in amebae. Journal of Cell Biology, 8(1):207–??, September 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/8/1/207.

Rose:1961:GCL

[Ros61] George G. Rose. The Golgi complex in living osteoblasts. Journal of Cell Biology, 9(2):463–??, February 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/9/2/463.

Rose:1962:UBS

[Ros62a] George G. Rose. An unusual behavior of strain HeLa cells in tissue culture. Journal of Cell Biology, 13(1):153–??, April 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/13/1/153.

Rosenbluth:1962:FSA

[Ros62b] Jack Rosenbluth. The fine structure of acoustic ganglia in the rat. Journal of Cell Biology, 12(2):329–??, February 1962. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/12/2/329.

Rosenbluth:1962:SCT

[Ros62c] Jack Rosenbluth. Subsurface cisterns and their relationship to the neuronal plasma membrane. Journal of Cell Biology, 13 (3):405–??, June 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/13/3/405.

Rosenberg:1963:REC

[Ros63a] Murray D. Rosenberg. The relative extensibility of cell surfaces. Journal of Cell Biology, 17(2):289–??, May 1963. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/17/2/289.

Rosenbluth:1963:FSE

[Ros63b] J. Rosenbluth. Fine structure of epineurial muscle cells in Aplysia Californica. Journal of Cell Biology, 17(2):455–??, May 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/17/2/
455.

Rosenbluth:1963:CBO

[Ros63c]  Jack Rosenbluth.  Contrast between osmium-fixed and

permanganate-fixed toad spinal ganglia. Journal of Cell Biology, 16(1):143–??, January 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/16/1/143.

Roth:1960:SFIa

[Rot60a] Jay S. Roth. Studies on the Function of Intracellular Ribonucleases. Journal of Cell Biology, 7(3):443–??, June 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/7/3/443.

Roth:1960:SFIb

[Rot60b] Jay S. Roth. Studies on the function of intracellular ribonucleases. Journal of Cell Biology, 8(3):665–??, December 1960. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/8/3/665.

Rose:1960:PCO

[RP60] George G. Rose and C. M. Pomerat. Phase contrast observations of the endoplasmic reticulum in living tissue cultures. Journal of Cell Biology, 8(2):423–??, October 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/8/2/423.

Rosenbluth:1961:FSN

[RP61] Jack Rosenbluth and Sanford L. Palay. The fine structure of nerve cell bodies and their myelin sheaths in the eighth nerve ganglion of the goldfish. Journal of Cell Biology, 9(4):853–??, April 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/9/
4/853.

Rabinovitch:1962:CDSa

[RP62a] M. Rabinovitch and W. Plaut. Cytoplasmic DNA synthesis in Amoeba Proteus. Journal of Cell Biology, 15(3):525–??, December 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/15/3/
525.

Rabinovitch:1962:CDSb

[RP62b] M. Rabinovitch and W. Plaut. Cytoplasmic DNA synthesis in Amoeba Proteus. Journal of Cell Biology, 15(3):535–??, December 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/15/3/
535.

Ris:1962:UDC

[RP62c] Hans Ris and Walter Plaut. Ultrastructure of DNA-containing areas in the chloroplast of chlamydomonas. Journal of Cell Biology, 13(3):383–??, June 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/13/3/383.

Roth:1964:YPU

[RP64] Thomas F. Roth and Keith R. Porter. Yolk protein uptake in the oocyte of the mosquito Aedes Aegypti. L. Journal of Cell Biology, 20(2):313–??, February 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/20/2/313.

Rodman:1963:PDD

[RPM63] N. F. Rodman, J. C. Painter, and N. B. McDevitt. Platelet disintegration during clotting. Journal of Cell Biology, 16(2): 225–??, February 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/16/2/225.

Rizki:1962:CRM

[RR62] T. M. Rizki and Rose M. Rizki. Cytodifferentiation in the rosy mutant of Drosophila Melanogaster. Journal of Cell Biology, 12(1):149–??, January 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/12/1/149.

Rizki:1963:IES

[RR63] T. M. Rizki and Rose M. Rizki. An inducible enzyme system in the larval cells of Drosophila Melanogaster. Journal of Cell Biology, 17(1):87–??, April 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/17/1/87.

Rizki:1964:FAI

[RR64] T. M. Rizki and Rose M. Rizki. Factors affecting the intracellualr synthesis of kynurenine. Journal of Cell Biology, 21 (1):27–??, April 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/21/1/27.

Ris:1961:EMS

[RS61a] Hans Ris and R. N. Singh. Electron microscope studies on bluegreen algae. Journal of Cell Biology, 9(1):63–??, January 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/9/1/63.

Rosen:1961:SOF

[RS61b] Walter G. Rosen and Kenneth A. Siegesmund. Some observations on the fine structure of a thermophilic, acidophilic alga. Journal of Cell Biology, 9(4):910–??, April 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/9/4/910.

Rebhun:1964:ISS

[RS64a] Lionel I. Rebhun and Thomas K. Sharpless. Isolation of spindles from the surf clam spisula solidissima. Journal of Cell Biology, 22(2):488–??, August 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/22/2/488.

Renaud:1964:DBB

[RS64b] Fernando L. Renaud and Hewson Swift. The development of basal bodies and flagella in allomyces arbusculus. Journal of Cell Biology, 23(2):339–??, November 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/23/2/339.

Roth:1964:SFC

[RS64c] L. E. Roth and Y. Shigenaka. The structure and formation of cilia and filaments in rumen protozoa. Journal of Cell Biology, 20(2):249–??, February 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/20/2/249.

Richardson:1962:SFM

[RT62] Thomas Richardson and A. L. Tappel. Swelling of fish mitochondria. Journal of Cell Biology, 13(1):43–??, April 1962. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/13/1/43.

Rosa:1963:UTB

[RT63] Charles G. Rosa and Kwan-Chung Tsou. The use of tetranitroblue tetrazolium for the cytochemical localization of succinic dehydrogenase. Journal of Cell Biology, 16(3):445–??, March 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/16/3/
445.

Ram:1963:PAF

[RTPM63] J. Sri Ram, Saroj S. Tawde, G. B. Pierce, and A. R. Midgley. Preparation of antibody-ferritin conjugates for immuno-electron microscopy. Journal of Cell Biology, 17(3):673–??, June 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/17/3/673.

Robertson:1962:CLG

[RV62] David M. Robertson and F. Stephen Vogel. Concentric lamination of glial processes in oligodendrogliomas. Journal of Cell Biology, 15(2):313–??, November 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/15/2/313.

Rosenbluth:1964:DEF

[RW64] Jack Rosenbluth and Steven L. Wissig. The distribution of exogenous ferritin in toad spinal ganglia and the mechanism of its uptake by neurons. Journal of Cell Biology, 23(2): 307–??, November 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/23/2/307.

Resnik:1960:OCC

[RWG60] Robert A. Resnik, Theodor Wanko, and Mary Ann Gavin. Observations on a Cytoplasmic Component in Lens Fibers. Journal of Cell Biology, 7(2):403–??, April 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/7/2/403.

Ryter:1960:EAM

[Ryt60] Antoinette Ryter. Étude au microscope électronique des transformations nucléaires de E. coli K12S et K12S(λ_{26}) après irradiation aux rayons ultraviolets et aux rayons X. Journal of Cell Biology, 8(2):399–??, October 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/8/2/399.

Schuel:1964:SICa

[SA64] Herbert Schuel and Norman G. Anderson. Studies on isolated cell components. Journal of Cell Biology, 21(3):309–??, June 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/21/3/
309.

Sakai:1960:SSGa

[Sak60a] Hikoichi Sakai. Studies on sulfhydryl groups during cell division of sea urchin egg. Journal of Cell Biology, 8(3):603–??, December 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
8/3/603.

Sakai:1960:SSGb

[Sak60b] Hikoichi Sakai. Studies on sulfhydryl groups during cell division of sea urchin egg. Journal of Cell Biology, 8(3):609–??, December 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
8/3/609.

Sano:1962:FST

[San62] Masao Sano. Further studies on the theta cell of the mouse anterior pituitary as revealed by electron microscopy, with special reference to the mode of secretion. Journal of Cell Biology, 15(1):85–??, October 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/15/1/85.

Satir:1962:ESP

[Sat62] Peter Satir. On the evolutionary stability of the 9 + 2 pattern. Journal of Cell Biology, 12(1):181–??, January 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/12/1/181.

Satir:1963:SC

[Sat63] Peter Satir. Studies on cilia. Journal of Cell Biology, 18(2): 345–??, August 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/18/2/345.

Shanthaveerappa:1962:PE

[SB62] T. R. Shanthaveerappa and G. H. Bourne. A perineural epithelium. Journal of Cell Biology, 14(2):343–??, August 1962. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/14/2/343.

Salpeter:1964:AEM

[SB64] M. M. Salpeter and L. Bachmann. Autoradiography with the electron microscope. Journal of Cell Biology, 22(2):469–??, August 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
22/2/469.

Sabatini:1963:CEM

[SBB63] David D. Sabatini, Klaus Bensch, and Russell J. Barrnett. Cytochemistry and electron microscopy. Journal of Cell Biology, 17(1):19–??, April 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/17/1/19.

Salisbury:1961:DNF

[SBdlTL61] G. W. Salisbury, W. J. Birge, L. de la Torre, and J. R. Lodge. Decrease in nuclear Feulgen-positive material (DNA) upon aging in in vitro storage of bovine spermatozoa. Journal of Cell Biology, 10(3):353–??, July 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/10/3/353.

Scarpelli:1963:FLN

[SC63a] Dante G. Scarpelli and Elson L. Craig. The fine localization of nucleoside triphosphatase activity in the retina of the frog. Journal of Cell Biology, 17(2):279–??, May 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/17/2/279.

Swartzendruber:1963:EMO

[SC63b] D. C. Swartzendruber and C. C. Congdon. Electron microscope observations on tingible body macrophages in mouse spleen. Journal of Cell Biology, 19(3):641–??, December 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/19/3/641.

Stasny:1964:ESO

[SC64] John T. Stasny and F. L. Crane. The effect of sonic oscillation on the structure and function of beef heart mitochondria. Journal of Cell Biology, 22(1):49–??, July 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/22/1/49.

Schreil:1964:SFA

[Sch64] Werner H. Schreil. Studies on the fixation of artificial and bacterial DNA plasms for the electron microscopy of thin sections. Journal of Cell Biology, 22(1):1–??, July 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/22/1/1.

Sabatini:1961:UZA

[SD61] David D. Sabatini and Eduardo D. P. De Robertis. Ultrastructural zonation of adrenocortex in the rat. Journal of Cell Biology, 9(1):105–??, January 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/9/1/105.

Smith:1964:SCI

[SD64] Janet A. Smith and Hector F. DeLuca. Structural changes in isolated liver mitochondria of rats during essential fatty acid deficiency. Journal of Cell Biology, 21(1):15–??, April 1964. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/21/1/15.

Seaman:1961:APA

[Sea61] Gerald R. Seaman. Acid phosphatase activity associated with phagotrophy in the ciliate, Tetrahymena. Journal of Cell Biology, 9(1):243–??, January 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/9/1/243.

Sedar:1961:EMOa

[Sed61a] Albert W. Sedar. Electron microscopy of the oxyntic cell in the gastric glands of the bullfrog (Rana Catesbiana). Journal of Cell Biology, 9(1):1–??, January 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/9/1/1.

Sedar:1961:EMOb

[Sed61b] Albert W. Sedar. Electron microscopy of the oxyntic cell in the gastric glands of the bullfrog, rana catesbiana. Journal of Cell Biology, 10(1):47–??, May 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/10/1/47.

Sedar:1962:EMO

[Sed62] Albert W. Sedar. Electron microscopy of the oxyntic cell in the gastric glands of the bullfrog, rana catesbiana. Journal of Cell Biology, 14(1):152–??, July 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/14/1/152.

Seed:1964:RBD

[See64] John Seed. The relations between DNA, RNA, and protein in normal embryonic cell nuclei and spontaneous tumour cell nuclei. Journal of Cell Biology, 20(1):17–??, January 1964. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/20/1/17.

Sedar:1961:CFS

[SF61] A. W. Sedar and M. H. F. Friedman. Correlation of the fine structure of the gastric parietal cell (dog) with functional activity of the stomach. Journal of Cell Biology, 11(2):349– ??, November 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/11/2/349.

Sedar:1964:ECD

[SF64] Albert W. Sedar and John G. Forte. Effects of calcium depletion on the junctional complex between oxyntic cells of gastric glands. Journal of Cell Biology, 22(1):173–??, July 1964. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/22/1/173.

Samuels:1964:CUC

[SGW64] Stanley Samuels, Nicholas K. Gonatas, and Martin Weiss. Chemical and ultrastructural comparison of synthetic and pathologic membrane system. Journal of Cell Biology, 21 (1):148–??, April 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/21/1/148.

Share:1960:PRL

[SH60] Leonard Share and Ronald W. Hansrote. Permeability of Rat Liver Microsomes to Sucrose and Carboxypoly-glucose in vitro. Journal of Cell Biology, 7(2):239–??, April 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/7/2/239.

Shatkin:1960:CCC

[Sha60] A. J. Shatkin. A Chlorophyll–Containing Cell Fraction from the Blue–Green Alga, Anabaena variabilis. Journal of Cell Biology, 7(3):583–??, June 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/7/3/583.

Shafiq:1963:EMSa

[Sha63a] S. Ahmad Shafiq. Electron microscopic studies on the indirect flight muscles of Drosophila Melanogaster. Journal of Cell Biology, 17(2):351–??, May 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/17/2/351.

Shafiq:1963:EMSb

[Sha63b] S. Ahmad Shafiq. Electron microscopic studies on the indirect flight muscles of Drosophila melanogaster. Journal of Cell Biology, 17(2):363–??, May 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/17/2/363.

Shalla:1964:AAT

[Sha64] Thomas A. Shalla. Assembly and aggregation of tobacco mosaic virus in tomato leaflets. Journal of Cell Biology, 21(2):253–??, May 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
21/2/253.

Shelton:1962:PSR

[She62] Emma Shelton. Prolonged survival of rabbit thoracic Duct lymphocytes in a diffusion chamber. Journal of Cell Biology, 12 (3):652–??, March 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/12/3/652.

Silk:1961:MIA

[SHSG61] M. H. Silk, A. O. Hawtrey, I. M. Spence, and J. H. S. Gear. A method for intracellular autoradiography in the electron microscope. Journal of Cell Biology, 10(4):577–??, August 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/10/4/577.

Seligman:1962:SIC

[SHSS62] Arnold M. Seligman, Jacob S. Hanker, Arlene R. Seaman, and Hannah E. Silberstein. Stability of the iodine–carbon bond of radioactive iodoalbumin in the 50 Kv electron beam. Journal of Cell Biology, 15(2):395–??, November 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/15/2/395.

Shapiro:1961:FSH

[SHT61] J. E. Shapiro, B. R. Hershenov, and G. S. Tulloch. The fine structure of Haematolœchus spermatozoan tail. Journal of Cell Biology, 9(1):211–??, January 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/9/1/211.

Sims:1964:IFH

[Sim64] R. T. Sims. The incorporation and fate of H^{3}-tyrosine in the hair cortex of rats observed by radioautography. Journal of Cell Biology, 22(2):403–??, August 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/22/2/403.

Siperstein:1963:IAP

[Sip63] Eleanor R. Siperstein. Identification of the adrenocorticotrophinproducing cells in the rat hypophysis by autoradiography. Journal of Cell Biology, 17(3):521–??, June 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/17/3/521.

Selman:1964:EMS

[SJ64] G. G. Selman and A. Jurand. An electron microscope study of the endoplasmic reticulum in newt notochord cells after disturbance with ultrasonic treatment and subsequent regeneration. Journal of Cell Biology, 20(1):175–??, January 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/20/1/175.

Simmel:1961:OUT

[SK61a] Eva B. Simmel and David A. Karnofsky. Observations on the uptake of tritiated thymidine in the pronuclei of fertilized sand dollar embryos. Journal of Cell Biology, 10(1):59–??, May 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/10/1/59.

Sisken:1961:TDS

[SK61b] Jesse E. Sisken and Riojun Kinosita. Timing of DNA synthesis in the mitotic cycle in vitro. Journal of Cell Biology, 9 (3):509–??, March 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/9/3/509.

Sheldon:1962:SC

[SK62] H. Sheldon and F. B. Kimball. Studies on cartilage. Journal of Cell Biology, 12(3):599–??, March 1962. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/12/3/599.

Stegwee:1963:HCR

[SKdBH63] D. Stegwee, E. C. Kimmel, J. A. de Boer, and S. Henstra. Hormonal control of reversible degeneration of flight muscle in the Colorado potato beetle, leptinotarsa decemlineata say (coleoptera). Journal of Cell Biology, 19(3):519–??, December 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/19/3/
519.

Spurlock:1963:TMM

[SKF63] Ben O. Spurlock, Victoria C. Kattine, and James A. Freeman. Technical modifications in maraglas embedding. Journal of Cell Biology, 17(1):203–??, April 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/17/1/203.

Skoglund:1961:FAS

[Sko61] C. R. Skoglund. Functional analysis of swim-bladder muscles engaged in sound production of the toadfish. Journal of Cell Biology, 10(4):187–??, August 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/10/4/187.

Smith:1960:CSE

[SL60] David S. Smith and Virginia C. Littau. Cellular specialization in the excretory epithelia of an insect, Macrosteles fascifrons stål (homoptera). Journal of Cell Biology, 8(1):103–??, September 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
8/1/103.

Selzman:1962:PCG

[SL62] H. M. Selzman and Robert A. Liebelt. Paneth cell granule of mouse intestine. Journal of Cell Biology, 15(1):136–??, October 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/15/1/
136.

Studzinski:1964:EPN

[SL64] G. P. Studzinski and R. Love. Effects of puromycin on the nucleoproteins of the HeLa cell. Journal of Cell Biology, 22(3):

493–??, September 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/22/3/493.

Slautterback:1963:CM

[Sla63] David B. Slautterback. Cytoplasmic microtubules. Journal of Cell Biology, 18(2):367–??, August 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/18/2/367.

Simpson:1964:BGN

[SM64] Nancy Sype Simpson and Robert Gilmore McKinnell. The burnsi gene as a nuclear marker for transplantation experiments in frogs. Journal of Cell Biology, 23(2):371–??, November 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/23/2/371.

Smith:1960:IFF

[Smi60] David S. Smith. Innervation of the fibrillar flight muscle of an insect: Tenebrio molitor (coleoptera). Journal of Cell Biology, 8(2):447–??, October 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/8/2/447.

Smith:1961:OFM

[Smi61a] David S. Smith. The organization of the flight muscle in a dragonfly, aeshna Sp. (odonata). Journal of Cell Biology, 11 (1):119–??, October 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/11/1/119.

Smith:1961:ROW

[Smi61b] David S. Smith. Reticular organizations within the striated muscle cell. Journal of Cell Biology, 10(4):61–??, August 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/10/4/61.

Smith:1961:SIF

[Smi61c] David S. Smith. The structure of insect fibrillar flight muscle. Journal of Cell Biology, 10(4):123–??, August 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/10/4/123.

Smith:1963:OIL

[Smi63a] David S. Smith. The organization and innervation of the luminescent organ in a firefly, photuris pennsylvanica (coleoptera). Journal of Cell Biology, 16(2):323–??, February 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/16/2/323.

Smith:1963:SFM

[Smi63b] David S. Smith. The structure of flight muscle sarcosomes in the blowfly Calliphora Erythrocephala (Diptera). Journal of Cell Biology, 19(1):115–??, October 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/19/1/115.

Sheldon:1964:EMA

[SMV64] H. Sheldon, J. M. McKenzie, and D. Van Nimwegan. Electron microscopic autoradiography. Journal of Cell Biology, 23 (1):200–??, October 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/23/1/200.

Steinert:1960:ECO

[SN60] M. Steinert and A. B. Novikoff. The existence of a cytostome and the occurrence of pinocytosis in the trypanosome, trypanosoma mega. Journal of Cell Biology, 8(2):563–??, October 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
8/2/563.

Savage:1960:RTC

[SNE60] J. R. K. Savage, G. J. Neary, and H. J. Evans. The rejoining time of chromatid breaks induced by gamma radiation in Vicia faba root tips at 3^{∘}C. Journal of Cell Biology, 7(1):79–??, February 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
7/1/79.

Simpson:1962:FSS

[SO62] F. O. Simpson and S. J. Oertelis. The fine structure of sheep myocardial cells; sarcolemmal invaginations and the transverse tubular system. Journal of Cell Biology, 12(1):91–??, January 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/12/1/
91.

Sorokin:1962:CFR

[Sor62] Sergei Sorokin. Centrioles and the formation of rudimentary cilia by fibroblasts and smooth muscle cells. Journal of Cell Biology, 15(2):363–??, November 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/15/2/363.

Siekevitz:1960:CSPa

[SP60a] Philip Siekevitz and George E. Palade. A Cytochemical Study on the Pancreas of the Guinea Pig. Journal of Cell Biology, 7(4):619–??, July 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/7/4/619.

Siekevitz:1960:CSPb

[SP60b] Philip Siekevitz and George E. Palade. A Cytochemical Study on the Pancreas of the Guinea Pig. Journal of Cell Biology, 7(4):631–??, July 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/7/4/631.

Siekevitz:1962:CSP

[SP62] Philip Siekevitz and George E. Palade. Cytochemical study on the pancreas of the guinea pig. Journal of Cell Biology, 13(2):217–??, May 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/13/2/217.

Stone:1964:CDD

[SP64] G. E. Stone and D. M. Prescott. Cell division and DNA synthesis in Tetrahymena Pyriformis deprived of essential amino acids. Journal of Cell Biology, 21(2):275–??, May 1964. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/21/2/275.

Smoller:1961:CIB

[SPB61] Carolyn G. Smoller, Dorothy R. Pitelka, and Howard A. Bern. Cytoplasmic inclusion bodies in cortisol–treated mammary tumors of C3H/Crgl mice. Journal of Cell Biology, 9(4):915–??, April 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/9/
4/915.

Sheldon:1960:SC

[SR60] Huntington Sheldon and Robert A. Robinson. Studies on cartilage. Journal of Cell Biology, 8(1):151–??, September 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/8/1/151.

Szollosi:1961:OSP

[SR61] Daniel G. Szollosi and Hans Ris. Observations on sperm penetration in the rat. Journal of Cell Biology, 10(2):275–??, June 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/10/2/
275.

Singer:1961:PSS

[SS61a] S. J. Singer and Anita F. Schick. The properties of specific stains for electron microscopy prepared by the conjugation of antibody molecules with ferritin. Journal of Cell Biology, 9(3):519–??, March 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/9/
3/519.

Stenger:1961:UMC

[SS61b] Richard J. Stenger and David Spiro. The ultrastructure of mammalian cardiac muscle. Journal of Cell Biology, 9(2): 325–??, February 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/9/2/325.

Schultz:1962:EDR

[SS62a] Richard L. Schultz and Phyllis W. Schultz. Enzyme distribution in the rat fetus and placenta following the adminstration of ethionine. Journal of Cell Biology, 15(2):211–??, November 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/15/2/
211.

Shepherd:1962:CEM

[SS62b] G. R. Shepherd and P. C. Sanders. The complementary effects of 2-mercaptoethanol and trypsin in the lysis of HeLa cells.

Journal of Cell Biology, 14(2):346–??, August 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/14/2/346.

Sheldon:1962:DAI

[SSK62] H. Sheldon, M. Silverberg, and I. Kerner. On the differing appearance of intranuclear and cytoplasmic glycogen in liver cells in glycogen storage disease. Journal of Cell Biology, 13(3):468– ??, June 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
13/3/468.

Sandritter:1960:WEG

[SSKD60] Von W. Sandritter, H. G. Schiemer, H. Kraus, and Und U. Dörrien. Das Wachstum von Einzelzellen in der Gewebekultur. Interferenzmikroskopische Untersuchungen an HeLa– Zellen. (German) [The growth of single cells in tissue culture. Interference microscopy investigation of HeLa cells]. Journal of Cell Biology, 7(3):579–??, June 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/7/3/579.

Stahl:1963:BM

[SSNB63] W. L. Stahl, J. C. Smith, L. M. Napolitano, and R. E. Basford. Brain mitochondria. Journal of Cell Biology, 19(2): 293–??, November 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/19/2/293.

Staubli:1963:NET

[Stä63] Willy Stäubli. A new embedding technique for electron microscopy, combining a water-soluble epoxy resin (durcupan) with water-insoluble araldite. Journal of Cell Biology, 16(1): 197–??, January 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/16/1/197.

Schuel:1964:SICb

[STA64] Herbert Schuel, Samuel R. Tipton, and Norman G. Anderson. Studies on isolated cell components. Journal of Cell Biology, 22 (2):317–??, August 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/22/2/317.

Steinert:1960:MAK

[Ste60] M. Steinert. Mitochondria associated with the kinetonucleus of trypanosoma mega. Journal of Cell Biology, 8(2):542–??, October 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/8/
2/542.

Stern:1961:PID

[Ste61] Herbert Stern. Periodic induction of deoxyribonuclease activity in relation to the mitotic cycle. Journal of Cell Biology, 9(2): 271–??, February 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/9/2/271.

Steinert:1964:CTM

[Ste64] M. Steinert. Le chondriome De trypanosoma mega. Journal of Cell Biology, 20(1):192–??, January 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/20/1/192.

Stoeckenius:1961:EMD

[Sto61] Walther Stoeckenius. Electron microscopy of DNA molecules “stained” with heavy metal salts. Journal of Cell Biology, 11 (2):297–??, November 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/11/2/297.

Stockinger:1962:EDM

[Sto62a] L. Stockinger. Elektronenoptische Darstellung Einer Makroglobulin-Retention in Plasmazellen Des Menschlichen Knochenmarks. (German) [Electron-optical imaging of macroglobulin retention in human bone marrow plasma cells]. Journal of Cell Biology, 15 (1):131–??, October 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/15/1/131.

Stoeckenius:1962:SEM

[Sto62b] Walther Stoeckenius. Some electron microscopical observations on liquid-crystalline phases in lipid-water systems. Journal of Cell Biology, 12(2):221–??, February 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/12/2/221.

Stoeckenius:1963:SON

[Sto63] Walther Stoeckenius. Some observations on negatively stained mitochondria. Journal of Cell Biology, 17(2):443–??, May 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/17/2/443.

Straus:1962:CIU

[Str62] W. Straus. Colorimetric investigation of the uptake of an intravenously injected protein (horseradish peroxidase) by rat kidney and effects of competition by egg white. Journal of Cell Biology, 12(2):231–??, February 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/12/2/231.

Strauss:1963:AFI

[Str63] Elliott W. Strauss. The absorption of fat by intestine of golden hamster in vitro. Journal of Cell Biology, 17(3):597–??, June 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/17/3/
597.

Straus:1964:COR

[Str64a] Werner Straus. Cytochemical observations on the relationship between lysosomes and phagosomes in kidney and liver by combined staining for acid phosphatase and intravenously injected horseradish peroxidase. Journal of Cell Biology, 20(3):497–??, March 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
20/3/497.

Straus:1964:OPP

[Str64b] Werner Straus. Occurrence of phagosomes and phago-lysosomes in different segments of the nephron in relation to the reabsorption, transport, digestion, and extrusion of intravenously injected horseradish peroxidase. Journal of Cell Biology, 21 (3):295–??, June 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/21/3/295.

Suganuma:1961:PMS

[Sug61] Atsushi Suganuma. The plasma membrane of staphylococcus aureus. Journal of Cell Biology, 10(2):292–??, June 1961. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/10/2/292.

Suganuma:1964:PMS

[Sug64] Atsushi Suganuma. The plasma membrane of Staphylococcus Aureus treated with HCl solutions. Journal of Cell Biology, 21(2):290–??, May 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/21/2/290.

Stempak:1964:ISM

[SW64] Jerome G. Stempak and Robert T. Ward. An improved staining method for electron microscopy. Journal of Cell Biology, 22(3): 697–??, September 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/22/3/697.

Sporn:1962:ICN

[SWD62] Michael B. Sporn, Theodor Wanko, and Wesley Dingman. The isolation of cell nuclei from rat brain. Journal of Cell Biology, 15(1):109–??, October 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/15/1/109.

Seno:1960:EMO

[SY60] Satimaru Seno and Kōyō Yoshizawa. Electron microscope observations on frozen-dried cells. Journal of Cell Biology, 8(3): 617–??, December 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/8/3/617.

Szollosi:1964:SFC

[Szo64] Daniel Szollosi. The structure and function of centrioles and their satellites in the jellyfish phialidium gregarium. Journal of Cell Biology, 21(3):465–??, June 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/21/3/465.

Taylor:1960:ADC

[Tay60] J. Herbert Taylor. Asynchronous Duplication of Chromosomes in Cultured Cells of Chinese Hamster. Journal of Cell Biology, 7(3):455–??, June 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/7/3/455.

Taylor:1962:RCP

[Tay62] A. Cecil Taylor. Responses of cells to pH changes in the medium. Journal of Cell Biology, 15(2):201–??, November 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/15/2/201.

Taylor:1963:RPS

[Tay63] Edwin W. Taylor. Relation of protein synthesis to the division cycle in mammalian cell cultures. Journal of Cell Biology, 19 (1):1–??, October 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/19/1/1.

Taylor:1964:ACM

[Tay64] J. Herbert Taylor. The arrangement of chromosomes in the mature sperm of the grasshopper. Journal of Cell Biology, 21(2):286–??, May 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/21/2/286.

Tice:1962:FSL

[TB62] Lois W. Tice and Russell J. Barrnett. Fine structural localization of adenosinetriphosphatase activity in heart muscle myofibrils. Journal of Cell Biology, 15(3):401–??, December 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/15/3/401.

Trujillo-Cenoz:1960:FSS

[TC60] O. Trujillo-Cenóz. The fine structure of a special type of nerve fiber found in the ganglia of Armadillidium vulgare (Crustacea– Isopoda). Journal of Cell Biology, 7(1):185–??, February 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/7/1/185.

Tonna:1964:EEI

[TC64] Edgar A. Tonna and Eugene P. Cronkite. The effects of extraperiosteal injections of blood components on periosteal cell proliferation. Journal of Cell Biology, 23(1):79–??, October 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/23/1/79.

Tahmisian:1961:IXR

[TD61] T. N. Tahmisian and R. L. Devine. The influence of X-rays on organelle induction and differentiation in grasshopper spermatogenesis. Journal of Cell Biology, 9(1):29–??, January 1961. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/9/1/29.

Thorp:1963:OIC

[TD63a] Frank K. Thorp and Albert Dorfman. The occurrence of intracellular chondroitin sulfate. Journal of Cell Biology, 18(1):13– ??, July 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
18/1/13.

Tooze:1963:OPS

[TD63b] John Tooze and Howard G. Davies. The occurrence and possible significance of haemoglobin in the chromosomal regions of mature erythrocyte nuclei of the newt triturus cristatus cristatus. Journal of Cell Biology, 16(3):501–??, March 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/16/3/501.

Telfer:1961:REL

[Tel61] William H. Telfer. The route of entry and localization of blood proteins in the oocytes of saturniid moths. Journal of Cell Biology, 9(4):747–??, April 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/9/4/747.

Todaro:1963:QSG

[TG63] George J. Todaro and Howard Green. Quantitative studies of the growth of mouse embryo cells in culture and their development into established lines. Journal of Cell Biology, 17(2):299– ??, May 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
17/2/299.

Travis:1964:SOR

[TG64] Dorothy F. Travis and Melvin J. Glimcher. The structure and organization of, and the relationship between the organic matrix and the inorganic crystals of embryonic bovine enamel.

Journal of Cell Biology, 23(3):447–??, December 1964. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/23/3/447.

Tunik:1961:DMA

[TH61] Bernard Tunik and Howard Holtzer. The distribution of muscle antigens in contracted myofibrils determined by fluoresceinlabeled antibodies. Journal of Cell Biology, 11(1):67–??, October 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/11/1/
67.

Tanzer:1964:EL

[TH64] Marvin L. Tanzer and Ronald D. Hunt. Experimental lathyrism. Journal of Cell Biology, 22(3):623–??, September 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/22/3/623.

Thaemert:1963:UDV

[Tha63] J. C. Thaemert. The ultrastructure and disposition of vesiculated nerve processes in smooth muscle. Journal of Cell Biology, 16(2):361–??, February 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/16/2/361.

Thoenes:1962:FSL

[Tho62] W. Thoenes. Fine structure of lipid granules in proximal tubule cells of mouse kidney. Journal of Cell Biology, 12(2): 433–??, February 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/12/2/433.

Thomas:1964:DCR

[Tho64a] P. K. Thomas. The deposition of collagen in relation to Schwann cell basement membrane during peripheral nerve regeneration. Journal of Cell Biology, 23(2):375–??, November 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/23/2/375.

Thomas:1964:ULM

[Tho64b] Richard S. Thomas. Ultrastructural localization of mineral matter in bacterial spores by microincineration. Journal of Cell Biology, 23(1):113–??, October 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/23/1/113.

Tedeschi:1963:PEO

[TJA63] Henry Tedeschi, Joseph M. James, and William Anthony. Photometric evidence for the osmotic behavior of rat liver microsomes. Journal of Cell Biology, 18(3):503–??, September 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/18/3/503.

Tomasz:1964:FSD

[TJO64] Alexander Tomasz, James D. Jamieson, and Elena Ottolenghi. The fine structure of diplococcus pneumoniae. Journal of Cell Biology, 22(2):453–??, August 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/22/2/453.

Takeya:1961:LEM

[TKMT61] Kenji Takeya, Masaatsu Koike, Ryoichi Mori, and Tadao Toda. Light and electron microscope studies of mycobacterium — mycobacteriophage interactions. Journal of Cell Biology, 11(2): 441–??, November 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/11/2/441.

Takeya:1961:FSP

[TMT^{+}61] Kenji Takeya, Ryoichi Mori, Tooru Tokunaga, Masaatsu Koike, and Kazuhito Hisatsune. Further studies on the paired fibrous structure of mycobacterial cell wall. Journal of Cell Biology, 9 (2):496–??, February 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/9/2/496.

Tonna:1961:CCS

[Ton61] Edgar A. Tonna. The cellular complement of the skeletal system studied autoradiographically with tritiated thymidine (h^{3} tdr) during growth and aging. Journal of Cell Biology, 9(4):813–??, April 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/9/
4/813.

Tooze:1964:MSC

[Too64] John Tooze. Measurements of some cellular changes during the fixation of amphibian erythrocytes with osmium tetroxide solutions. Journal of Cell Biology, 22(3):551–??, September 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/22/3/551.

Tormey:1963:FSC

[Tor63] John Mcd. Tormey. Fine structure of the ciliary epithelium of the rabbit, with particular reference to “infolded membranes,” “vesicles,” and the effects of diamox. Journal of Cell Biology, 17(3):641–??, June 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/17/3/641.

Tormey:1964:DMC

[Tor64] John McD. Tormey. Differences in membrane configuration between osmium tetroxide-fixed and glutaraldehyde-fixed ciliary epithelium. Journal of Cell Biology, 23(3):658–??, December 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/23/3/
658.

Tyler:1964:LMS

[TP64] W. S. Tyler and J. Pangborn. Laminated membrane surface and osmiophilic inclusions in avian lung epithelium. Journal of Cell Biology, 20(1):157–??, January 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/20/1/157.

Trier:1963:SSI

[Tri63] Jerry S. Trier. Studies on small intestinal crypt epithelium. Journal of Cell Biology, 18(3):599–??, September 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/18/3/599.

Trotter:1964:FSS

[Tro64] Nancy L. Trotter. A fine structure study of lipid in mouse liver regenerating after partial hepatectomy. Journal of Cell Biology, 21(2):233–??, May 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/21/2/233.

Takats:1963:TDP

[TS63a] S. T. Takats and R. M. S. Smellie. Thymidine degradation products in plant tissues labeled with tritiated thymidine. Journal of Cell Biology, 17(1):59–??, April 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/17/1/59.

Tamarin:1963:ADS

[TS63b] Arnold Tamarin and Leo M. Sreebny. An analysis of desmosome shape, size, and orientation by the use of histometric and densitometric methods with electron microscopy. Journal of Cell Biology, 18(1):125–??, July 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/18/1/125.

Thomas:1964:TAD

[TS64] P. K. Thomas and H. Sheldon. Tubular arrays derived from myelin breakdown during wallerian degeneration of peripheral nerve. Journal of Cell Biology, 22(3):715–??, September 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/22/3/715.

Tsuda:1961:ICE

[TT61] Seizo Tsuda and E. L. Tatum. Intracellular crystalline ergosterol in neurospora. Journal of Cell Biology, 11(1):171–??, October 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/11/1/
171.

Thornton:1964:CCB

[TT64] Robert M. Thornton and Kenneth V. Thimann. On a crystalcontaining body in cells of the oat coleoptile. Journal of Cell Biology, 20(2):345–??, February 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/20/2/345.

Thomas:1961:LDP

[TW61] Richard S. Thomas and Robley C. Williams. Localization of DNA and protein in tipula iridescent virus (tiv) by enzymatic digestion and electron microscopy. Journal of Cell Biology, 11 (1):15–??, October 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/11/1/15.

Tokuyasu:1960:FSR

[TY60] Kiyoteru Tokuyasu and Eichi Yamada. The Fine Structure of the Retina. V. Abnormal Retinal Rods and their Morphogenesis. Journal of Cell Biology, 7(1):187–??, February 1960. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/7/1/187.

Uchizono:1962:SPP

[Uch62] Koji Uchizono. The structure of possible photoreceptive elements in the sixth abdominal ganglion of the crayfish. Journal of Cell Biology, 15(1):151–??, October 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/15/1/151.

Uzman:1960:CNR

[UV60] Betty Geren Uzman and Gloria M. Villegas. A Comparison of Nodes of Ranvier in Sciatic Nerves with Node–Like Structures in Optic Nerves of the Mouse. Journal of Cell Biology, 7(4):761– ??, July 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/7/
4/761.

VanIterson:1961:SFR

[Van61] Woutera Van Iterson. Some features of a remarkable organelle in Bacillus Subtilis. Journal of Cell Biology, 9(1):183–??, January 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/9/
1/183.

vonderDecken:1963:LAA

[vdD63] Alexandra von der Decken. Labeling with ^{14}C amino acids of albumin-like protein by rat liver ribonucleoprotein particles. Journal of Cell Biology, 16(3):471–??, March 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/16/3/471.

Veratti:1961:IFS

[Ver61] Emilio Veratti. Investigations on the fine structure of striated muscle fiber. Journal of Cell Biology, 10(4):1–??, August 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/10/4/1.

vanIterson:1964:CLRb

[vIL64a] Woutera van Iterson and W. Leene. A cytochemical localization of reductive sites in a Gram-negative bacterium. Journal of Cell Biology, 20(3):377–??, March 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/20/3/377.

vanIterson:1964:CLRa

[vIL64b] Woutera van Iterson and W. Leene. A cytochemical localization of reductive sites in a Gram-positive bacterium. Journal of Cell Biology, 20(3):361–??, March 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/20/3/361.

Vitols:1961:SOMb

[VL61] Eberhards Vitols and Anthony W. Linnane. Studies on the oxidative metabolism of Saccharomyces Cerevisiae. Journal of Cell Biology, 9(3):701–??, March 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/9/3/701.

Vitols:1961:SOMa

[VNL61] Eberhards Vitols, R. J. North, and Anthony W. Linnane. Studies on the oxidative metabolism of Saccharomyces Cerevisiae. Journal of Cell Biology, 9(3):689–??, March 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/9/3/689.

Vial:1960:EMO

[VO60] Juan D. Vial and Héctor Orrego. Electron Microscope Observations on the Fine Structure of Parietal Cells. Journal of Cell Biology, 7(2):367–??, April 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/7/2/367.

Voelz:1964:SSB

[Voe64] Herbert Voelz. The “spindle-shaped body” in fibroblasts. Journal of Cell Biology, 20(2):333–??, February 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/20/2/333.

VanIterson:1961:OEM

[VR61] Woutera Van Iterson and C. F. Robinow. Observations with the electron microscope on the fine structure of the nuclei of two spherical bacteria. Journal of Cell Biology, 9(1):171–??, January 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/9/
1/171.

vanTubergen:1961:URE

[vT61] R. P. van Tubergen. The use of radioautography and electron microscopy for the localization of tritium label in bacteria. Journal of Cell Biology, 9(1):219–??, January 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/9/1/219.

Weiss:1960:SCM

[WA60] L. Weiss and J. A. Armstrong. Structural changes in mammalian cells associated with cooling to -79^{∘}C. Journal of Cell Biology, 7(4):673–??, July 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/7/4/673.

Watson:1961:MUI

[WA61] Michael L. Watson and William G. Aldridge. Methods for the use of indium as an electron stain for nucleic acids. Journal of Cell Biology, 11(2):257–??, November 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/11/2/257.

Walker:1963:SDV

[Wal63] Donald G. Walker. A survey of dehydrogenases in various epithelial cells in the rat. Journal of Cell Biology, 17(2):255–??, May 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
17/2/255.

Ward:1962:OPFa

[War62a] Robert T. Ward. The origin of protein and fatty yolk in rana pipiens. Journal of Cell Biology, 14(2):303–??, August 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/14/2/303.

Ward:1962:OPFb

[War62b] Robert T. Ward. The origin of protein and fatty yolk in rana pipiens. Journal of Cell Biology, 14(2):309–??, August 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/14/2/309.

Watson:1960:ENE

[Wat60] Michael L. Watson. The Extracellular Nature of Enamel in the Rat. Journal of Cell Biology, 7(3):489–??, June 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/7/3/489.

Watson:1962:OGA

[Wat62] Michael L. Watson. Observations on a granule associated with chromatin in the nuclei of cells of rat and mouse. Journal of Cell Biology, 13(1):162–??, April 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/13/1/162.

Watabe:1963:DTS

[Wat63] Norimitsu Watabe. Decalcification of thin sections for electron microscope studies of crystal-matrix relationships in mollusc shells. Journal of Cell Biology, 18(3):701–??, September 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/18/3/701.

Weissmann:1964:SL

[WBT64] Gerald Weissmann, Beverly Becher, and Lewis Thomas. Studies on lysosomes. Journal of Cell Biology, 22(1):115–??, July 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/22/1/
115.

Wellings:1961:MPD

[WD61] S. R. Wellings and K. B. Deome. Milk protein droplet formation in the Golgi apparatus of the C3H/Crgl mouse mammary epithelial cells. Journal of Cell Biology, 9(2):479–??, February 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
9/2/479.

Weber:1964:UCR

[Web64] Rudolf Weber. Ultrastructural changes in regressing tail muscles of Xenopus larvae at metamorphosis. Journal of Cell Biology, 22(2):481–??, August 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/22/2/481.

Wessells:1964:DSM

[Wes64] Norman K. Wessells. DNA synthesis, mitosis, and differentiation in pancreatic acinar cells in vitro. Journal of Cell Biology, 20(3):415–??, March 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/20/3/415.

Wetzel:1961:SPF

[Wet61] Bruce K. Wetzel. Sodium permanganate fixation for electron microscopy. Journal of Cell Biology, 9(3):711–??, March 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/9/3/711.

Wolken:1961:PSR

[WG61] J. J. Wolken and P. D. Gupta. Photoreceptor structures. the retinal cells of the cockroach eye. Journal of Cell Biology, 9 (3):720–??, March 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/9/3/720.

Weisblum:1962:CPA

[WHF62] Bernard Weisblum, Lawrence Herman, and Patrick J. Fitzgerald. Changes in pancreatic acinar cells during protein deprivation. Journal of Cell Biology, 12(2):313–??, February 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/12/2/313.

Wilsenach:1963:DCA

[Wil63] R. Wilsenach. Differentiation of the chloroplast of anthoceros. Journal of Cell Biology, 18(2):419–??, August 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/18/2/419.

Williamson:1964:AT

[Wil64] Joseph R. Williamson. Adipose tissue. Journal of Cell Biology, 20(1):57–??, January 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/20/1/57.

Wischnitzer:1960:OAL

[Wis60a] S. Wischnitzer. Observations on the annulate lamellae of immature amphibian oocytes. Journal of Cell Biology, 8(2):558–??, October 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/8/
2/558.

Wissig:1960:ASF

[Wis60b] Steven L. Wissig. The Anatomy of Secretion in the Follicular Cells of the Thyroid Gland. Journal of Cell Biology, 7(3):419– ??, June 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/7/
3/419.

Wissig:1963:ASF

[Wis63] Steven L. Wissig. The anatomy of secretion in the follicular cells of the thyroid gland. Journal of Cell Biology, 16(1):93–??, January 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
16/1/93.

Wallace:1963:SAY

[WK63] Robin A. Wallace and Shuichi Karasaki. Studies on amphibian yolk. Journal of Cell Biology, 18(1):153–??, July 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/18/1/153.

Weibel:1964:MST

[WK64] Ewald R. Weibel and Bruce W. Knight. A morphometric study on the thickness of the pulmonary air-blood barrier. Journal of Cell Biology, 21(3):367–??, June 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/21/3/367.

Wilson:1963:MCL

[WL63] J. Walter Wilson and Elizabeth H. Leduc. Mitochondrial changes in the liver of essential fatty acid-deficient mice. Journal of Cell Biology, 16(2):281–??, February 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/16/2/281.

Warshawsky:1963:SMP

[WLD63] H. Warshawsky, C. P. Leblond, and B. Droz. Synthesis and migration of proteins in the cells of the exocrine pancreas as revealed by specific activity determination from radioautographs. Journal of Cell Biology, 16(1):1–??, January 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/16/1/1.

Whaley:1963:GAC

[WM63] W. Gordon Whaley and Hilton H. Mollenhauer. The Golgi apparatus and cell plate formation — a postulate. Journal of Cell Biology, 17(1):216–??, April 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/17/1/216.

Wachtel:1961:EMH

[WMG61] Allen Wachtel, Robert Mathewson, and Harry Grundfest. Electron microscopic and histochemical comparison of the two types of electroplaques of narcine brasiliensis. Journal of Cell Biology, 11(3):663–??, December 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/11/3/663.

Whaley:1960:SON

[WML60] W. Gordon Whaley, Hilton H. Mollenhauer, and James H. Leech. Some observations on the nuclear envelope. Journal of Cell Biology, 8(1):233–??, September 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/8/1/233.

Wakid:1961:EEI

[WN61] N. W. Wakid and Dorothy M. Needham. Effect of estradiol injection on the cytoplasmic fractions of the myometrium in the ovariectomized rat. Journal of Cell Biology, 10(1):136–??, May 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/10/1/
136.

Wooding:1964:DSW

[WN64] F. B. P. Wooding and D. H. Northcote. The development of the secondary wall of the xylem in Acer Pseudoplatanus. Journal of Cell Biology, 23(2):327–??, November 1964. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/23/2/327.

Wyss:1961:DGA

[WNS61] Orville Wyss, Marilyn G. Neumann, and M. D. Socolofsky. Development and germination of the azotobacter cyst. Journal of Cell Biology, 10(4):555–??, August 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/10/4/555.

Wolf:1964:DNT

[Wol64] Merrill K. Wolf. Differentiation of neuronal types and synapses in myelinating cultures of mouse cerebellum. Journal of Cell Biology, 22(1):259–??, July 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/22/1/259.

Warth:1963:CSB

[WOM63a] A. D. Warth, D. F. Ohye, and W. G. Murrell. The composition and structure of bacterial spores. Journal of Cell Biology, 16 (3):579–??, March 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/16/3/579.

Warth:1963:LCS

[WOM63b] A. D. Warth, D. F. Ohye, and W. G. Murrell. Location and composition of spore mucopeptide in bacillus species. Journal of Cell Biology, 16(3):593–??, March 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/16/3/593.

Weibel:1964:NCC

[WP64a] Ewald R. Weibel and George E. Palade. New cytoplasmic components in arterial endothelia. Journal of Cell Biology, 23(1): 101–??, October 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/23/1/101.

Wolstenholme:1964:CDS

[WP64b] D. R. Wolstenholme and W. Plaut. Cytoplasmic DNA synthesis in Amoeba Proteus. Journal of Cell Biology, 22(3): 505–??, September 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/22/3/505.

Wimber:1960:ATI

[WQSW60] Donald E. Wimber, Henry Quastler, Otto L. Stein, and Doris R. Wimber. Analysis of tritium incorporation into individual cells by autoradiography of squash preparations. Journal of Cell Biology, 8(2):327–??, October 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/8/2/327.

Weintraub:1964:EMS

[WR64] M. Weintraub and H. W. J. Ragetli. An electron microscope study of tobacco mosaic virus lesions in Nicotiana Glutinosa L. Journal of Cell Biology, 23(3):499–??, December 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/23/3/499.

Woodard:1961:NAP

[WRS61] John Woodard, Ellen Rasch, and Hewson Swift. Nucleic acid and protein metabolism during the mitotic cycle in Vicia Faba. Journal of Cell Biology, 9(2):445–??, February 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/9/2/445.

Walker:1961:FFC

[WS61a] Donald G. Walker and Arnold M. Seligman. Formalin fixation in the cytochemical demonstration of succinic dehydrogenase of mitochondria. Journal of Cell Biology, 9(2):415–??, February 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
9/2/415.

Wollenberger:1961:MAM

[WS61b] Albert Wollenberger and Wolfgang Schulze. Mitochondrial alterations in the myocardium of dogs with aortic stenosis. Journal of Cell Biology, 10(2):285–??, June 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/10/2/285.

Walker:1963:UFF

[WS63] Donald G. Walker and Arnold M. Seligman. The use of formalin fixation in the cytochemical demonstration of succinic and dpnand tpn-dependent dehydrogenases in mitochondria. Journal of Cell Biology, 16(3):455–??, March 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/16/3/455.

Worden:1964:PPB

[WS64] Patricia Barron Worden and W. R. Sistrom. The preparation and properties of bacterial chromatophore fractions. Journal of Cell Biology, 23(1):135–??, October 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/23/1/135.

Wollman:1964:LEA

[WSB64] Seymour H. Wollman, Samuel S. Spicer, and Marvin S. Burstone. Localization of esterase and acid phosphatase in granules and colloid droplets in rat thyroid epithelium. Journal of Cell Biology, 21(2):191–??, May 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/21/2/191.

Wiener:1964:SEG

[WSL64] Joseph Wiener, David Spiro, and Werner R. Loewenstein. Studies on an epithelial (gland) cell junction. Journal of Cell Biology, 22(3):587–??, September 1964. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/22/3/587.

Weier:1962:GSF

[WT62] T. Elliot Weier and William W. Thomson. The grana of starchfree chloroplasts of nicotiana rustica. Journal of Cell Biology, 13(1):89–??, April 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/13/1/89.

Watabe:1961:SSF

[WW61] Norimitsu Watabe and Karl M. Wilbur. Studies on Shell formation. Journal of Cell Biology, 9(4):761–??, April 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/9/4/761.

Yamada:1960:CFW

[Yam60] Eichi Yamada. Collagen Fibrils within the Renal Glomerulus. Journal of Cell Biology, 7(2):407–??, April 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/7/2/407.

Yamagishi:1962:IBN

[Yam62] Hideo Yamagishi. Interaction between nucleic acids and berberine sulfate. Journal of Cell Biology, 15(3):589–??, December 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/15/3/
589.

Yamamoto:1963:TUM

[Yam63a] Toshiyuki Yamamoto. On the thickness of the unit membrane. Journal of Cell Biology, 17(2):413–??, May 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/17/2/413.

Yamamoto:1963:SOF

[Yam63b] Toshiyuki Yamamoto. Some observations on the fine structure of the sympathetic ganglion of bullfrog. Journal of Cell Biology, 16(1):159–??, January 1963. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/16/1/159.

Yasuzumi:1960:SARa

[Yas60] G. Yasuzumi. Spermatogenesis in Animals as Revealed by Electron Microscopy. Journal of Cell Biology, 7(1):73–??, February 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
7/1/73.

Yasuzumi:1962:EMX

[Yas62a] G. Yasuzumi. Electron microscopy and X-ray scanning microanalysis of needle biopsy material from human liver. Journal of Cell Biology, 14(3):421–??, September 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/14/3/421.

Yasuzumi:1962:XRS

[Yas62b] G. Yasuzumi. X-ray scanning microanalysis of elemental iron localized in testicular nutritive cells of cipangopaludina malleata reeve. Journal of Cell Biology, 14(3):495–??, September 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/14/3/495.

Young:1962:CMS

[YFJ62] I. Elizabeth Young and Philip C. Fitz-James. Chemical and morphological studies of bacterial spore formation. Journal of Cell Biology, 12(1):115–??, January 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/12/1/115.

Young:1962:CPS

[You62] Richard W. Young. Cell proliferation and specialization during endochondral osteogenesis in Young rats. Journal of Cell Biology, 14(3):357–??, September 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/14/3/357.

Yasuzumi:1960:SARb

[YTT60] G. Yasuzumi, Hiroaki Tanaka, and Osamu Tezuka. Spermatogenesis in Animals as Revealed by Electron Microscopy. Journal of Cell Biology, 7(3):499–??, June 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/7/3/499.

Zalokar:1961:EMC

[Zal61] M. Zalokar. Electron microscopy of centrifuged hyphae of neurospora. Journal of Cell Biology, 9(3):609–??, March 1961. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/9/3/609.

Zacks:1961:OFS

[ZB61a] S. I. Zacks and J. M. Blumberg. Observations on the fine structure and cytochemistry of mouse and human intercostal neuromuscular junctions. Journal of Cell Biology, 10(4):517–??, August 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
10/4/517.

Zobel:1961:ES

[ZB61b] C. Richard Zobel and Michael Beer. Electron stains. Journal of Cell Biology, 10(3):335–??, July 1961. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/10/3/335.

Zeigel:1962:SBM

[ZD62] R. F. Zeigel and A. J. Dalton. Speculations based on the morphology of the Golgi systems in several types of proteinsecreting cells. Journal of Cell Biology, 15(1):45–??, October 1962. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/15/1/
45.

Zebrun:1960:EMO

[ZM60] William Zebrun and Hilton H. Mollenhauer. Electron Microscopic Observations on Mitochondria of Rat Testes Fixed in Potassium Permanganate. Journal of Cell Biology, 7(2):311–??, April 1960. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/7/
2/311.